### v1

# Construction of a Basis of Kähler differentials for generic plane branches

Abstract.
Let ${\mathcal{C}}$ be a fixed equisingularity class of irreducible germs of complex analytic plane curves. We compute a basis of the ${\mathbb{C}}[[x]]$-module of Kähler differentials for generic $\Gamma\in{\mathcal{C}}$, algorithmically, and study its behaviour under blow-up.
As an application, we give an alternative proof for a formula of Genzmer, that provides the dimension of the generic component of the moduli of analytic classes in the equisingularity class of $\Gamma$.
Key words and phrases:
plane branch, approximate roots, moduli, Kähler differentials, flows of holomorphic vector fields
2010 Mathematics Subject Classification:
32S05, 32S65, 14H20, 32G13
1. Introduction
What is the structure of the generic component of the analytic moduli of a singular plane branch? What is its behavior under blow-up? These are the two questions we address in this work, by means of explicitly and algorithmically computing a basis of the semi-module associated to a generic singularity in a topological class.
Let us explain our work in detail. We are interested in problems regarding the analytic classification of irreducible germs of complex analytic plane curves, the so called plane branches. The topological classification of germs of plane curves has a simple classical solution (cf. [5, Theorem 8.4.21]). Consider, in $(\mathbb{C}^{2},0)$ with coordinates $(x,y)$, an irreducible germ of complex analytic plane curve $\Gamma_{\circ}$. Given a plane branch $\Gamma$, it is said to be topologically (resp. analytically) conjugated to $\Gamma_{\circ}$, denoted $\Gamma\sim_{\mathrm{top}}\Gamma_{\circ}$ (resp. $\Gamma\sim_{\mathrm{ana}}\Gamma_{\circ}$) if there exists a germ of homeomorphism (resp. biholomorphism) $\phi$ defined in a neighborhood of the origin of $\mathbb{C}^{2}$ such that $\phi(\Gamma_{\circ})=\Gamma$. It is well-known that the set
$\tilde{\mathcal{C}}:=\{\Gamma:\ \Gamma\ {\rm is\ a\ plane\ branch\ such\ that}\ \Gamma_{\circ}\sim_{\mathrm{top}}\Gamma\}$
consists of the plane branches with the same graph of resolution of singularities as $\Gamma_{\circ}$ and hence it is called the equisingularity class of $\Gamma$. Alternatively, $\Gamma\in\tilde{\mathcal{C}}$ if and only if the semigroup of values ${\mathcal{S}}(\Gamma_{\circ})$ and ${\mathcal{S}}(\Gamma)$ coincide, where
${\mathcal{S}}(\Gamma)=\left\{\nu_{\Gamma}(f):=\operatorname{ord}_{t}(f(\Gamma(t)))\ :\ f(x,y)\in\mathbb{C}[[x,y]]\right\}$
and $\Gamma(t)$ is an irreducible Puiseux parametrization of $\Gamma$.
It is natural to refine this topological classification by considering the analytic classification, both providing a complete system of analytic invariants and describing the properties of the moduli space $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$. The salience of this problem was noted by Zariski in his 1973 Course at l’École Polytechnique in Paris [19], where he also described the moduli space for several topological classes. The problem of analytic classification of plane curves was solved by Hefez, Hernandes and Rodrigues Hernandes in a series of papers [14, 15, 16] starting in 2011. Let us focus in the case of plane branches [14]. The main invariant of the classification is the semimodule
$\Lambda_{\Gamma}=\big{\{}\nu_{\Gamma}(\omega):=\operatorname{ord}_{t}(\Gamma(t)^{\ast}\omega)+1\ |\ \big{.}\big{.}\omega\in\hat{\Omega}(\mathbb{C}^{2},0)\big{\}},$
where $\hat{\Omega}(\mathbb{C}^{2},0):=\mathbb{C}[[x,y]]dx+\mathbb{C}[[x,y]]dy$ is the set of formal $1$-forms and we define $\operatorname{ord}_{t}(h(t)dt)=\operatorname{ord}_{t}(h(t))$ for $h(t)\in{\mathbb{C}}[[t]]$. Thus, in order to continue the study of the properties of the moduli spaces $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$, it is interesting not only to describe the semimodule, but also to explicitly compute bases of $1$-forms generating it in some sense.
The following, by no means exhaustive, is a summary of references in the literature. The moduli space $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$ of a plane branch $\Gamma_{\circ}$ can be identified with the set of orbits of a regular action of a connected solvable algebraic group $G$ on an affine space ${\mathbb{C}}^{p}$ by a theorem of Ebey [10]. As a consequence, that moduli space can also be identified with a constructible subset of an affine space. There exists a Zariski open subset $U$ of ${\mathbb{C}}^{p}$ consisting of orbits of $G$ such that all orbits of $G$ in $U$ have the same dimension $d$ by a theorem of Rosenlicht [18]. Indeed, $U/G$ can be identified with a Zariski open subset of an algebraic variety $W$ of dimension $p-d$ (cf. [19, Chapter VI.1]) that represents the generic component of the moduli space $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$. The generic dimension $p-d$ of the moduli was studied in the case of one Puiseux characteristic exponent by Delorme [9] and Briançon, Granger and Maisonobe [4] who gave an algorithm for its calculation. Genzmer provided a general formula for the generic dimension of the moduli space of any plane branch that depends on simple combinatorial data associated to the topological class of $\Gamma_{\circ}$ [13].
Delorme studied the generation of the semimodule $\Lambda_{\Gamma_{\circ}}$ over the semigroup $\mathcal{S}(\Gamma_{\circ})$ in the case of a unique Puiseux characteristic exponent, revealing interesting properties for the generating $1$-forms [9]. This approach was adopted by Cano, Corral and Senovilla-Sanz in [6], where they describe geometrical properties of Delorme’s generating $1$-forms from the viewpoint of foliation theory.
Consider, for simplicity, the subset $\mathcal{C}$ of the topological class $\tilde{\mathcal{C}}$ of $\Gamma_{\circ}$ of plane branches $\Gamma$ whose tangent cone is different from $x=0$. It is obvious that every analytic class of an element of $\tilde{\mathcal{C}}$ intersects $\mathcal{C}$ and hence $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}=\mathcal{C}/\sim_{\mathrm{ana}}$. Given $\Gamma\in\mathcal{C}$, set ${\mathcal{I}}_{\Gamma}=\{\omega\in\hat{\Omega}(\mathbb{C}^{2},0):\Gamma(t)^{*}\omega\equiv 0\}$. We shall compute a basis of the ${\mathbb{C}}[[x]]$-module of Kähler differentials of $\Gamma$; more precisely, if $n$ is the multiplicity of $\Gamma_{\circ}$ at the origin, we shall find an Apery basis of $\Gamma$, i.e. germs of $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ such that
  • $\hat{\Omega}(\mathbb{C}^{2},0)=\mathbb{C}[[x]]\Omega_{1,\Gamma}\oplus\ldots\oplus\mathbb{C}[[x]]\Omega_{n,\Gamma}\oplus{\mathcal{I}}_{\Gamma}$;
  • The values $\nu_{\Gamma}(\Omega_{1,\Gamma}),\ldots,\nu_{\Gamma}(\Omega_{n,\Gamma})$ define an Apery set of $\Lambda_{\Gamma}$ [2], i.e. the classes of $\nu_{\Gamma}(\Omega_{j,\Gamma})$ and $\nu_{\Gamma}(\Omega_{k,\Gamma})$ modulo $n$ are different for all $1\leq j<k\leq n$.
To begin with, define $\overline{\Omega}_{2j+1}=y^{j}dx$ and $\overline{\Omega}_{2j+2}=y^{j}dy$ for any $j\in\mathbb{Z}_{\geq 0}$. We prove that can choose the elements of the basis $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ with sharp extra properties for any generic $\Gamma\in\mathcal{C}$:
minimality:
$\Omega_{j,\Gamma}\in\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$;
monotonicity:
$\nu_{\Gamma}(\Omega_{j,\Gamma})<\nu_{\Gamma}(\Omega_{k,\Gamma})$ for all $1\leq j<k\leq n$.
Let us comment these properties. By the properties of an Apery basis, it follows that $\oplus_{k=1}^{j}{\mathbb{C}}[[x]]\Omega_{k,\Gamma}$ is a free $\mathbb{C}[[x]]$-module of rank $j$. Moreover, we have a filtration
$\mathbb{C}[[x]]\overline{\Omega}_{1}\subset\mathbb{C}[[x]]\overline{\Omega}_{1}\oplus\mathbb{C}[[x]]\overline{\Omega}_{2}\subset\ldots\subset\oplus_{k=1}^{j}\mathbb{C}[[x]]\overline{\Omega}_{k}\subset\ldots$
of $\mathbb{C}[[x]]$-modules of rank $1,2,\ldots,j,\ldots$ The most economical scenario for the Apery basis with respect to the filtration happens when
$\Omega_{j,\Gamma}\in\oplus_{k=1}^{j}\mathbb{C}[[x]]\overline{\Omega}_{k}\setminus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$
for any $1\leq j\leq n$. This property derives from the minimality property. Thus, the filtration gives an order on the $1$-forms of an Apery basis. Monotonicity implies that this is the same order that is given by the natural valuation.
Let us remark that the minimality property is related to “maximal contact” properties in the spirit of Hironaka [17] for the resolution of algebraic singularities and of Aroca, Hironaka and Vicente [3] for the complex analytic case. Maximal contact varieties are essential for the resolution of singularities in characteristic 00. The $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ have maximal contact in the following sense: $\nu_{\Gamma}(\Omega_{j,\Gamma})$ is the maximum of the $\Gamma$-orders of the $1$-forms in $\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$ (Proposition 3.7).
There are two other properties that will be key in our construction of an Apery basis for a generic $\Gamma\in\mathcal{C}$:
simplicity:
We divide our algorithm of construction of the Apery basis in $g+1$ levels, where $g$ is the genus of $\Gamma_{\circ}$, i.e. its number of Puiseux characteristic exponents. We define the terminal family of $1$-forms of level 00 as the set consisting of the elementary $1$-forms $dx$ and $dy$. Given a terminal family of level $l$, the initial family of level $l+1$, with $0\leq l<g$, is obtained by combining that terminal family of level $l$ with data of topological flavor. Then we describe a straightforward algorithm that provides a terminal family of level $l+1$. At the end of level $g$ we obtain an Apery basis.
richness:
The algorithm of construction of an Apery basis preserves a very rich structure. For instance, the $\Gamma$-values of all the $1$-forms involved in the algorithm can be calculated in terms of combinatorial data of the equisingularity class of $\Gamma_{\circ}$ (Propositions 3.3 and 3.4). In particular, we obtain a combinatorial algorithm that provides the generic semimodule $\Lambda_{\Gamma}$ for $\Gamma\in\mathcal{C}$.
Being more precise, the initial family of level $l+1$, with $0\leq l<g$, is obtained by combining the terminal family of level $l$ with the approximate root $f_{l+1,\Gamma}$ (cf. [1]) whose $\Gamma$-order is the generator $\overline{\beta}_{l+1}$ of the semigroup $\mathcal{S}(\Gamma)$ (sections 2.4 and 3.5.2). The main problem is that the $\Gamma$-orders of the $1$-forms in the initial family of level $l+1$ are not pairwise different modulo $n$. Our algorithm is intended to provide a terminal family of level $l+1$ with that property, and the others above. Philosophically, our method factors the contribution of the approximate roots $f_{l,\Gamma}=0$, with $1\leq l\leq g$, in the construction of the semimodule, one by one. Let us mention that, in a similar spirit, de Abreu and Hernandes, give some relations between the semimodules of approximate roots of a plane branch $\Gamma$ and the semimodule of $\Gamma$ [8].
In order to understand the properties of the $1$-forms that appear along the steps of the algorithm, we consider the contributions of the coefficients of the Puiseux expansion of $\Gamma$ one by one, for generic $\Gamma\in\mathcal{C}$. A similar idea was considered recently by Casas-Alvero [7]. This viewpoint leads us to define two new concepts of independent interest, namely the leading variable (or coefficient) and the $c$-sequence associated to some $\omega\in\hat{\Omega}(\mathbb{C}^{2},0)$ (Definition 3.8). They provide sufficient control to calculate the generic semigroup $\Lambda_{\Gamma}$, for $\Gamma\in\mathcal{C}$, without calculating the $1$-forms in the Apery basis or any other $1$-forms along the process.
The existence of an Apery basis with the minimality and monotonicity properties, for generic $\Gamma\in\mathcal{C}$, is proved in Theorem 3.1. Simplicity and richness are natural byproducts of our algorithm (Propositions 3.3 and 3.4). Moreover, the Apery bases given by Theorem 3.1 depend polynomially on $\Gamma$ for generic $\Gamma\in\mathcal{C}$.
Our second aim is to relate the semimodule of $\Gamma$ with analogue concepts for the strict transform $\tilde{\Gamma}$ of $\Gamma$ by the blow-up of the origin. We are forced to consider a slightly more general setting and to study the semimodule $\Lambda_{\Gamma}^{E}$ (cf. Definition 2.3) of orders of Kähler differentials on $\Gamma$ associated to $1$-forms that preserve $E$, where $E$ is a normal crossings divisor at the origin of $\mathbb{C}^{2}$. The construction is similar to the previous one, where $E=\emptyset$, and is studied in section 4. Theorem 4.1 is the analogue of Theorem 3.1 for this setting. Minimality, monotonicity, simplicity and richness are all used to provide a simple comparison of the semimodules $\Lambda_{\Gamma}^{E}$ and $\Lambda_{\tilde{\Gamma}}^{\tilde{E}}$ (Proposition 6.1), where $\tilde{E}$ is the total transform of $E$.
Let us provide an application of our construction. Consider the sequence $P_{0}=(0,0)\in\mathbb{C}^{2},P_{1},P_{2}\ldots$ of infinitely near points of a plane branch $\Gamma$ (Remark 5.3). Let $\pi_{j}$ be the blow-up of the point $P_{j-1}$. Denote by $\Gamma_{j}$ the strict transform of $\Gamma$ by $\tau_{j}:=\pi_{1}\circ\ldots\circ\pi_{j}$ where $\Gamma_{0}=\Gamma$ and $\tau_{0}=\mathrm{id}$. Fix $j\in\mathbb{Z}_{\geq 0}$. Now, let us consider actions provided by the flow of a vector field on $\Gamma_{j}$. Actions of vector fields as a tool for the study of the analytic classification problem were introduced by the first author in [11]. Consider a holomorphic vector field $X$ defined in a neighborhood of $P_{j}$ such that $E_{j}:=\tau_{j}^{-1}(0,0)$ is $X$-invariant but $\Gamma_{j}$ is not. We obtain a family of curves $\Gamma_{j}^{\epsilon}=\mathrm{exp}(\epsilon X)(\Gamma_{j})$ defined for $\epsilon\in\mathbb{C}$ close to 00. By blow-down by $\tau_{j}$, we get a family ${(\Gamma^{\epsilon})}_{\epsilon\in(\mathbb{C},0)}$ in the equisingularity class of $\Gamma$. The Puiseux expansion of $\Gamma^{\epsilon}$ is of the form
$\Gamma^{\epsilon}(t)=\left(t^{n},\sum a_{\beta}(\epsilon)t^{\beta}\right)$
where the map $(\epsilon,t)\mapsto\Gamma^{\epsilon}(t)$ is holomorphic. Let $\beta$ be the first exponent such that $a_{\beta}(\epsilon)$ is non-constant. Any $\beta$ obtained this way, by varying $X$, is called a flow-fanning exponents at index $j$, and its set is denoted by $\Theta_{j}(\Gamma)$ (Definition 5.6). The generic dimension of the moduli is equal to
$\sum_{j\geq 1}\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma))$
by Remark 5.11 and Proposition 5.4, where we assume that $\Gamma$ is generic in its equisingularity class $\mathcal{C}$ (otherwise we replace $\Gamma$ with a generic element). Such formula is obtained by introducing a correspondence between exponents of the Puiseux expansion of $\Gamma$ and divisors of the desingularization process (Proposition 5.1). We get a geometric interpretation of the generic dimension of the moduli in which $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ is the set of exponents of a Puiseux parametrization that can be changed by the action of a vector field defined in a neighborhood of $P_{j}$ but can not be changed by the action of a vector field defined in a neighborhood of any of the points $P_{0},P_{1},\ldots,P_{j-1}$.
It is possible to compute $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ in terms of the semimodules $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 5.4) by applying the results on the correspondence between orders of Kähler differentials on $\Gamma$ and the tangency orders of vector fields with $\Gamma$ described in [12]. For the case where $\Gamma$ is generic in its equisingularity class, our construction of the Apery basis and its properties allow to calculate in parallel $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 6.2), to obtain $\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma))$ for $j\in\mathbb{Z}_{\geq 1}$. This term is indeed the $j$th-term $\sigma(\nu_{P_{j-1}}(\Gamma_{j-1}\cup\tau_{j}^{-1}(0,0)))$ in Genzmer’ formula (Proposition 6.1) (cf. Definitions 6.2 and 2.1). So, we obtain a geometric interpretation of Genzmer’s formula for the generic dimension of the moduli space $\mathcal{C}/\sim_{\mathrm{ana}}$ (Theorem 6.1).
We introduce the setting of the paper in section 2. The construction of the Apery basis of a generic plane branch in an equisingularity class is carried out in section 3. This is generalized for the case of Kähler differentials preserving a normal crossings divisor $E$ in section 4. Section 5 is devoted to introducing the correspondence between exponents of the Puiseux expansion and divisors of the desingularization process of a plane branch $\Gamma$ and to use such a setting to introduce the sets of flow-fanning exponents and describe their properties. Section 6 gives, as a direct application of our results, an alternative proof of Genzmer’s formula.
2. Equisingularity class of a pair $(\Gamma_{\circ},E)$
We consider, in $(\mathbb{C}^{2},0)$ with coordinates $(x,y)$, an irreducible germ of complex analytic plane curve $\Gamma_{\circ}$ and a normal crossings divisor $E\subset\{xy=0\}$. It is well-known that the equisingularity class $\tilde{\mathcal{C}}$ of $(\Gamma_{\circ},E)$ is equivalent to the equisingularity class of $\Gamma_{\circ}$ together with the intersection numbers $(\Gamma_{\circ},D)_{(0,0)}$ at the origin for each irreducible component $D$ of $E$ (cf. [5, Theorem 8.4.21]).
Definition 2.1.
We consider the notation $\nu_{P}(\Gamma)$ for the multiplicity of a germ of curve $\Gamma$ at $P$. Sometimes we just write $\nu(\Gamma)$ when $P$ is implicit.
2.1. The set of exponents of an equisingularity class
Define $n$ as the multiplicity $\nu(\Gamma_{\circ})$ of $\Gamma_{\circ}$ at the origin if $\{x=0\}\not\subset E$, otherwise we define $n$ as the intersection number $(\Gamma_{\circ},x=0)_{(0,0)}$. We define
$\beta_{0}=\max\{(\overline{\Gamma},y=0)_{(0,0)}\ |\ (\overline{\Gamma},E)\in\tilde{\mathcal{C}}\ \mathrm{and}\ (\overline{\Gamma},x=0)_{(0,0)}=n\}.$
For any curve $\overline{\Gamma}\in\tilde{\mathcal{C}}$, there exists an irreducible curve $\Gamma$ that is analytically conjugated to $\overline{\Gamma}$ by a map preserving every irreducible component of $E$ and admits a parametrization of the form
(1) $\Gamma(t)=\left(t^{n},\sum_{\beta\geq\beta_{0}}a_{\beta,\Gamma}t^{\beta}\right).$
So, up to a change of coordinates, we suppose that $\Gamma_{\circ}$ has a Puiseux expansion of the form (1). The calculation of the $\mathbb{C}[[x]]$-module of Kähler differentials on $\Gamma_{\circ}$ of $1$-forms preserving $E$ is simple if $\Gamma_{\circ}$ is smooth and transverse to $x=0$. Indeed, it is generated by $dx$, $dy$ or $xdy$ depending on whether $\{y=0\}\not\subset E$, $E=\{y=0\}$ or $E=\{xy=0\}$ respectively. Thus, we will assume from now on $n>1$.
We can associate Puiseux characteristic exponents $\beta_{1}<\ldots<\beta_{g}$ to the pair $(\Gamma_{\circ},E)$ by considering the natural numbers $\beta$ such that
$\gcd(\{n\}\cup\{\tilde{\beta}:a_{\tilde{\beta},\Gamma_{\circ}}\neq 0\ \wedge\ \tilde{\beta}<\beta\})\neq\gcd(\{n\}\cup\{\tilde{\beta}:a_{\tilde{\beta},\Gamma_{\circ}}\neq 0\ \wedge\ \tilde{\beta}\leq\beta\}).$
Indeed if $x=0$ is not the tangent cone of $\Gamma_{\circ}$ these numbers are the usual Puiseux characteristic exponents of $\Gamma_{\circ}$. Notice that either $\beta_{0}\in(n)$ or $\beta_{0}=\beta_{1}$ that in particular $\beta_{0}=\beta_{1}$ holds if $\beta_{0}<n$. Moreover, the Puiseux exponents of $(\Gamma_{\circ},E)$ just depend on the pair $(\Gamma_{\circ},n)$.
Definition 2.2.
We define the set of exponents ${\mathcal{E}}_{E}(\Gamma_{\circ})$ of $(\Gamma_{\circ},E)$ as
${\mathcal{E}}_{E}(\Gamma_{\circ})=\cup_{j=0}^{g}\{m\in(n,\beta_{0},\beta_{1},\ldots,\beta_{j}):m\geq\beta_{j}\}.$
Given any $\Gamma\in\tilde{\mathcal{C}}$ that admits a parametrization of the form (1), we have $a_{\beta,\Gamma}=0$ if $\beta\not\in{\mathcal{E}}_{E}(\Gamma_{\circ})$. Moreover $a_{\beta_{0},\Gamma}\ldots a_{\beta_{g},\Gamma}\neq 0$ holds. In particular, the Puiseux exponents of $(\Gamma,E)$ are $\beta_{1}<\ldots<\beta_{g}$ and ${\mathcal{E}}_{E}(\Gamma)={\mathcal{E}}_{E}(\Gamma_{\circ})$. In general $E$ is fixed and we just write ${\mathcal{E}}(\Gamma)$.
We consider the subset
${\mathcal{C}}=\left\{\Gamma\ |\ \Gamma(t)=\left(t^{n},\sum_{\beta\in{\mathcal{E}}_{E}(\Gamma_{\circ})}a_{\beta,\Gamma}t^{\beta}\right)\ \mathrm{and}\ a_{\beta_{0},\Gamma}\ldots a_{\beta_{g},\Gamma}\neq 0\right\}$
of the equisingularity class $\tilde{\mathcal{C}}$ of $(\Gamma_{\circ},E)$. We say that ${\mathcal{C}}$ is an arranged equisingularity class of curves. It intersects any analytic class contained in $\tilde{\mathcal{C}}$. We just remove from $\tilde{\mathcal{C}}$ the curves $\Gamma$ such that $n\neq(\Gamma,x=0)_{(0,0)}$ or $\beta_{0}\neq(\Gamma,y=0)_{(0,0)}$. For simplicity, we drop the term “arranged” from now on.
Remark 2.1.
The above discussion allows us to define ${\mathcal{E}}_{E}({\mathcal{C}})=\mathcal{E}_{E}(\Gamma)$ for $\Gamma$ in $\mathcal{C}$, as ${\mathcal{E}}$ is constant in ${\mathcal{C}}$. This way, we can consider ${\{a_{\beta}\}}_{\beta\in{\mathcal{E}}({\mathcal{C}})}$ as variables that parametrize the equisingularity class ${\mathcal{C}}$. These parameters are the main tool allowing us to attain our purpose. Thus, from now on, and unless explicitly said otherwise, we forget the subindex 0 and will work with $\Gamma$ generic in its equisingularity class.
2.2. The semigroup and the semi-module
Recall that the main topological invariant associated to $\Gamma$ is its semigroup:
${\mathcal{S}}(\Gamma)=\left\{\nu_{\Gamma}(f):=\operatorname{ord}_{t}(f(\Gamma(t)))\ :\ f(x,y)\in\mathbb{C}[[x,y]]\right\},$
whereas it is known since [14] that the main analytic invariant of $\Gamma$ is the set $\Lambda_{\Gamma}$ of contacts of holomorphic (or formal) $1$-forms with $\Gamma$.
Definition 2.3.
We define
$\Lambda_{\Gamma}^{E}=\big{\{}\nu_{\Gamma}(\omega):=\operatorname{ord}_{t}(\Gamma^{\ast}\omega)+1\ |\ \big{.}\big{.}\omega\in\hat{\Omega}_{E}(\mathbb{C}^{2},0)\big{\}},$
where
$\hat{\Omega}_{E}(\mathbb{C}^{2},0)=\left\{Adx+Bdy\ |\ \ A,B\in\mathbb{C}[[x,y]]\ \mathrm{and}\ E\ \mathrm{is}\ \omega-\mathrm{invariant}\right\}$
is the $\mathbb{C}[[x,y]]$-module of formal $1$-forms in $(\mathbb{C}^{2},0)$ preserving $E$. We also define ${\Omega}_{E}(\mathbb{C}^{2},0)$ as the subset of $\hat{\Omega}_{E}(\mathbb{C}^{2},0)$ consisting of germs of analytic $1$-forms in $(\mathbb{C}^{2},0)$, and denote
$\hat{\Omega}(\mathbb{C}^{2},0)=\hat{\Omega}_{\emptyset}(\mathbb{C}^{2},0),\ {\Omega}(\mathbb{C}^{2},0)={\Omega}_{\emptyset}(\mathbb{C}^{2},0)\ \mathrm{and}\ \Lambda_{\Gamma}=\Lambda_{\Gamma}^{\emptyset}.$
We intend to compute $\Lambda_{\Gamma}^{E}$ for generic $(\Gamma,E)$ in a given equisingularity class ${\mathcal{C}}$: what we call $\Lambda_{\mathcal{C}}^{E}$.
The set $\Lambda_{\Gamma}^{E}$ is not a semigroup, but it is a so called semimodule: if $\ell\in\Lambda_{\Gamma}^{E}$ and $s\in\mathcal{S}(\Gamma)$, then obviously $\ell+s\in\Lambda_{\Gamma}^{E}$ (because $f\omega$ belongs to $\hat{\Omega}_{E}(\mathbb{C}^{2},0)$ if $\omega$ does). However, we do not need such generality. We shall only use the fact that if $\ell\in\Lambda_{\Gamma}^{E}$, then $n+\ell\in\Lambda_{\Gamma}^{E}$. If we write $[\ell]_{m}$ for the class modulo $m$ of $\ell$, then: if $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ is a family of $1$-forms such that $\nu_{\Gamma}(\Omega_{i})$ is minimum in $[\nu_{\Gamma}(\Omega_{i})]_{n}\cap\Lambda_{\Gamma}^{E}$ for $i=1,\ldots,n$ and $[\nu_{\Gamma}(\Omega_{i})]_{n}\neq[\nu_{\Gamma}(\Omega_{j})]_{n}$ for $i\neq j$, then certainly $\mathbf{\Omega}$ gives, by push-back to $\Gamma$, a basis of Kähler $1$-forms over $\Gamma$ and preserving $E$, i.e.
$\Gamma^{*}\hat{\Omega}_{E}(\mathbb{C}^{2},0)=\Gamma^{*}({\mathbb{C}}[[x]]\Omega_{1}+\ldots+{\mathbb{C}}[[x]]\Omega_{n}).$
This will be, in essence, the aim of our construction: providing a simultaneous construction for bases of Kähler differentials on generic curves of a fixed equisingularity class ${\mathcal{C}}$ of a pair $(\Gamma_{\circ},E)$.
2.3. Combinatorial invariants of $(\Gamma,E)$
The following are the basic invariants of $(\Gamma,E)$ (or, equivalently, its equisingularity class ${\mathcal{C}}$), following [19].
Definition 2.4.
We define the sequences $(e_{j})_{j\geq 1}$, $(n_{j})_{j\geq 1}$ and $(\nu_{j})_{j\geq 1}$ as
  1. (1)
    Let $e_{0}=n$ and $e_{m}=\gcd(e_{m-1},\beta_{m})$ if $1\leq m\leq g$.
  2. (2)
    Then $n_{0}=1$ and $n_{m}=e_{m-1}/e_{m}$ if $1\leq m\leq g$.
  3. (3)
    Finally, $\nu_{m}=n_{0}\ldots n_{m}=n/e_{m}$ for $0\leq m\leq g$.
2.4. Approximate roots of $(\Gamma,E)$
In order to provide bases of Kähler differentials for generic curves in $\mathcal{C}$, we shall need to use both the elementary $1$-forms $dx$ and $dy$ and the approximate roots of $\Gamma$ where $\Gamma\in{\mathcal{C}}$.
Fix $11\leq j\leq g+$ 1. We define
$f_{j}=\prod_{k=0}^{\nu_{j-1}-1}(y-\sum_{\beta<\beta_{j}}e^{\frac{2\pi ik\beta}{n}}a_{\beta}x^{\beta/n}).$
Since $f_{j}$ is invariant by the transformation $x\mapsto e^{\frac{2\pi i}{n}}x$, it belongs to $\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x]][y]$. Given $\Gamma\in\mathcal{C}$, let $\Gamma_{<\beta_{j}}$ be the truncation
$\Gamma_{<\beta_{j}}=\bigg{(}t^{n},\sum_{\beta<\beta_{j}}a_{\beta}t^{\beta}\bigg{)}$
of $\Gamma$ up to (but not including) $\beta_{j}$. By convention, we denote $\beta_{g+1}=\infty$ and we obtain $\Gamma_{<\beta_{g+1}}=\Gamma$. By construction, $f_{j,\Gamma}\circ\Gamma_{<\beta_{j}}\equiv 0$ holds for all $\Gamma\in\mathcal{C}$ and $11\leq j\leq g+$ 1.
Remark 2.2.
When $n=\nu(\Gamma)$, the polynomials $f_{1,\Gamma},\ldots,f_{g,\Gamma}$ are called, since [1], the approximate roots of $\Gamma$. In our context, we call them the approximate roots of the pair $(\Gamma,E)$. The value $\overline{\beta}_{j}=\nu_{\Gamma}(f_{j,\Gamma})$ (i.e. $\overline{\beta}_{j}=\operatorname{ord}_{t}(f_{j,\Gamma}(\Gamma))$) is a topological invariant of $\Gamma$ and thus it does not depend on $\Gamma\in\mathcal{C}$ for $1\leq j\leq g$. It is known that ${\mathcal{S}}(\Gamma)$ is the semigroup $\langle n,\overline{\beta}_{1},\ldots,\overline{\beta}_{g}\rangle$ generated by $n,\overline{\beta}_{1},\ldots,\overline{\beta}_{g}$; moreover, we have $\overline{\beta}_{1}=\beta_{1}$ and
(2) $\overline{\beta}_{j}=n_{j-1}\overline{\beta}_{j-1}-\beta_{j-1}+\beta_{j}$
for $2\leq j\leq g$ (cf. [19, Chapter II, Theorem 3.9]). Notice that $\overline{\beta}_{j}$ is the first element of the semigroup of $\Gamma$ that is not a multiple of $e_{j-1}$. Moreover, we have $\overline{\beta}_{j}-{\beta}_{j}\in(e_{j-1})$.
Remark 2.3.
The $n$ expressions
$k_{1}\overline{\beta}_{1}+\ldots+k_{g}\overline{\beta}_{g}$
with $0\leq k_{1}<n_{1}$, $\ldots$, $0\leq k_{g}<n_{g}$ define $n$ different classes modulo $n$. The analogous property holds if we replace $\overline{\beta}_{j}$ with $\beta_{j}$ for $1\leq j\leq g$.
Thus, using $f_{0,\Gamma},\ldots,f_{g,\Gamma}$, where $f_{0,\Gamma}=x$, we can compute the whole semigroup $\mathcal{S}(\Gamma)$ as the set
(3) $\{\nu_{\Gamma}(f_{0,\Gamma}^{k_{0}}f_{1,\Gamma}^{k_{1}}\cdots f_{g,\Gamma}^{k_{g}}):0\leq k_{0},k_{1},\ldots,k_{g}\}.$
Paralleling the structure of the semigroup $\mathcal{S}(\Gamma)$, which is generated by the $\Gamma$-orders of the $g$ approximate roots of $\Gamma$, we shall divide the construction of a basis of Kähler $1$-forms $\mathbf{\Omega}$ on $\Gamma$ into $g$ levels, one for each Puiseux characteristic exponent, each level $l$ giving rise to a family $\mathcal{T}_{l}$ of $1$-forms, so that $\mathcal{T}_{l}\subset\mathcal{T}_{l+1}$ and $\mathbf{\Omega}\subset\mathcal{T}_{g}$. The value $2{\nu}_{l}$ will be the cardinal of the set of $1$-forms of each level $l$ in $\mathbf{\Omega}$.
2.5. Parametrized $1$-forms
As we are going to use the approximate roots, which are of the form $y^{j}+\cdots$, where the dots have exponents in $y$ of order less than $j$, we will need to study a distinguished family of $\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x]]$-modules.
Definition 2.5.
Given $D\in{{\mathbb{C}}[a_{\beta}]}_{\beta\in{\mathcal{E}}({\mathcal{C}})}$, we denote by $D(\Gamma)$ the value $D({(a_{\beta})}_{\beta\in{\mathcal{E}}({\mathcal{C}})})$. Given $g=\sum D_{r}t^{r}$, where $D_{r}\in{{\mathbb{C}}[a_{\beta}]}_{\beta\in{\mathcal{E}}({\mathcal{C}})}$ for any $r$, we define $g_{\Gamma}=\sum D_{r}(\Gamma)t^{r}$.
Definition 2.6.
We set:
$\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)=\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]dx+\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]dy.$
Given $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ and $\Gamma\in\mathcal{C}$ corresponding to the family ${(a_{\beta})}_{\beta\in\mathcal{E}(\mathcal{C})}$ of complex coefficients, we define the form $\Omega_{\Gamma}\in\hat{\Omega}\mbox{(${\mathbb{C}}^{2},0$)}$ obtained by evaluating $\Omega$ in ${(a_{\beta,\Gamma})}_{\beta\in\mathcal{E}(\mathcal{C})}$ analogously as in Definition 2.5. We consider the set ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ consisting of $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ such that $\Omega_{\Gamma}\in{\Omega}\mbox{(${\mathbb{C}}^{2},0$)}$ for any $\Gamma\in\mathcal{C}$, and denote $\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)=\Omega_{E}(\mathbb{C}^{2},0)\cap\Omega_{\mathcal{C}}(\mathbb{C}^{2},0)$.
So, we can consider the elements of $\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ as formal $1$-forms parametrized by ${\mathcal{C}}$. We focus on the properties of $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ for generic $\Gamma\in\mathcal{C}$.
Definition 2.7.
Given $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$, define
$\nu(\Omega)=\nu(\Omega_{\Gamma}),\ \ \nu_{\mathcal{C}}(\Omega)=\nu(t\Gamma^{*}\Omega_{\Gamma})\ \ \mathrm{and}\ \ \Lambda_{\mathcal{C}}^{E}=\Lambda_{\Gamma}^{E}$
for a generic $\Gamma\in\mathcal{C}$.
The genericity condition will become clearer when constructing the families $\mathcal{T}_{l}$.
3. Case $E=\emptyset$
We divide our study into two cases: $E=\emptyset$, $E\neq\emptyset$. The former will in fact contain all the important definitions and arguments, whereas the latter will just consist in a generalization in which everything except the bootstrapping process carries over without modification.
The case $E=\emptyset$ is the foundational one. For the sake of simplicity, and because we are essentially doing the ground work, we omit $E$ everywhere in this section, as it is irrelevant.
3.1. Distinguished modules of $1$-forms
We shall make frequent use of the following $1$-forms and modules.
Definition 3.1.
We define
${\mathcal{M}}_{2\nu_{l}}=\left(\bigoplus_{j=0}^{\nu_{l}-1}\mathbb{C}[[x]]y^{j}dx\right)\bigoplus\left(\bigoplus_{j=0}^{\nu_{l}-1}\mathbb{C}[[x]]y^{j}dy\right)$
for $0\leq l\leq g$.
Let
$\overline{\Omega}_{2j+1}=y^{j}dx\quad\mathrm{and}\quad\overline{\Omega}_{2j+2}=y^{j}dy$
for $j\geq 0$. Thus, $(\overline{\Omega}_{1},\ldots,\overline{\Omega}_{2\nu_{l}})$ is a basis of the free ${\mathbb{C}}[[x]]$-module ${\mathcal{M}}_{2\nu_{l}}$.
Definition 3.2.
We define ${\mathcal{I}}_{\Gamma}$ as the set of $1$-forms $\Omega\in\hat{\Omega}(\mathbb{C}^{2},0)$ such that $\Gamma^{*}\Omega\equiv 0$.
Now, let us focus on parametrized forms. Setting $M_{0}=\{0\}$, we define, recursively,
$M_{j+1}=\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x]]\overline{\Omega}_{j+1}+M_{j},\ \ \hat{M}_{j+1}=(\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}\setminus\{0\})\overline{\Omega}_{j+1}+M_{j},$
for $j\geq 0$. Clearly $\hat{M}_{j}\subset M_{j}$ for $j\geq 1$.
The subspaces $M_{j}$ are relevant for our computations because they satisfy the following important properties:
Lemma 3.1.
Let $P=\sum_{j=0}^{k}\alpha_{j}y^{j}\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x]][y]$ where $\alpha_{j}\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x]]$ for any $0\leq j\leq k$. Then:
$PM_{j}\subset M_{j+2k}\ \ \mathrm{for\ any}\ j\geq 1.$
Moreover, if $\alpha_{k}\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}\setminus\{0\}$, we also have
$P\hat{M}_{j}\subset\hat{M}_{j+2k}\ \ \mathrm{for\ any}\ j\geq 1.$
As a matter of fact, the most important spaces are the ones corresponding to the indices $2{\nu}_{j}$:
Definition 3.3.
A $1$-form $\omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ is of level $l$ if $\omega\in M_{2{\nu}_{l}}$ for some $l\in\{0,\ldots,g\}$.
All the $1$-forms we shall construct will be of some level $l$, as the reader will see. Certainly, $dx$ and $dy$ are both of level 00.
3.2. Ordered families
Our aim is to find nice bases for generic curves $\Gamma$ in a given equisingularity class ${\mathcal{C}}$.
Definition 3.4.
Consider a family $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{r})$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ where $0\leq l\leq g$ and $r\in\{2\nu_{l},\min(2\nu_{l},n)\}$. We say that $\mathbf{\Omega}$ is an ordered family of level $l$ for $\mathcal{C}$ if
  1. (i)
    $\Omega_{j}\in\hat{M}_{j}$ for any $1\leq j\leq r$;
  2. (ii)
    $\nu_{\mathcal{C}}(\Omega_{j})\leq\nu_{\mathcal{C}}(\Omega_{k})$ for all $1\leq j\leq k\leq r$.
We say that $\mathbf{\Omega}$ is a separated family of level $l$ if it is ordered and
  1. (iii)
    $\nu_{\mathcal{C}}(\Omega_{1}),\ldots,\nu_{\mathcal{C}}(\Omega_{\min(2\nu_{l},n)})$ define pairwise different classes modulo $n$.
For $l=g$ and $r=n$, we say that $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$, with $\Omega_{i}\in{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$, is a nice basis if it is a separated family and
  1. (iv)
    $\hat{\Omega}(\mathbb{C}^{2},0)=\mathbb{C}[[x]]\Omega_{1,\Gamma}\oplus\ldots\oplus\mathbb{C}[[x]]\Omega_{n,\Gamma}\oplus{\mathcal{I}}_{\Gamma}$ for generic $\Gamma\in\mathcal{C}$.
Remark 3.1.
Condition (i) implies
(4) ${\mathcal{M}}_{2\nu_{l}}=\mathbb{C}[[x]]\Omega_{1,\Gamma}+\ldots+\mathbb{C}[[x]]\Omega_{2\nu_{l},\Gamma}$
for generic $\Gamma\in\mathcal{C}$ if $r=2\nu_{l}$. Thus, ${\bf\Omega}$ is a basis of the ${\mathbb{C}}[[x]]$-module ${\mathcal{M}}_{2\nu_{l}}$. Moreover, it guarantees that the minimality property (cf. section 1) is satisfied for generic $\Gamma\in\mathcal{C}$. A nice basis ${\bf\Omega}$ induces an Apery basis of $\Gamma$ (cf. section 1) for generic $\Gamma\in\mathcal{C}$ by Conditions (iii) and (iv). Finally, a nice basis satisfies the monotonicity condition (cf. section 1) for generic $\Gamma\in\mathcal{C}$ by Conditions (ii) and (iii).
At level 00, the first condition will be accomplished by considering the 1-forms $dx$ and $df_{1}$, that constitute a separated family. Then we describe a process in which $dx$ and $df_{1}$ appear alternatively and the approximate roots of $\Gamma$, with $\Gamma\in\mathcal{C}$, play a prominent role. We obtain $f_{1}=y$, since $\beta_{0}=\beta_{1}$ for the case $E=\emptyset$. Anyway, we have $f_{1}\neq y$ if $\beta_{0}<\beta_{1}$. So, in order to have uniform notations for the cases $E=\emptyset$ and $E\neq\emptyset$, we prefer to use $f_{1}$ in all cases.
3.3. The basic step — leading variables
Fix an equisingularity class ${\mathcal{C}}$ of germs of branches. Our construction essentially follows this process: assume we have found a terminal family (Definition 3.10) $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ of level $l<g$. Such family is separated but we do not delve now into its properties, which will be introduced in their definition. For now, a terminal family of level $l$ can be considered just as the product of our algorithm at the end of level $l$. We define $\mathcal{T}_{0}=\{dx,df_{1}\}$ as the terminal family of level 00. We are going to construct, iteratively, a terminal family $\mathcal{T}_{l+1}$ of level $l+1$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$, with $2{\nu}_{l+1}$ elements. In the end, we shall have $\mathcal{T}_{g}$ with $2n$ elements whose first $n$ elements constitute a nice basis for $\mathcal{C}$. The construction of $\mathcal{T}_{l+1}$ requires several stages. The first one (stage 00) is elementary: we combine the $1$-forms in $\mathcal{T}_{l}$ with the approximate root $f_{l+1}$. More precisely, we define
(5) $\Omega_{2a\nu_{l}+j}^{0}=f_{l+1}^{a}\Omega_{j}$
for all $0\leq a<n_{l+1}$ and $1\leq j\leq 2\nu_{l}$ (recall that $f_{l+1,\Gamma}$ is the $l+1$-th approximate root of $\Gamma$). Notice the aforementioned alternate role of $dx$ and $df_{1}$ since $\Omega_{k}^{0}\in\hat{M}_{k}$ for $k\geq 1$.
Definition 3.5.
Fix a terminal family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ We say that
[…]
is a the initial family (of stage 00) of level $l+1$.
These $1$-forms belong to ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$, see equation (14), and are $2{\nu}_{l+1}$ in number. They will be called the stage 0 forms of level $l+1$.
Remark 3.2.
Since $\nu_{\mathcal{C}}(\Omega_{2a\nu_{l}+j}^{0})=a\overline{\beta}_{l+1}+\nu_{\mathcal{C}}(\Omega_{j})$, the $\mathcal{C}$-orders of elements of $\mathcal{T}_{l}$ determine the $\mathcal{C}$-orders of the elements of $\mathcal{G}_{l+1}^{0}$.
Definition 3.6.
Given $l\in{\mathbb{Z}}$ and $m\in{\mathbb{Z}}_{\geq 1}$, we denote by $[l]_{m}$ the class of $l$ modulo $m$. Moreover, given $S\subset{\mathbb{Z}}$, we define $[S]_{m}=\{[l]_{m}\ |\ \ l\in S\}$.
Unfortunately, the stage 0 forms of level $l+1$ are ordered but usually not separated (condition (iii) in Definition 3.4). We will construct families $\mathcal{G}_{l+1}^{0},\mathcal{G}_{l+1}^{1},\ldots$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ until we find a (separated) family $\mathcal{G}_{l+1}^{r}$, which will be the terminal one.
Definition 3.7.
Assume a family $\mathcal{G}_{l+1}^{s}=(\Omega_{1}^{s},\ldots,\Omega_{2{\nu}_{l+1}}^{s})$ is given, with $0\leq l<g$ and $s\geq 0$ (i.e. a family of level $l+1$ at some stage). Assume that $\mathcal{G}_{l+1}^{s}$ is non-separated. When this happens, we consider the pairs $j,k$ with $1\leq j<k\leq 2{\nu}_{l+1}$ such that $[\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$: whenever this equality holds, there is some $d\in\mathbb{Z}_{\geq 0}$ with $\nu_{\mathcal{C}}(\Omega^{s}_{k})=\nu_{\mathcal{C}}(\Omega_{j}^{s})+dn$. Consider the pull-backs
$\Gamma^{\ast}\Omega_{k,\Gamma}^{s}=(C_{k,\eta_{k}}(\Gamma)t^{\eta_{k}}+\cdots)dt,\ \Gamma^{\ast}\Omega^{s}_{j,\Gamma}=(C_{j,\eta_{j}}(\Gamma)t^{\eta_{j}}+\cdots)dt$
of $\Omega_{k}^{s}$ and $\Omega_{j}^{s}$, where $C_{k,\eta_{k}},C_{j,\eta_{j}}\in{{\mathbb{C}}[a_{\beta}]}_{\beta\in{\mathcal{E}}({\mathcal{C}})}\setminus\{0\}$, for all curves in an equisingularity class. For any such a class, $C_{k,\eta_{k}}$ and $C_{j,\eta_{j}}$ are both generically non-vanishing. Then we replace $\Omega^{s}_{k}$ with
(6) $\Omega^{s+1}_{k}:=C_{j,\eta_{j}}\Omega^{s}_{k}-C_{k,\eta_{k}}x^{d}\Omega^{s}_{j};$
Setting also $\Omega^{s+1}_{j}=\Omega^{s}_{j}$, for the other $1$-form, we obtain the $2{\nu}_{l+1}$ $1$-forms in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ of stage $s+1$ (hence the exponent in $\Omega^{s+1}_{j}$), in the same level $l+1$.
Remark 3.3.
The families $\mathcal{G}_{l+1}^{1},\mathcal{G}_{l+1}^{2},\ldots$ are well-defined (Proposition 3.2). There are some points in the definition that need to be proved, for instance that $d$ is non-negative. Arithmetic arguments will show how this construction can be iterated and at some stage it gives a separated family $\mathcal{G}_{l+1}^{r}$ of $1$-forms of level $l+1$.
Remark 3.4.
We have $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})>\nu_{\mathcal{C}}(\Omega^{s}_{k})$ and $[\nu_{\mathcal{C}}(\Omega^{s+1}_{k})]_{n}\neq[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ as we shall see (Propositions 3.3 and 3.4).
Remark 3.5.
The algorithm described above to obtain a separated family $\mathcal{G}_{l+1}^{r}$ was called simple in section 1. One reason is that $\mathcal{G}_{l+1}^{0}$ is built from $\mathcal{T}_{l}$ and data related to the equisingularity class, namely the approximate root $f_{l+1}$. Another reason is that the algorithm consists in multiple applications of the elementary substitution (6). Indeed, the simplicity and expressiveness of Equation (6) will be essential to control the properties of our construction.
In order to control $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})$ in substitution (6), we require the following definition, which will allow us to explicitly compute that contact order. For the sake of simplicity, if $\beta\in\mathcal{E}({\mathcal{C}})$, then ${\mathfrak{p}}(\beta)$ is the previous element in $\mathcal{E}({\mathcal{C}})$ and ${\mathfrak{n}}(\beta)$ is the following one.
Definition 3.8.
We say that $g(t)\in\mathbb{C}[a_{\beta_{0}},a_{{\mathfrak{n}}(\beta_{0})},a_{{\mathfrak{n}}^{2}(\beta_{0})},\ldots][[t]]$ has leading variable $\beta$ (or $a_{\beta}$) of degree $m\in\mathbb{Z}_{\geq 0}$, or equivalently, that $g(t)$ starts at $\beta$ with degree $m$ if $\beta\in\mathcal{E}({\mathcal{C}})$ and
  • There exists $k\in\mathbb{Z}_{\geq 0}$ with
    $g(t)=t^{k}\sum_{r=0}^{\infty}C_{{\mathfrak{n}}^{r}(\beta)}t^{{\mathfrak{n}}^{r}(\beta)}$
    where $C_{{\mathfrak{n}}^{r}(\beta)}\in{{\mathbb{C}}[a_{\beta^{\prime}}]}_{\beta^{\prime}\in{\mathcal{E}}({\mathcal{C}})}$ for $r\geq 0$;
  • There exist $m\in\mathbb{Z}_{\geq 0}$, a polynomial $R\in\mathbb{C}{[a_{\beta^{\prime}}]}_{\beta_{0}\leq\beta^{\prime}<\beta}\setminus\{0\}$, a unit $\tilde{c}\in{\mathbb{C}}^{*}$ and a sequence $c_{\beta}=1,c_{{\mathfrak{n}}(\beta)},c_{{\mathfrak{n}}^{2}(\beta)},\ldots$ in ${\mathbb{C}}$ such that
    (7) $C_{{\mathfrak{n}}^{r}(\beta)}=\tilde{c}c_{{\mathfrak{n}}^{r}(\beta)}a_{{\mathfrak{n}}^{r}(\beta)}R(a_{\beta_{0}},\ldots,a_{{\mathfrak{p}}(\beta)})a_{\beta}^{m}+Q_{{\mathfrak{n}}^{r}(\beta)},$
    for all $r\geq 0$, where $Q_{{\mathfrak{n}}^{r}(\beta)}\in\mathbb{C}{[a_{\beta^{\prime}}]}_{\beta_{0}\leq\beta^{\prime}<{\mathfrak{n}}^{r}(\beta)}$.
  • If $\beta\not\in\left\{\beta_{1},\ldots,\beta_{g}\right\}$ (i.e. it is not a Puiseux characteristic exponent) and $\ell=\max\{l\geq 0:\beta_{l}\leq\beta\}$, then $m=0$ and $[k]_{e_{\ell}}=[0]_{e_{\ell}}$.
  • If $\beta=\beta_{\ell+1}$ (i.e. it is a Puiseux characteristic exponent), then $[k]_{e_{\ell}}=[m\beta]_{e_{\ell}}$.
The sequence $(c_{{\mathfrak{n}}^{r}(\beta)})_{r\geq 0}$ will be called the $c$-sequence of $g$.
Notice that not all power series have a leading variable.
Remark 3.6.
Consider the equisingularity class whose arranged representatives are of the form $\Gamma(t)=(t^{6},a_{9}t^{9}+a_{10}t^{10}+\ldots)$ where $a_{9}a_{10}\neq 0$. Its Puiseux characteristic exponents are $\beta_{1}=9$ and $\beta_{2}=10$. Let $f_{1}=y$ and
$f_{2}=(y-a_{9}x^{3/2})(y+a_{9}x^{3/2})=y^{2}-a_{9}^{2}x^{3}$
be its approximate roots. Moreover, we have that
$t\frac{\partial(f_{1,\Gamma}\circ\Gamma(t))}{\partial t}=9a_{9,\Gamma}t^{9}+10a_{10,\Gamma}t^{10}+11a_{11,\Gamma}t^{11}+\ldots$
and
$f_{2}^{2}\circ\Gamma(t)={\left(a_{10}t^{10}+a_{11}t^{11}+\ldots\right)}^{2}{\left(2a_{9}t^{9}+a_{10}t^{10}+a_{11}t^{11}+\ldots\right)}^{2}$
belong to $\mathbb{C}[a_{9},a_{10},\ldots][[t]]$, where we skip the subindexes $\Gamma$ in the second expression. In the former case, $t\frac{\partial(f_{1}\circ\Gamma(t))}{\partial t}$ has leading variable $a_{9}$ of degree 00 and we have $k=0$, $e_{0}=6$, $\tilde{c}=9$, $R\equiv 1$, $Q_{9}\equiv Q_{10}\equiv\ldots\equiv 0$, $c_{9}=1$, $c_{10}=10/9$, $c_{11}=11/9$ and so on.
In the latter case, one can verify that $f_{2}^{2}\circ\Gamma(t)$ has leading variable $a_{10}$ of degree $1$. More precisely, we have $k=28$, $e_{1}=3$, $\tilde{c}=1$, $R=4a_{9}^{2}$, $Q_{10}\equiv 0$, $Q_{11}=4a_{9}a_{10}^{3}$, $c_{10}=1$, $c_{11}=c_{12}=\ldots=2$.
Remark 3.7.
The nonzero polynomial $R(a_{\beta_{0}},\ldots,a_{{\mathfrak{p}}(\beta)})$ is the same for all the coefficients $C_{\beta}$ (this is what makes the definition of the $c-$sequence possible). Also, despite saying that $a_{\beta}$ is of “degree $m$”, the first element $C_{\beta}$ is
$C_{\beta}=\tilde{c}a_{\beta}R(a_{\beta_{0}},\ldots,a_{{\mathfrak{p}}(\beta)})a_{\beta}^{m}+Q_{\beta}.$
But the terminology reflects the fact that $a_{\beta}$ is the maximal variable in the coefficient of the least order (the “leader”), and if $m>0$, then it subsequently appears explicitly always to the power $m$ in $C_{{\mathfrak{n}}^{r}(\beta)}$, for any $r>0$. If $\beta$ is not a Puiseux characteristic exponent, then $m=0$.
As in Remark 3.6, the notion of leading variable can be extended to power series $f\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$ just by considering $g(t)=f\circ\Gamma(t)$, where $\Gamma\in\mathcal{C}$, since $g$ can be interpreted as an element of $\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[t]]$. Analogously, it can be extended to $1$-forms $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ by considering the power series $g(t)$, where $\Gamma\in\mathcal{C}$, satisfying $t\Gamma^{*}\Omega_{\Gamma}=g(t)dt$.
Remark 3.8.
The use of leading variables will enable us to control $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})$ in substitution (6). As a matter of fact, we will prove, in Propositions 3.3 and 3.4 that
$\nu_{\mathcal{C}}(\Omega_{k}^{s+1})=\nu_{\mathcal{C}}(\Omega_{k}^{s})+({\mathfrak{n}}^{s+1}(\beta_{l+1})-{\mathfrak{n}}^{s}(\beta_{l+1})).$
This property, together with Remark 3.2, determine the $\mathcal{C}$-orders of all the $1$-forms appearing in the construction of a nice basis of $\mathcal{C}$. We obtain an algorithm to calculate such $\mathcal{C}$-orders without calculating the corresponding $1$-forms. This is the property of rich structure of the construction mentioned in section 1.
Definition 3.9.
The set of power series $g(t)$ starting at $\beta$ with degree $m$ will be denoted $\Delta^{m}_{\beta}$. We shall also make use of
  • The set $\Delta_{\beta,\leq}^{m}$ of power series $g(t)\in\Delta^{m}_{\beta}$ whose $c$-sequence is increasing in ${\mathbb{Q}}^{+}$;
  • The set $\Delta^{m}_{\beta,<}$ of power series $g(t)\in\Delta^{m}_{\beta}$ whose $c$-sequence is strictly increasing in ${\mathbb{Q}}^{+}$.
Given $1\leq j\leq g$, we define the set $\tilde{\Delta}_{\beta_{j}}$ consisting of $g\in\Delta_{\beta_{j}}$ such that $Q_{\beta_{j}}\equiv 0$. We denote $\tilde{\Delta}_{\beta_{j},\leq}^{m}={\Delta}_{\beta_{j},\leq}^{m}\cap\tilde{\Delta}_{\beta_{j}}$ and $\tilde{\Delta}_{\beta_{j},<}^{m}={\Delta}_{\beta_{j},<}^{m}\cap\tilde{\Delta}_{\beta_{j}}$. For each $\beta\in\mathcal{E}(\mathcal{C})$, the union for all degrees $m$ of each of those sets is also important. Thus, we denote
$\Delta_{\beta}=\cup_{m\geq 0}\Delta_{\beta}^{m},\ \Delta_{\beta,\leq}=\cup_{m\geq 0}\Delta_{\beta,\leq}^{m}\ \mathrm{and}\ \Delta_{\beta,<}=\cup_{m\geq 0}\Delta_{\beta,<}^{m}.$
We also denote $\tilde{\Delta}_{\beta_{j},\leq}=\cup_{m\geq 0}\tilde{\Delta}_{\beta_{j},\leq}^{m}$ and $\tilde{\Delta}_{\beta_{j},<}=\cup_{m\geq 0}\tilde{\Delta}_{\beta_{j},<}^{m}$.
Remark 3.9.
Clearly, $\Delta_{\beta}\cap\Delta_{\beta^{\prime}}=\emptyset$ if $\beta\neq\beta^{\prime}$ so that the leading variable of $g\in\cup\Delta_{\beta}$ is unique.
Remark 3.10.
The calculations in Remark 3.6 imply $df_{1}\in\Delta_{9,<}^{0}$ and $f_{2}^{2}\in\Delta_{10,\leq}^{1}\setminus\Delta_{10,<}^{1}$.
We shall also use the same notation $\Delta^{m}_{\beta}$, etc. for the corresponding subsets of $\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$ and $\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ because there is no possibility of confusion. For completeness, we add $dx$ to the set $\Delta_{n,<}$ in the case of $1$-forms.
The next result will be critical to deal with $1$-forms whose leading variable is a Puiseux characteristic exponent.
Proposition 3.1.
Assume $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ (resp. $f\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$) belongs to $\Delta_{\beta_{l},\leq}^{m}$ for some $1\leq l\leq g$ and $20\leq m\leq n_{l}-$ 2. Then $\Omega\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$ (resp. $f\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$).
Proof.
With the notations of Definition 3.8, consider $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$. The proof for $f\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$ is analogous but it can also be deduced from the case of $1$-forms by considering $df=\frac{\partial f}{\partial x}dx+\frac{\partial f}{\partial y}dy$.
By definition, we know that $[k]_{e_{l-1}}=[m\beta_{l}]_{e_{l-1}}$. Let $\Gamma$ be a generic curve with irreducible parametrization
$\Gamma(t)=\left(t^{n/e_{l-1}},\sum_{\beta_{0}\leq\beta<\beta_{l}}a_{\beta,\Gamma}t^{\beta/e_{l-1}}\right).$
Denote $\Gamma^{*}\Omega=h(t)dt/t$. Notice that, using Landau’s notation,
$e_{l-1}h(t^{e_{l-1}})\frac{dt}{t}=(\Gamma(t^{e_{l-1}}))^{*}\Omega=\left(C_{\beta_{l}}t^{k+\beta_{l}}+O(t^{k+\beta_{l}+1})\right)\frac{dt}{t}.$
Assume, aiming at contradiction, that $Q_{\beta_{l}}\not\equiv 0$. This implies $C_{\beta_{l}}\neq 0$ for a generic curve as above. As the left hand side has order a multiple of $e_{l-1}$, we obtain $[0]_{e_{l-1}}=[k+\beta_{l}]_{e_{l-1}}$ and thus $(m+1)\beta_{l}$ is a multiple of $e_{l-1}$, which is impossible because $m+1$ is not a multiple of $n_{l}$. ∎
3.4. Leading variables of products
Because our argument hinges on the substitutions (6) and their iterations, we need to study how some elementary operations affect the leading variables of power series and $1$-forms. This is the aim of the following technical lemmas.
We first prove that product of several (say $a$) power series starting at a Puiseux characteristic exponent $\beta_{l}$ with degree 00 and $b$ power series starting at $\beta<\beta_{l}$ gives a power series with leading variable $\beta_{l}$ of degree $a-1$. We can also control the $c$-sequence of such product.
Lemma 3.2.
Fix $1\leq l\leq g$. Suppose
$g_{1}(t),\ldots,g_{a}(t)\in\tilde{\Delta}_{\beta_{l},\leq}^{0}\ \mathrm{and}\ g_{a+1}(t),\ldots,g_{b}(t)\in\cup_{\beta^{\prime}<\beta_{l}}\Delta_{\beta^{\prime}}$
with $a\geq 1$. Then $\prod_{j=1}^{b}g_{j}(t)\in\tilde{\Delta}_{\beta_{l},\leq}^{a-1}$. Moreover, the $c$-sequence of $\prod_{j=1}^{b}g_{j}(t)$ is equal to the sum of the $c$-sequences of $g_{1},\ldots,g_{a}$ except at its first term, which is $1$.
Proof.
For each $j$, let $\beta(j)$ be such that $g_{j}\in\Delta_{\beta(j)}$ for $1\leq j\leq b$. Obviously, $\beta(j)=\beta_{l}$ for $1\leq j\leq a$. We have
(8) $g_{j}(t)=t^{k_{j}}\sum_{r=0}^{\infty}C_{{\mathfrak{n}}^{r}(\beta(j)),j}t^{{\mathfrak{n}}^{r}(\beta(j))},$
for $1\leq j\leq b$ by Definition 3.8. We need to prove, first, that $\prod_{j=1}^{b}g_{j}(t)$ is of the form
(9) $\prod_{j=1}^{b}g_{j}(t)=t^{k}\sum_{r=0}^{\infty}C_{{\mathfrak{n}}^{r}(\beta_{l})}t^{{\mathfrak{n}}^{r}(\beta_{l})}$
where $C_{{\mathfrak{n}}^{r}(\beta_{l})}\in{\mathbb{C}}[a_{\beta_{0}},a_{{\mathfrak{n}}(\beta_{0})},\ldots]$ for any $r\geq 0$. By equation (8), if the coefficient of $t^{\ell}$ in $\prod_{j=1}^{b}g_{j}(t)$ is non-vanishing, then
(10) $\ell=(k_{1}+\beta(1)+m_{1})+(k_{2}+\beta(2)+m_{2})+\ldots+(k_{b}+\beta(b)+m_{b}),$
where $\beta(j)+m_{j}\in{\mathcal{E}}({\mathcal{C}})$ and $m_{j}\in\mathbb{Z}_{\geq 0}$ for any $1\leq j\leq b$ by the first property of Definition 3.8. Notice that
$k=\nu_{0}\left(\prod_{j=1}^{b}g_{j}(t)\right)-\beta_{l}=\sum_{j=1}^{b}(k_{j}+\beta(j))-\beta_{l}\in[(a-1)\beta_{l}]_{e_{l-1}}$
holds by the first property of Definition 3.8. By subtracting $k$ from Expression (10), we deduce that in order to prove (9) we just need to show the following claim: if $\beta(j)+m_{j}\in{\mathcal{E}}({\mathcal{C}})$ for any $1\leq j\leq b$ then $\beta_{l}+\sum_{j=1}^{b}m_{j}\in{\mathcal{E}}({\mathcal{C}})$.
Let us prove the claim. Let $\beta^{\prime}=\max_{1\leq j\leq b}(\beta(j)+m_{j})$, and set
$e(\beta)=\gcd\{\beta^{\prime\prime}\in\mathcal{E}(\mathcal{C})\ :\ \beta^{\prime\prime}\leq\beta\}.$
Notice that $\beta^{\prime}+re(\beta^{\prime})\in{\mathcal{E}}(\mathcal{C})$ for all $\beta^{\prime}\in\mathcal{E}(\mathcal{C})$ and $r\in{\mathbb{Z}}_{\geq 0}$ by Definition of $\mathcal{E}(\mathcal{C})$. Moreover, $e(\beta^{\prime})$ divides $m_{1},\ldots,m_{b}$, $\beta^{\prime}-\beta_{l}$ and hence it also divides $\sum_{j=1}^{b}m_{j}-(\beta^{\prime}-\beta_{l})$. As a consequence,
$\beta_{l}+\sum_{j=1}^{b}m_{j}=\beta^{\prime}+\sum_{j=1}^{b}m_{j}-(\beta^{\prime}-\beta_{l})\in{\mathcal{E}}(\mathcal{C}),$
as desired.
Let us study the polynomials $R$ and $Q_{{\mathfrak{n}}^{r}(\beta_{l})}$, for $r\geq 0$, in Definition 3.8. Now, if the coefficient of $t^{\ell}$ in $\prod_{j=1}^{b}g_{j}(t)$ does not vanish then $\ell=k+{\mathfrak{n}}^{r}(\beta_{l})$ for some $r\geq 0$ by the above claim. Moreover, the coefficient of order $\ell$ is a sum of terms of the form
(11) $C_{\beta(1)+m_{1},1}C_{\beta(2)+m_{2},2}\ldots C_{\beta(b)+m_{b},b}$
where
(12) ${\mathfrak{n}}^{r}(\beta_{l})=\ell-k=\beta_{l}+m_{1}+\ldots+m_{b}.$
It is straightforward to verify that the expression
(13) $D_{{\mathfrak{n}}^{r}(\beta_{l})}:=\sum_{s=1}^{a}\left(C_{{\mathfrak{n}}^{r}(\beta_{l}),s}\prod_{1\leq j\leq b,\ j\neq s}C_{\beta(j),j}\right)$
is of the form (7) setting $R=\prod_{j=1}^{a}R_{j}\prod_{j=a+1}^{b}C_{\beta(j),j}$, where $R_{j}$ is the corresponding factor in equation (7) associated to $g_{j}$. By direct computation, we obtain that the sum $t^{k}\sum_{r\geq 0}D_{{\mathfrak{n}}^{r}(\beta_{l})}t^{{\mathfrak{n}}^{r}(\beta_{l})}$ belongs to $\tilde{\Delta}_{\beta_{l},\leq}$ and its $c$-sequence $(c_{{\mathfrak{n}}^{r}(\beta_{l})})_{r\geq 0}$ satisfies $c_{\beta_{l}}=1$ and $c_{{\mathfrak{n}}^{r}(\beta_{l})}=\sum_{j=1}^{a}c_{{\mathfrak{n}}^{r}(\beta_{l})}^{j}$ for any $r\geq 1$, where $(c_{{\mathfrak{n}}^{r}(\beta_{l})}^{j})_{r\geq 0}$ is the $c$-sequence of $g_{j}$ for $1\leq j\leq a$.
The terms of the form (11) that are not in (13) satisfy $\beta(j)+m_{j}<{\mathfrak{n}}^{r}(\beta_{l})$ for any $1\leq j\leq b$. Thus, the sum of such terms defines a polynomial $Q_{{\mathfrak{n}}^{r}(\beta_{l})}\in\mathbb{C}{[a_{\beta^{\prime}}]}_{\beta_{0}\leq\beta^{\prime}<{\mathfrak{n}}^{r}(\beta)}$. We obtain that $\prod_{j=1}^{b}g_{j}(t)\in{\Delta}_{\beta_{l},\leq}$ and its $c$-sequence is still $(c_{{\mathfrak{n}}^{r}(\beta_{l})})_{r\geq 0}$. Moreover, there are no such terms for $r=0$ and thus $\prod_{j=1}^{b}g_{j}(t)\in\tilde{\Delta}_{\beta_{l},\leq}$. ∎
The first important consequence applies to powers of approximate roots of $\Gamma$:
Corollary 3.1.
Let $1\leq l\leq g$ and $m\geq 1$. Then $f_{l}^{m}$ belongs to $\tilde{\Delta}_{\beta_{l},\leq}^{m-1}$ and its $c$-sequence is equal to $1,m,\ldots,m,\ldots$
Proof.
The power series $f_{l}$ is given by
(14) $f_{l}=\prod_{k=0}^{\nu_{l-1}-1}(y-\sum_{\beta<\beta_{l}}e^{\frac{2\pi ik\beta}{n}}a_{\beta}x^{\beta/n}).$
In particular $\Gamma^{\ast}f_{l}=f_{l}\circ\Gamma$ is equal to
$\Gamma^{\ast}f_{l}=g_{0}(t)g_{1}(t)\cdots g_{\nu_{l-1}-1}(t)$
where
$g_{0}(t)=\sum_{\beta\geq\beta_{l}}a_{\beta}t^{\beta}\in\tilde{\Delta}^{0}_{\beta_{l},\leq}$
which has constant $c$-sequence equal to $1$. Since
$(e^{\frac{2\pi ik\beta_{1}}{n}},e^{\frac{2\pi ik\beta_{2}}{n}},\ldots,e^{\frac{2\pi ik\beta_{l-1}}{n}})=(1,1,\ldots,1)$
implies that $k$ is a multiple of $n/\gcd(n,\beta_{1},\ldots,\beta_{l-1})=\nu_{l-1}$, we conclude that
[…]
The result is, thus, a direct consequence of Lemma 3.2, setting $a=m$ and $b=(\nu_{l-1}-1)m$, and the multinomial formula. ∎
Remark 3.11.
Remarks 3.6 and 3.10 can be obtained as an application of Corollary 3.1 where $f_{2}^{2}\in{\Delta}_{\beta_{2},\leq}^{1}={\Delta}_{10,\leq}^{1}$.
Corollary 3.1 translates to $1$-forms which are the product of a power of an approximate root of $\Gamma$ and a $1$-form with lesser leading variable:
Corollary 3.2.
Let $1\leq l\leq g$, $m\geq 1$ and $\Omega\in\Delta_{\beta}$ with $\beta<\beta_{l}$. Then $f_{l}^{m}\Omega\in\tilde{\Delta}_{\beta_{l},\leq}^{m-1}$ and the $c$-sequence of $f_{l}^{m}\Omega$ is equal to $1,m,m,\ldots$.
Finally, for products of powers of approximate roots and $1$-forms with the same leading variable with degree 00, we can also control the outcome, and the $c$-sequence. The next result is a consequence of Corollary 3.1, Lemma 3.2 and Proposition 3.1.
Corollary 3.3.
Let $1\leq l\leq g$, $m\geq 0$ and $\Omega\in{\Delta}_{\beta_{l},\leq}^{0}$. Then $f_{l}^{m}\Omega\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$ and the $c$-sequence $(c_{{\mathfrak{n}}^{r}(\beta_{l})}^{\prime})_{r\geq 0}$ of $f_{l}^{m}\Omega$ has $c_{\beta_{l}}^{\prime}=1$ and $c_{{\mathfrak{n}}^{r}(\beta_{l})}^{\prime}=c_{{\mathfrak{n}}^{r}(\beta_{l})}+m$ for $r\geq 1$ where $(c_{{\mathfrak{n}}^{r}(\beta_{l})})_{r\geq 0}$ is the $c$-sequence of $\Omega$.
All these results will be relevant for what will be stage 00 of each level $l+1$, in which we shall make use of both the $1$-forms constructed in levels up to $l$, and the approximate root $f_{l+1}$.
3.5. Terminal families
As described in subsection 3.3, given $0\leq l<g$, the initial family $\mathcal{G}_{l+1}^{0}$ of level $l+1$ is built from the terminal family $\mathcal{T}_{l}$ of level $l$ and the approximate root $f_{l+1}$. In this section we properly define terminal families, study their properties and then prove that our construction provides a terminal family $\mathcal{T}_{l+1}$ of level $l+1$.
3.5.1. Definition and first properties of terminal families
Definition 3.10.
We say that $\mathcal{T}_{l}=\{\Omega_{1},\ldots\Omega_{2{\nu}_{l}}\}\subset{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ is a terminal family of level $l$ if the following conditions hold:
  1. (a)
    $\mathcal{T}_{l}$ is a separated family of level $l$;
  2. (b)
    the last contact satisfies $\nu_{\mathcal{C}}(\Omega_{2{\nu}_{l}})\leq\overline{\beta}_{l+1}$ when $0\leq l<g$;
  3. (c)
    ${\mathcal{T}}_{l}$ is contained in $\cup_{\beta^{\prime}\in{\mathcal{E}}({\mathcal{C}})}\Delta_{\beta^{\prime},<}$. Moreover, in the case $0\leq l<g$, we have:
    $\mathcal{T}_{l}\subset\bigg{(}\bigcup_{\begin{subarray}{c}\beta^{\prime}\in{\mathcal{E}}({\mathcal{C}})\\ \beta^{\prime}<\beta_{l+1}\end{subarray}}\Delta_{\beta^{\prime},<}\bigg{)}\bigcup\Delta_{\beta_{l+1},<}^{0}$
Condition (b) indicates that the sequence of multiplicities “does not go too far” and indeed that all of them are less than or equal to the multiplicity of $df_{l+1}$ which is a form of level $l$ since it belongs to $\hat{M}_{2{\nu}_{l}}$. Condition (c) implies, on one hand, that the leading variables are not greater than $\beta_{l+1}$ for $l<g$, and that if they are $\beta_{l+1}$, then they appear with degree 00; and on the other hand, that the $c$-sequences are required to be strictly increasing. The last property is inherent to the study the semi-module of contacts of holomorphic $1$-forms; notice that Expression (3) describes the semigroup of the curves in ${\mathcal{C}}$ but $f_{0}^{k_{0}}f_{1}^{k_{1}}\cdots f_{r}^{k_{r}}\in{\Delta}_{\beta_{r},\leq}^{k_{r}-1}\setminus{\Delta}_{\beta_{r},<}^{k_{r}-1}$ if $r\geq 1$ and $k_{r}>0$ and thus those products do not have strictly increasing $c$-sequences. Condition (c) is used in the inductive steps of our construction.
The family $\mathcal{T}_{0}=\{dx,df_{1}\}$ is terminal of level 00 because $\Omega_{1}\in\Delta_{n,<}$ and $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$. This is our starting point.
Remark 3.12.
In a terminal family ${\mathcal{T}}_{l}$ of level $l<g$, the multiplicity $\nu_{\mathcal{C}}(\Omega)$ of any $1$-form $\Omega\in\mathcal{T}_{l}$ belongs to $[\beta_{l+1}]_{e_{l}}$ or $[0]_{e_{l}}$ (because of their corresponding leading variables). This is a key property for our later arguments. More precisely, we have
  • If $\Omega$ is of level $l-1$, then $\nu_{\mathcal{C}}(\Omega)\in(e_{l})$;
  • Let be $\Omega$ of level $l$. Then $\Omega\in\cup_{\beta<\beta_{l+1}}\Delta_{\beta,<}\Leftrightarrow\nu_{\mathcal{C}}(\Omega)\in(e_{l})$;
  • Let be $\Omega$ of level $l$. Then $\Omega\in\Delta_{\beta_{l+1},<}^{0}\Leftrightarrow\nu_{\mathcal{C}}(\Omega)\in[\beta_{l+1}]_{e_{l}}$
Hence, a terminal family of level $l<g$ “covers” all the congruences modulo $n$ in $[(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ (cf. Definition 3.6).
Corollary 3.4.
We have $[\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}=[(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ for $l<g$.
Proof.
This holds because, on one hand, there are $2\nu_{l}$ congruences modulo $n$ in the set $(e_{l})\cup\{\beta_{l+1}+(e_{l})\}$ ($\nu_{l}$ from $(e_{l})$ and other $\nu_{l}$ from $\beta_{l+1}+(e_{l})$); and on the other hand, $[\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}$ has $2\nu_{l}$ elements by definition of terminal family. ∎
3.5.2. From level $l$ to level $l+1$
Suppose $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ is a terminal family of level $l$ for $0\leq l<g$. Our aim is to construct a terminal family of level $l+1$ from $\mathcal{T}_{l}$. This requires several stages. The first one (stage 00) is the construction of the initial family $\mathcal{G}_{l+1}^{0}$ of stage 00 (Definition 3.5).
Lemma 3.3.
$\mathcal{G}_{l+1}^{0}$ is an ordered family of level $l+1$.
Proof.
Property (i) holds by construction.
Notice that $\nu_{\mathcal{C}}(f_{l+1})=\overline{\beta}_{l+1}$, $\nu(f_{l+1})=\nu_{l}$ and $n\leq\nu_{\mathcal{C}}(\Omega_{j})\leq\overline{\beta}_{l+1}$ for $1\leq j\leq 2\nu_{l}$ by the properties of a terminal family of level $l$. Since the sequence $(\nu_{\mathcal{C}}(\Omega_{j}))_{1\leq j\leq 2\nu_{l}}$ satisfies property (ii) and
$\nu_{\mathcal{C}}(\Omega_{2\nu_{l}+1}^{0})-\nu_{\mathcal{C}}(\Omega_{2\nu_{l}}^{0})\geq(\overline{\beta}_{l+1}+n)-\overline{\beta}_{l+1}=n,$
property (ii) holds for $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{1\leq j\leq 2\nu_{l+1}}$. Thus $\mathcal{G}_{l+1}^{0}$ is ordered. ∎
Remark 3.13.
[…]
$\nu_{\mathcal{C}}(f_{l+1}^{a}\Omega_{j})=a\overline{\beta}_{l+1}+\nu_{\mathcal{C}}(\Omega_{j})\in(e_{l+1})$
for any $(a,j)\in{\mathfrak{S}}$, the set $[\nu_{\mathcal{C}}(\mathcal{G}_{l+1}^{0})]_{n}$ contains at most $\nu_{l+1}=n/e_{l+1}$ elements. Notice also that, since $m\overline{\beta}_{l+1}\not\in(e_{l})$ for any $1\leq m<n_{l+1}$, we have
$\nu_{\mathcal{C}}(f_{l+1}^{a}\Omega_{j})-\nu_{\mathcal{C}}(f_{l+1}^{b}\Omega_{k})\not\in(n)$
if $(a,j),(b,k)\in{\mathfrak{S}}$, $(a,j)\neq(b,k)$ and $\nu_{\mathcal{C}}(\Omega_{j})$ and $\nu_{\mathcal{C}}(\Omega_{k})$ belong both to $(e_{l})$ or both to $[\beta_{l+1}]_{e_{l}}$. In particular any element of $[(e_{l+1})]_{n}$ has at most two representatives in $(\nu_{\mathcal{C}}({\Omega}_{j}^{0}))_{j=1}^{2\nu_{l+1}}$. Hence, $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{j=1}^{2\nu_{l+1}}$ defines exactly the $\nu_{l+1}$ classes modulo $n$ in $[(e_{l+1})]_{n}$ and each class has two representatives in it.
This property will allow us to choose, from each class modulo $n$, the $1$-form whose representative appears later in the sequence and substitute it for another one using the key formula (6), in order to obtain the $2{\nu}_{l+1}$ new forms of the next stage $1$. There are two cases, namely $l+1<g$ and $l+1=g$. In the former case, this property of having two representatives in each class modulo $n$ is stable up to a final stage in which there is only one representative in each class modulo $n$; the forms at this stage will be the terminal ones. In the case $l+1=g$, we will see that at some stage the family provides a nice basis.
We construct families $\mathcal{G}_{l+1}^{0},\mathcal{G}_{l+1}^{1},\ldots$ (see subsection 3.3) until we obtain a terminal family $\mathcal{G}_{l+1}^{r}$. The following definition summarizes all the properties we shall make use of when passing from stage $s$ to stage $s+1$, and they will guarantee that we can find a terminal family at some stage. Assume a family $\mathcal{G}_{l+1}^{s}=(\Omega_{1}^{s},\ldots,\Omega_{2{\nu}_{l+1}}^{s})$ of level $l+1$ at stage $s$ is given, with $0\leq l<g$ and $s\geq 1$. Recall that $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ is a terminal family of level $l$.
Definition 3.11.
The family $\mathcal{G}_{l+1}^{s}$ is a good family of stage $s\geq 1$ (in level $l+1$) if the following properties hold:
  1. (I)
    The first $2{\nu}_{l}$ elements come from the previous terminal family ${\mathcal{T}}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$: $\Omega_{j}^{s}=\Omega_{j}$ for any $1\leq j\leq 2{\nu}_{l}$.
  2. (II)
    $(\Omega_{1}^{s},\ldots,\Omega_{2{\nu}_{l+1}}^{s})$ is an ordered family of level $l+1$.
  3. (III)
    There is a partition of $\{1,2,\ldots,2\nu_{l+1}\}$ in two sets ${\mathcal{L}}_{s}$ and ${\mathcal{U}}_{s}$ of the same cardinal such that $\Omega_{j}^{s}\in\cup_{\beta^{\prime}<{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$ if $j\in{\mathcal{L}}_{s}$ and $\Omega_{j}^{s}\in\Delta_{{\mathfrak{n}}^{s}(\beta_{l+1}),<}^{0}$ if $j\in{\mathcal{U}}_{s}$.
  4. (IV)
    We have
    $\begin{split}\{[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}:j\in{\mathcal{L}}_{s}\}=[(e_{l+1})]_{n}&\ \ \mathrm{and}\\ \{[\nu_{\mathcal{C}}(\Omega_{j}^{s})-{\mathfrak{n}}^{s}(\beta_{l+1})]_{n}:j\in{\mathcal{U}}_{s}\}=[(e_{l+1})]_{n}.&\end{split}$
    In particular $[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}$ if $j\neq k$ and both $j,k$ belong to ${\mathcal{L}}_{s}$ or both belong to ${\mathcal{U}}_{s}$.
3.5.3. From stage $s$ to $s+1$ inside level $l+1$
Assume, from now on, that the families $\mathcal{G}^{0}_{l+1},\ldots,\mathcal{G}^{s}_{l+1}$ of level $l+1$ have been constructed for $s\geq 0$, $\mathcal{G}^{j}_{l+1}$ is good for $1\leq j\leq s$ and that
(15) $\{\beta_{l+1}={\mathfrak{n}}^{0}(\beta_{l+1}),{\mathfrak{n}}^{1}(\beta_{l+1}),\ldots,{\mathfrak{n}}^{s}(\beta_{l+1})\}\subset(e_{l+1}).$
Now, we explain how to construct a good family ${\mathcal{G}}_{l+1}^{s+1}$. If ${\mathcal{G}}_{l+1}^{s}$ is non-terminal, ${\mathcal{G}}_{l+1}^{s+1}$ is closer than ${\mathcal{G}}_{l+1}^{s}$ to be terminal. Later on, we will see that one of the good families constructed is indeed terminal.
Proposition 3.2.
$\mathcal{G}^{s+1}_{l+1}$ is well-defined.
Proof.
Consider $1\leq k\leq 2{\nu}_{l+1}$. There exists a unique $1\leq j\leq 2{\nu}_{l+1}$, with $j\neq k$, such that $\nu_{\mathcal{C}}(\Omega_{k}^{s})-\nu_{\mathcal{C}}(\Omega_{j}^{s})=dn$ for some $d\in{\mathbb{Z}}$. This holds by Remark 3.13 for the case $s=0$, and by Property (IV) for $s\geq 1$. Notice that ${\mathcal{G}}_{l+1}^{s}$ is ordered by Lemma 3.3 and Property (II). Hence, we get $d\leq 0$ if $j>k$ and $d\geq 0$ if $j<k$. Therefore, $\Omega_{k}^{s+1}=\Omega_{k}^{s}$ holds in the former case whereas $\Omega_{k}^{s+1}$ is given by substitution (6) in the latter case. We deduce that $\Omega_{k}^{s+1}$ is a well-defined element of ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$. ∎
Remark 3.14.
The equality of $\mathbb{C}[[x]]$-modules
$\sum_{j=1}^{2{\nu}_{l+1}}{\mathbb{C}}[[x]]\Omega_{j,\Gamma}^{s}=\sum_{j=1}^{2{\nu}_{l+1}}{\mathbb{C}}[[x]]\Omega_{j,\Gamma}^{s+1}=\sum_{j=1}^{2{\nu}_{l+1}}\mathbb{C}[[x]]\overline{\Omega}_{j}$
holds for generic $\Gamma\in\mathcal{C}$ (cf. Definition 2.6), by construction.
The next remark fixes the notation for the next results.
Remark 3.15.
Consider $1\leq j<k\leq 2{\nu}_{l+1}$ with $\nu_{\mathcal{C}}(\Omega_{k}^{s})-\nu_{\mathcal{C}}(\Omega_{j}^{s})=dn$ for some $d\in{\mathbb{Z}}_{\geq 0}$. Let $\beta^{\prime}$ and $\beta^{\prime\prime}$ be such that $\Omega_{j}^{s}\in\Delta_{\beta^{\prime},\leq}$, $\Omega_{k}^{s}\in\Delta_{\beta^{\prime\prime},\leq}$, and write
(16) $\begin{split}t\Gamma^{*}(\Omega_{j,\Gamma}^{s})&=\left(C_{\beta^{\prime},j}(\Gamma)t^{\nu_{\mathcal{C}}(\Omega_{j}^{s})}+\sum_{r\geq 1}C_{{\mathfrak{n}}^{r}(\beta^{\prime}),j}(\Gamma)t^{\nu_{\mathcal{C}}(\Omega_{j}^{s})+{\mathfrak{n}}^{r}(\beta^{\prime})-\beta^{\prime}}\right)dt\\ t\Gamma^{*}(\Omega_{k,\Gamma}^{s})&=\left(C_{\beta^{\prime\prime},k}(\Gamma)t^{\nu_{\mathcal{C}}(\Omega_{k}^{s})}+\sum_{r\geq 1}C_{{\mathfrak{n}}^{r}(\beta^{\prime\prime}),k}(\Gamma)t^{\nu_{\mathcal{C}}(\Omega_{k}^{s})+{\mathfrak{n}}^{r}(\beta^{\prime\prime})-\beta^{\prime\prime}}\right)dt,\end{split}$
where $C_{\beta^{\prime},j},C_{\beta^{\prime\prime},k}\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}\setminus\{0\}$. We have
(17) $\Omega_{k}^{s+1}=C_{\beta^{\prime},j}\Omega_{k}^{s}-C_{\beta^{\prime\prime},k}x^{d}\Omega_{j}^{s}.$
The following elementary results will be used when studying the properties of stage $s+1$ given stage $s$. Consider a family $\mathcal{G}_{l+1}^{s}$, either the initial family of level $l+1$ if $s=0$ or a good family if $s\geq 1$.
Lemma 3.4.
Assume that $[\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ for $j\neq k$, and that $\Omega^{s}_{k}\in\Delta_{{\mathfrak{n}}^{s}(\beta_{l+1}),<}$ and $\Omega^{s}_{j}\in\cup_{\beta^{\prime}<{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$. Let $(c_{{\mathfrak{n}}^{r}(\beta_{l+1})})_{r\geq 0}$ be the $c$-sequence of $\Omega^{s}_{k}$. Then:
$\Omega_{\max(j,k)}^{s+1}\in\Delta_{{\mathfrak{n}}^{s+1}(\beta_{l+1}),<}^{0},\,\,\,\Omega_{\min(j,k)}^{s+1}=\Omega_{\min(j,k)}^{s}$
and the $c$-sequence of $\Omega_{\max(j,k)}^{s+1}$ is equal to
$((c_{{\mathfrak{n}}^{s+1}(\beta_{l+1})})^{-1}(c_{{\mathfrak{n}}^{s+r}(\beta_{l+1})}))_{r\geq 1}.$
In particular,
$\nu_{\mathcal{C}}(\Omega_{\max(j,k)}^{s+1})-\nu_{\mathcal{C}}(\Omega_{\max(j,k)}^{s})={\mathfrak{n}}^{s+1}(\beta_{l+1})-{\mathfrak{n}}^{s}(\beta_{l+1})$
holds.
Proof.
This follows straightforwardly from equations (16) and (17). ∎
Lemma 3.5.
Suppose that $[\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}$ for $j<k$, and that $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ belong to $\tilde{\Delta}_{\beta_{l+1},\leq}$. Denote by $(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),j})_{r\geq 0}$ and $(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),k})_{r\geq 0}$ the $c$-sequences of $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ respectively. If $c_{{\mathfrak{n}}(\beta_{l+1}),j}\neq c_{{\mathfrak{n}}(\beta_{l+1}),k}$ and
$((c_{{\mathfrak{n}}(\beta_{l+1}),j}-c_{{\mathfrak{n}}(\beta_{l+1}),k})^{-1}(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),j}-c_{{\mathfrak{n}}^{r}(\beta_{l+1}),k}))_{r\geq 1}$
is an increasing sequence in ${\mathbb{Q}}^{+}$, then $\Omega_{k}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{l+1}),<}^{0}$ and the above sequence is the $c$-sequence of $\Omega_{k}^{1}$. In particular,
(18) $\nu_{\mathcal{C}}(\Omega_{k}^{1})-\nu_{\mathcal{C}}(\Omega_{k}^{0})={\mathfrak{n}}(\beta_{l+1})-\beta_{l+1}$
holds.
Proof.
With the notation (16), we have
$C_{{\mathfrak{n}}^{r}(\beta_{l+1}),\delta}=\tilde{c}_{\delta}c_{{\mathfrak{n}}^{r}(\beta_{l+1}),\delta}a_{{\mathfrak{n}}^{r}(\beta_{l+1})}R_{\delta}(a_{\beta_{1}},\ldots,a_{{\mathfrak{p}}(\beta_{l+1})})a_{\beta_{l+1}}^{m_{\delta}}+Q_{{\mathfrak{n}}^{r}(\beta_{l+1}),\delta}$
for $\delta\in\{j,k\}$ and $r\geq 0$, following Definition 3.8, where $Q_{\beta_{l+1},j}\equiv Q_{\beta_{l+1},k}\equiv 0$ by hypothesis.
If we now denote
$t\Gamma^{*}(\Omega_{k,\Gamma}^{1})=\left(\sum_{r\geq 1}C_{{\mathfrak{n}}^{r}(\beta_{l+1})}(\Gamma)t^{\nu_{\Gamma}(\Omega_{k}^{s})+{\mathfrak{n}}^{r}(\beta_{l+1})-\beta_{l+1}}\right)dt,$
then
$C_{{\mathfrak{n}}^{r}(\beta_{l+1})}=(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),k}-c_{{\mathfrak{n}}^{r}(\beta_{l+1}),j})[\tilde{c}_{j}\tilde{c}_{k}a_{\beta_{l+1}}^{m_{j}+m_{k}+1}R_{j}R_{k}]a_{{\mathfrak{n}}^{r}(\beta_{l+1})}+Q_{r},$
for $r\geq 1$, where $Q_{r}\in\mathbb{C}{[a_{\beta}]}_{\beta_{1}\leq\beta<{\mathfrak{n}}^{r}(\beta_{l+1})}$, and the result follows. ∎
We want to show that $\mathcal{G}^{s+1}_{l+1}$ is good. To this end, we need to distinguish the case $s=0$ and the rest. The former is the most complicated step of our construction.
Definition 3.12.
We define the initial sets ${\mathcal{L}}_{1}$ and ${\mathcal{U}}_{1}$ as follows:
  • Let $1\leq j\leq 2\nu_{l}$ such that $\Omega_{j}\in\Delta_{\beta^{\prime},<}$ with $\beta^{\prime}<\beta_{l+1}$. Then $j\in{\mathcal{L}}_{1}$ and $j+2r\nu_{l}\in{\mathcal{U}}_{1}$ for $1\leq r<n_{l+1}$.
  • Let $1\leq j\leq 2\nu_{l}$ such that $\Omega_{j}\in\Delta_{\beta^{\prime},<}$ with $\beta^{\prime}=\beta_{l+1}$. Then $j+2(n_{l+1}-1)\nu_{l}\in{\mathcal{U}}_{1}$ and $j+2r\nu_{l}\in{\mathcal{L}}_{1}$ for $20\leq r\leq n_{l+1}-$ 2.
Recall that the sets $M_{j}$ and $\hat{M}_{j}$ were defined in page 3.1.
Proposition 3.3.
The family ${\mathcal{G}}_{l+1}^{1}$ satisfies
  • ${\mathcal{U}}_{1}=\{1\leq k\leq 2\nu_{l+1}:\Omega_{k}^{1}\neq\Omega_{k}^{0}\}$.
  • $\Omega_{j}^{1}\in\cup_{\beta^{\prime}\leq\beta_{l+1}}\Delta_{\beta^{\prime},<}\cap\hat{M}_{j}$ for any $j\in{\mathcal{L}}_{1}$.
  • $\Omega_{k}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{l+1}),<}^{0}\cap\hat{M}_{k}$ for any $k\in{\mathcal{U}}_{1}$.
  • $\nu_{\mathcal{C}}(\Omega_{k}^{1})-\nu_{\mathcal{C}}(\Omega_{k}^{0})={\mathfrak{n}}(\beta_{l+1})-\beta_{l+1}$ for any $k\in{\mathcal{U}}_{1}$.
  • $\{[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}:j\in{\mathcal{L}}_{1}\}=[(e_{l+1})]_{n}$.
  • $\{[\nu_{\mathcal{C}}(\Omega_{k}^{1})-{\mathfrak{n}}(\beta_{l+1})]_{n}:k\in{\mathcal{U}}_{1}\}=[(e_{l+1})]_{n}$.
Proof.
Suppose that $\nu_{\mathcal{C}}(\Omega_{k}^{0})-\nu_{\mathcal{C}}(\Omega_{j}^{0})=dn$ for some $d\in{\mathbb{Z}}_{\geq 0}$ and $1\leq j<k\leq 2\nu_{l+1}$. First, we want to see that $j\in{\mathcal{L}}_{1}$ and $k\in{\mathcal{U}}_{1}$. As $\Omega_{j}^{0}=f_{l+1}^{m_{j}}\Omega_{a}$ and $\Omega_{k}^{0}=f_{l+1}^{m_{k}}\Omega_{b}$, we deduce that
$(m_{k}-m_{j})\overline{\beta}_{l+1}+(\nu_{\mathcal{C}}(\Omega_{b})-\nu_{\mathcal{C}}(\Omega_{a}))=dn.$
Moreover, $j<k$ implies $m_{j}\leq m_{k}$, and since $0\leq m_{j},m_{k}<n_{l+1}$, Remark 3.12 implies that this last equality can only happen in two cases
  • (a)
    $m_{k}=m_{j}+1$, $\nu_{\mathcal{C}}(\Omega_{b})\in(e_{l})$ and $\nu_{\mathcal{C}}(\Omega_{a})\in[\beta_{l+1}]_{e_{l}}$ or
  • (b)
    $m_{k}=n_{l+1}-1$, $m_{j}=0$, $\nu_{\mathcal{C}}(\Omega_{b})\in[\beta_{l+1}]_{e_{l}}$ and $\nu_{\mathcal{C}}(\Omega_{a})\in(e_{l})$.
Since any $[x]_{n}\in[(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{1\leq j\leq 2\nu_{l+1}}$ by Remark 3.13, we deduce ${\mathcal{U}}_{1}=\{1\leq r\leq 2\nu_{l+1}:\Omega_{r}^{1}\neq\Omega_{r}^{0}\}$.
Next, let us show $\Omega_{j}^{1}\in\cup_{\beta^{\prime}\leq\beta_{l+1}}\Delta_{\beta^{\prime},<}\cap\hat{M}_{j}$, $\Omega_{k}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{l+1}),<}^{0}\cap\hat{M}_{k}$ and equation (18).
Case (a). If $m_{k}=m_{j}+1$, then $\Omega_{a}\in\Delta_{\beta_{l+1},<}^{0}$ and $\Omega_{b}\in\Delta_{\beta,<}$ for some $\beta<\beta_{l+1}$ by Remark 3.12. Denote by $(c_{{\mathfrak{n}}^{r}(\beta_{l+1})})_{r\geq 0}$ the $c$-sequence of $\Omega_{a}$. Since $\beta<\beta_{l+1}$, notice that $\Omega_{k}^{0}=f_{l+1}^{m_{j}+1}\Omega_{b}$ belongs to $\tilde{\Delta}_{\beta_{l+1},\leq}^{m_{j}}\setminus\Delta_{\beta_{l+1},<}^{m_{j}}$ and its $c$-sequence is $1,m_{j}+1,\ldots,m_{j}+1,\ldots$ by Corollary 3.2. We have $\Omega_{j}^{1}=\Omega_{j}^{0}\in\tilde{\Delta}_{\beta_{l+1},<}^{m_{j}}\cap\hat{M}_{j}$ and that its $c$-sequence is
$1,c_{{\mathfrak{n}}(\beta_{l+1})}+m_{j},c_{{\mathfrak{n}}^{2}(\beta_{l+1})}+m_{j},\ldots$
by Corollary 3.3. Since $(c_{{\mathfrak{n}}^{r}(\beta_{l+1})})_{r\geq 0}$ is strictly increasing, we deduce that $\Omega_{k}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{l+1}),<}^{0}\cap\hat{M}_{k}$, that its $c$-sequence is
$((c_{{\mathfrak{n}}(\beta_{l+1})}-1)^{-1}(c_{{\mathfrak{n}}^{r}(\beta_{l+1})}-1))_{r\geq 1}$
and equation (18) by Lemma 3.5.
Case (b). That is, $m_{k}=n_{l+1}-1$, $m_{j}=0$. We know that $\Omega_{j}^{0}\in\Delta_{\beta,<}$ for some $\beta<\beta_{l+1}$ and $f_{l+1}^{n_{l+1}-1}\Omega_{b}\in\Delta_{\beta_{l+1},<}^{n_{l+1}-1}$. Lemma 3.4 implies that $\Omega_{k}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{l+1}),<}^{0}\cap\hat{M}_{k}$ and equation (18).
It remains to show the last two properties. We have $\nu_{\mathcal{C}}(\Omega_{j}^{0})\in(e_{l+1})$ for any $1\leq j\leq 2\nu_{l+1}$. This equality and equation (18) imply $\nu_{\mathcal{C}}(\Omega_{j}^{1})\in(e_{l+1})$ for any $j\in{\mathcal{L}}_{1}$ and $\nu_{\mathcal{C}}(\Omega_{k}^{1})-{\mathfrak{n}}(\beta_{l+1})\in(e_{l+1})$ for any $k\in{\mathcal{U}}_{1}$. By construction, we have
$[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{1})]_{n}$
if $j\neq j^{\prime}$ and $j,j^{\prime}\in{\mathcal{L}}_{s}$. Thus $\{[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}:j\in{\mathcal{L}}_{1}\}$ and $[(e_{l+1})]_{n}$ are sets of cardinal $\nu_{l+1}$ such that the former is included in the latter and as a consequence they coincide. Analogously, $[\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k^{\prime}}^{0})]_{n}$ holds if $k\neq k^{\prime}$ and $k,k^{\prime}\in{\mathcal{U}}_{1}$. This property together with equation (18) imply that $\{[\nu_{\mathcal{C}}(\Omega_{k}^{1})-{\mathfrak{n}}(\beta_{l+1})]_{n}:k\in{\mathcal{U}}_{1}\}$ is a set of $\nu_{l+1}$ elements contained in $[(e_{l+1})]_{n}$ and hence they coincide. ∎
The step from stage $s$ to stage $s+1$ for $s>0$ is easier. Suppose that the families $\mathcal{G}_{l+1}^{1},\ldots,\mathcal{G}_{l+1}^{s}$, with $\mathcal{G}_{l+1}^{r}=(\Omega_{j}^{r}:j=1,\ldots,2{\nu}_{l+1})$ are good for $1\leq r\leq s$ and Property (15) holds.
Definition 3.13.
We define ${\mathcal{U}}_{s+1}=\{1\leq k\leq 2\nu_{l+1}:\Omega_{k}^{s+1}\neq\Omega_{k}^{s}\}$ and ${\mathcal{L}}_{s+1}$ as its complement in $\{1,\ldots,2\nu_{l+1}\}$.
Proposition 3.4.
The family ${\mathcal{G}}_{l+1}^{s+1}$ satisfies
  • Both ${\mathcal{L}}_{s+1}$ and ${\mathcal{U}}_{s+1}$ have $\nu_{l+1}$ elements.
  • $\Omega_{j}^{s+1}\in\cup_{\beta^{\prime}<{\mathfrak{n}}^{s+1}(\beta_{l+1})}\Delta_{\beta^{\prime},<}\cap\hat{M}_{j}$ for any $j\in{\mathcal{L}}_{s+1}$.
  • $\Omega_{k}^{s+1}\in\Delta_{{\mathfrak{n}}^{s+1}(\beta_{l+1}),<}^{0}\cap\hat{M}_{k}$ for any $k\in{\mathcal{U}}_{s+1}$.
  • $\nu_{\mathcal{C}}(\Omega_{k}^{s+1})-\nu_{\mathcal{C}}(\Omega_{k}^{s})={\mathfrak{n}}^{s+1}(\beta_{l+1})-{\mathfrak{n}}^{s}(\beta_{l+1})$ for any $k\in{\mathcal{U}}_{s+1}$.
Proof.
As $\mathcal{G}_{l+1}^{s}$ is good, $\Omega_{j}^{s}\in\cup_{\beta^{\prime}\leq{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$ for any $1\leq j\leq 2{\nu}_{l+1}$. Since ${\mathfrak{n}}^{s}(\beta_{l+1})\in(e_{l+1})$ by hypothesis, any element of $[(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{s}))_{1\leq j\leq 2\nu_{l+1}}$ by property (IV) for $s$. Thus ${\mathcal{L}}_{s+1}$ and ${\mathcal{U}}_{s+1}$ have both $\nu_{l+1}$ elements.
Suppose $\nu_{\mathcal{C}}(\Omega_{k}^{s})-\nu_{\mathcal{C}}(\Omega_{j}^{s})=dn$ for some $d\in{\mathbb{Z}}_{\geq 0}$ and $1\leq j<k\leq 2{\nu}_{l+1}$. As $\nu_{\mathcal{C}}(\Omega_{j}^{s})-\nu_{\mathcal{C}}(\Omega_{k}^{s})\in(n)$, the $1$-forms $\Omega_{j}^{s},\Omega_{k}^{s}$ have different leading variables, also by property (IV) for $s$. As a consequence,
$\Omega_{j}^{s+1}=\Omega_{j}^{s}\in\cup_{\beta^{\prime}<{\mathfrak{n}}^{s+1}(\beta_{l+1})}\Delta_{\beta^{\prime},<}\quad\mathrm{and}\quad\Omega_{k}^{s+1}\in\Delta_{{\mathfrak{n}}^{s+1}(\beta_{l+1}),<}^{0}\cap\hat{M}_{k}$
by Lemma 3.4. The last property also follows from Lemma 3.4. ∎
We can now show that being good is a hereditary property.
Proposition 3.5.
Given $\mathcal{G}^{0}_{l+1},\ldots,\mathcal{G}^{s}_{l+1}$, assume that $\mathcal{G}^{j}_{l+1}$ is good for any $1\leq j\leq s$ and Property (15) holds. Then $\mathcal{G}^{s+1}_{l+1}$ is good.
Proof.
Property (I) in Definition 3.11 and property (i) of ordered families hold by construction.
Property (III) for $\mathcal{G}_{l+1}^{s+1}$ (for the leading variables and properties of the $c$-sequences) follows for $s=0$ applying Proposition 3.3. For $s>0$, apply Proposition 3.4 using that $\mathcal{G}_{l+1}^{s}$ is good.
Let us show Property (IV) for $\mathcal{G}_{l+1}^{s+1}$. It is a consequence of Proposition 3.3 if $s=0$ so we can assume $s\geq 1$. By construction, we obtain
$\{[\nu_{\mathcal{C}}(\Omega_{j}^{s+1})]_{n}:j\in{\mathcal{L}}_{s+1}\}=\{[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}:j\in{\mathcal{L}}_{s}\}=[(e_{l+1})]_{n}$
and
$\{[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}:k\in{\mathcal{U}}_{s+1}\}=\{[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}:k\in{\mathcal{U}}_{s}\}.$
This property, Proposition 3.4 and Property (IV) for $\mathcal{G}_{l+1}^{s}$ imply
$\{[\nu_{\mathcal{C}}(\Omega_{k}^{s+1})-{\mathfrak{n}}^{s+1}(\beta_{l+1})]_{n}:k\in{\mathcal{U}}_{s+1}\}=\{[\nu_{\mathcal{C}}(\Omega_{k}^{s})-{\mathfrak{n}}^{s}(\beta_{l+1})]_{n}:k\in{\mathcal{U}}_{s}\},$
concluding the proof of Property (IV) for $\mathcal{G}_{l+1}^{s+1}$.
Since ${\mathfrak{n}}^{s}(\beta_{l+1})\in(e_{l+1})$, it follows that $\nu_{\mathcal{C}}(\Omega_{j}^{s})\in(e_{l+1})$ for any $1\leq j\leq 2\nu_{l+1}$ by Property (IV) for $\mathcal{G}_{l+1}^{s}$. Since
$\nu_{\mathcal{C}}(\Omega_{k}^{s})-\nu_{\mathcal{C}}(\Omega_{j}^{s})\in(e_{l+1})\ \mathrm{and}\ \nu_{\mathcal{C}}(\Omega_{k}^{s})\geq\nu_{\mathcal{C}}(\Omega_{j}^{s})$
for all $1\leq j\leq k\leq 2{\nu}_{l+1}$, we obtain
(19) $\nu_{\mathcal{C}}(\Omega_{k}^{s})-\nu_{\mathcal{C}}(\Omega_{j}^{s})\geq\delta_{j,k}e_{l+1}$
where $\delta_{j,k}=0$ if $\nu_{\mathcal{C}}(\Omega_{k}^{s})-\nu_{\mathcal{C}}(\Omega_{j}^{s})\in(n)$ and otherwise $\delta_{j,k}=1$. Since ${\mathfrak{n}}^{s+1}(\beta_{l+1})-{\mathfrak{n}}^{s}(\beta_{l+1})\leq e_{l+1}$, we get
(20) $\nu_{\mathcal{C}}(\Omega_{k}^{s+1})\geq\nu_{\mathcal{C}}(\Omega_{k}^{s})\quad\mathrm{and}\quad\nu_{\mathcal{C}}(\Omega_{j}^{s+1})\leq\nu_{\mathcal{C}}(\Omega_{j}^{s})+e_{l+1}$
and therefore equations (19) and (20) imply
(21) $\nu_{\mathcal{C}}(\Omega_{k}^{s+1})-\nu_{\mathcal{C}}(\Omega_{j}^{s+1})\geq 0$
if $1\leq j\leq k\leq 2{\nu}_{l+1}$ and $\delta_{j,k}=1$. Assume $\delta_{j,k}=0$. Since
$\nu_{\mathcal{C}}(\Omega_{k}^{s+1})>\nu_{\mathcal{C}}(\Omega_{k}^{s})\quad\mathrm{and}\quad\nu_{\mathcal{C}}(\Omega_{j}^{s+1})=\nu_{\mathcal{C}}(\Omega_{j}^{s}),$
Equation (19) implies equation (21). Therefore, property (ii) holds, so that $\mathcal{G}_{l+1}^{s+1}$ is ordered, and we are done. ∎
Finally, we prove that after a finite number of stages one obtains a terminal family.
Proposition 3.6.
Suppose that $(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ is a terminal family of level $l$. Then there exists $s\geq 1$ such that $\mathcal{G}_{l+1}^{s}=(\Omega_{j}^{s}:j=1,\ldots,2{\nu}_{l+1})$ is terminal of level $l+1$, with $\Omega_{j}^{s}=\Omega_{j}$ for any $1\leq j\leq 2{\nu}_{l}$. Moreover, if $l+1<g$ then ${\mathfrak{n}}^{s}(\beta_{l+1})=\beta_{l+2}$.
Proof.
We need to verify the three properties of a terminal family (see Definition 3.10) for some $s\geq 1$. Consider a good family $\mathcal{G}_{l+1}^{s}$ of level $l+1$. Since any good family is ordered by Lemma 3.3 and Definition 3.11, it follows that $\mathcal{G}_{l+1}^{s}$ satisfies condition (a) for terminal families if and only if satisfies condition (iii) in Definition 3.4.
We divide the verification of this and the remaining properties in two cases: $l+1<g$ and $l+1=g$.
Case $l+1<g$. Let $s_{0}$ be such that ${\mathfrak{n}}^{s_{0}}(\beta_{l+1})=\beta_{l+2}$. We can construct good families ${\mathcal{G}}_{l+1}^{1},\ldots,{\mathcal{G}}_{l+1}^{s_{0}}$ by applying $s_{0}$ times Proposition 3.5. Suppose $s<s_{0}$. Since ${\mathfrak{n}}^{s}(\beta_{l+1})\in(e_{l+1})$, it follows that
$\left\{[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\ |\ \Omega_{j}^{s}\in\mathcal{G}_{l+1}^{s}\right\}=[(e_{l+1})]_{n}$
by Property (IV) for $\mathcal{G}_{l+1}^{s}$ and hence $\mathcal{G}_{l+1}^{s}$ is non-terminal because it is non-separated.
We claim that ${\mathcal{G}}_{l+1}^{s_{0}}$ is a terminal family. Since $\beta_{l+2}\not\in(e_{l+1})$, we deduce that all the classes modulo $n$ in $\{[\nu_{\mathcal{C}}(\Omega_{j}^{s_{0}})]_{n}\ |\ j=1,\ldots,2{\nu}_{l+1}\}$ are different by Property (IV). Therefore, ${\mathcal{G}}_{l+1}^{s_{0}}$ is separated, and thus condition (a) in the definition of terminal family 3.10 holds.
By definition, the maximum ${\mathcal{C}}$-order of a $1$-form in a terminal family $\mathcal{T}_{l}$ of level $l$ is $\overline{\beta}_{l+1}$. Since a $1$-form $\Omega_{k}^{0}$ of stage 00 and level $l+1$ is of the form $f_{l+1}^{a}\Omega$ where $0\leq a<n_{l+1}$ and $\Omega\in\mathcal{T}_{l}$, we obtain
$\nu_{\mathcal{C}}(\Omega_{k}^{0})\leq\overline{\beta}_{l+1}+(n_{l+1}-1)\overline{\beta}_{l+1}=n_{l+1}\overline{\beta}_{l+1}$
for any $1\leq k\leq 2{\nu}_{l+1}$. As ${\mathfrak{n}}^{s_{0}}(\beta_{l+1})=\beta_{l+2}$, the construction gives
$\nu_{\mathcal{C}}(\Omega_{k}^{s_{0}})\leq\nu_{\mathcal{C}}(\Omega_{k}^{0})+(\beta_{l+2}-\beta_{l+1})\leq n_{l+1}\overline{\beta}_{l+1}+(\beta_{l+2}-\beta_{l+1})=\overline{\beta}_{l+2}$
for any $1\leq k\leq 2{\nu}_{l+1}$ by equation (2). Thus ${\mathcal{G}}_{l+1}^{s_{0}}$ satisfies condition (b). Finally, condition (c) is a consequence of property (III).
Case $l+1=g$. Since $e_{g}=1$, we obtain good families ${\mathcal{G}}_{g}^{1},{\mathcal{G}}_{g}^{2},\ldots$ by iterative applications of Proposition 3.5. Let ${\mathcal{G}}_{g}^{s}$ be a good family. Condition (b) is trivially satisfied. Let $\mathcal{B}_{s}$ be the set of congruences $[x]_{n}$ modulo $n$ such that there exist $\Omega^{s}_{j},\Omega_{k}^{s}$ with $[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}=[x]_{n}$ and $j<k\leq n$. We claim that $\mathcal{G}_{g}^{s}$ is terminal if and only if $\mathcal{B}_{s}=\emptyset$. Let us show the non-trivial implication, assume $\mathcal{B}_{s}=\emptyset$. Since this implies
[…]
${\mathcal{G}}_{g}^{s}$ is separated. Condition (c) is again a consequence of property (III).
Let us conclude by proving that there is some $s_{0}\geq 1$ such that $\mathcal{B}_{s_{0}}=\emptyset$. In this case, ${\mathfrak{n}}^{s}(\beta_{g})=\beta_{g}+s$ since $e_{g}=1$. The construction implies
$\begin{split}\{[\nu_{\mathcal{C}}(\Omega_{j}^{s+1})]_{n}\ |\ j=1,\ldots,n\}&=\\ \{[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\ &|\ j=1,\ldots,n\}\cup\{[x]_{n}+[1]_{n}\ |\ [x]_{n}\in\mathcal{B}_{s}\}.\end{split}$
A finiteness argument shows that there must be an $s_{0}\geq 1$ such that $\mathcal{B}_{s_{0}}=\emptyset$, and we are done. ∎
3.6. The terminal family of level $g$
Thus, beginning with the $1$-forms $\Omega_{1}=dx$ and $\Omega_{2}=df_{1}$ of level 00, we obtain a terminal family of $1$-forms $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{2n})$ of level $g$ by applying $g$ times Proposition 3.6.
Theorem 3.1.
The family $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ is a nice basis for ${\mathcal{C}}$.
Proof.
Consider $\Gamma\in{\mathcal{C}}$. Given $\hat{f}\in{\mathbb{C}}[[x,y]]$, there exists $R=\sum_{l=0}^{n-1}\kappa_{l}(x)y^{l}$ in ${\mathbb{C}}[[x]][y]$ such that $\hat{f}-R\in(f_{g+1,\Gamma})$ (cf. subsection 2.4) by the Weierstrass division theorem. It follows that
(22) $\hat{\Omega}\mbox{(${\mathbb{C}}^{2},0$)}={\mathcal{M}}_{2\nu_{g}}+{\mathcal{I}}_{\Gamma}$
for any $\Gamma\in{\mathcal{C}}$, cf. Definitions 3.1 and 3.2.
Assume, from now on, that $\Gamma\in{\mathcal{C}}$ is generic. Since
$\sum_{l=1}^{2n}\mathbb{C}[[x]]\Omega_{l,\Gamma}={\mathcal{M}}_{2\nu_{g}}$
by equation (4), we deduce $\hat{\Omega}\mbox{(${\mathbb{C}}^{2},0$)}=\sum_{l=1}^{2n}\mathbb{C}[[x]]\Omega_{l,\Gamma}+{\mathcal{I}}_{\Gamma}$ by equation (22). Since $\mathbf{\Omega}$ is separated, it is easy to see that
$\Omega_{j,\Gamma}\in\sum_{l=1}^{n}\mathbb{C}[[x]]\Omega_{l,\Gamma}+{\mathcal{I}}_{\Gamma}$
for any $n<j\leq 2n$ and hence
$\sum_{l=1}^{n}\mathbb{C}[[x]]\Omega_{l,\Gamma}+{\mathcal{I}}_{\Gamma}=\sum_{l=1}^{2n}\mathbb{C}[[x]]\Omega_{l,\Gamma}+{\mathcal{I}}_{\Gamma}=\hat{\Omega}\mbox{(${\mathbb{C}}^{2},0$)}$
for generic $\Gamma\in\mathcal{C}$. Therefore, $({\Omega}_{1},\ldots,{\Omega}_{n})$ is a nice basis for ${\mathcal{C}}$. ∎
Proposition 3.7.
Let $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ be a nice basis for ${\mathcal{C}}$ and $1\leq k\leq n$. Then
$\nu_{\mathcal{C}}(\Omega_{k})=\max\{\nu_{\mathcal{C}}(\Omega^{\prime}):\Omega^{\prime}\in\hat{M}_{k}\}.$
Proof.
By construction, given $\Omega^{\prime}\in\hat{M}_{k}$ there exists $P\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}\setminus\{0\}$ such that $P\Omega^{\prime}$ is of the form
$c_{k}\Omega_{k}+\sum_{j=1}^{k-1}c_{j}\Omega_{j},$
where $c_{k}\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}\setminus\{0\}$ and $c_{j}\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x]]$ for any $1\leq j<k$. Since $\nu_{\mathcal{C}}(c_{b}\Omega_{b})-\nu_{\mathcal{C}}(c_{a}\Omega_{a})\not\in(n)$ for all $1\leq a<b\leq k$, it follows that
$\nu_{\mathcal{C}}(\Omega^{\prime})=\nu_{\mathcal{C}}(P\Omega^{\prime})=\min_{1\leq j\leq k,\ c_{j}\not\equiv 0}\nu_{\mathcal{C}}(c_{j}\Omega_{j})\leq\nu_{\mathcal{C}}(c_{k}\Omega_{k})=\nu_{\mathcal{C}}(\Omega_{k}).$
Since $\Omega_{k}\in\hat{M}_{k}$, we are done. ∎
Remark 3.16.
The analogue result also holds for terminal families of any level.
Corollary 3.5.
The $n$-tuple $(\nu_{\mathcal{C}}(\Omega_{1}),\ldots,\nu_{\mathcal{C}}(\Omega_{n}))$ does not depend on the nice basis $(\Omega_{1},\ldots,\Omega_{n})$ for $\mathcal{C}$.
Remark 3.17.
Consider a nice basis $(\Omega_{1},\ldots,\Omega_{n})$ for $\mathcal{C}$. Let $l<g$. We have $df_{l+1}\in\hat{M}_{2\nu_{l}}\cap\Delta_{\beta_{l+1},<}^{0}$ and $\nu_{\mathcal{C}}(df_{l+1})=\overline{\beta}_{l+1}$. Note that $\nu_{\mathcal{C}}(\Omega_{2\nu_{l}})\leq\overline{\beta}_{l+1}$ by the properties of the $1$-forms of level $l$ and Corollary 3.5. Since
$\nu_{\mathcal{C}}(\Omega_{2\nu_{l}})=\max\{\nu_{\mathcal{C}}(\Omega^{\prime}):\Omega^{\prime}\in\hat{M}_{2\nu_{l}}\}\geq\nu_{\mathcal{C}}(df_{l+1})=\overline{\beta}_{l+1}$
by Proposition 3.7, we obtain $\nu_{\mathcal{C}}(\Omega_{2\nu_{l}})=\overline{\beta}_{l+1}$. Thus, we could replace $\Omega_{2\nu_{l}}$ with $df_{l+1}$ in the nice basis.
Remark 3.18.
Since $\nu_{\mathcal{C}}(f_{l})=\overline{\beta}_{l}$ for $1\leq l\leq g$, it is easy to build an algorithm that provides $\Lambda_{\Gamma}$, for generic $\Gamma\in\mathcal{C}$, and whose input is $n,\beta_{1},\ldots,\beta_{g}$, or equivalently $n,\overline{\beta}_{1},\ldots,\overline{\beta}_{g}$, by Remark 3.8.
4. Case $E\neq\emptyset$
We now study the equisingularity class $\mathcal{C}$ of a pair $(\Gamma,E)$ where $\Gamma$ is a germ of irreducible curve and $E\subset\{xy=0\}$ is a nonempty normal crossings divisor. We give a procedure to construct a basis of the module of Kähler differentials preserving $E$ on any generic element in ${\mathcal{C}}$ which will allow us to provide a constructive proof of Genzmer’s formula [13] for the dimension of the moduli space. This procedure is quite similar to the one for $E=\emptyset$, so that we shall only indicate the differences.
4.1. Setting for the case $E\neq\emptyset$
In order to study $\Lambda_{\mathcal{C}}^{E}$ (Definition 2.7), we need to adapt the technique of the preciding section, as the previously defined $\Omega_{1}$ and $\Omega_{2}$ no longer need to leave $E$ invariant.
Definition 4.1.
Let $E=\{x^{\delta_{x}}y^{\delta_{y}}=0\}$ be a non-empty normal-crossings divisor where $\delta_{x},\delta_{y}\in\{0,1\}$. We define:
  • If $E=\{x=0\}$, then
    $\overline{\Omega}_{1}=dx,\quad\overline{\Omega}_{2}=xdy.$
  • Else $\{y=0\}\subset E$, and then
    $\overline{\Omega}_{1}=x^{\delta_{x}}dy,\quad\overline{\Omega}_{2}=ydx.$
Given $\overline{\Omega}_{1}$ and $\overline{\Omega}_{2}$, we define for $j\geq 1$:
$\overline{\Omega}_{2j+1}=y^{j}\overline{\Omega}_{1},\quad\overline{\Omega}_{2j+2}=y^{j}\overline{\Omega}_{2}.$
Finally, $M_{0}^{E}=\{0\}$ and, recursively:
$M_{j+1}^{E}=\mathbb{C}[a_{\beta}]_{\beta\in\mathcal{E}(\mathcal{C})}[[x]]\overline{\Omega}_{j+1}+M_{j}^{E},\ \hat{M}_{j+1}^{E}=(\mathbb{C}[a_{\beta}]_{\beta\in\mathcal{E}(\mathcal{C})}\setminus\{0\})\overline{\Omega}_{j+1}+M_{j}^{E}.$
Definition 4.2.
The set of $1$-forms of $E$-level $l$ is now defined as $M_{2{\nu}_{l}}^{E}$ for $l\in\{0,\ldots,g\}$. Let $\mathcal{I}_{E,\Gamma}$ be the set of $1$-forms $\Omega\in\hat{\Omega}_{E}(\mathbb{C}^{2},0)$ such that $\Gamma^{\ast}\Omega\equiv 0$.
We need to refine the definition of ordered and separated family:
Definition 4.3.
A family $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{r})$ in $\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, with $r\in\{2\nu_{l},\min(2\nu_{l},n)\}$, is $E$-ordered of level $0\leq l\leq g$ for $\Gamma$ if:
  • $E$-(i)
    $\Omega_{j}\in\hat{M}_{j}^{E}$ for $1\leq j\leq r$;
  • $E$-(ii)
    $\nu_{C}(\Omega_{k})\geq\nu_{C}(\Omega_{j})$ for $1\leq j\leq k\leq r$.
Moreover, $\mathbf{\Omega}$ is $E$-separated of level $l$ if it is ordered and
  • $E$-(iii)
    $\nu_{\mathcal{C}}(\Omega_{1}),\ldots,\nu_{\mathcal{C}}(\Omega_{\min(2\nu_{l},n)})$ define pairwise different classes modulo $n$.
Finally, for $l=g$ and $r=n$, $\mathbf{\Omega}$ is an $E$-nice basis if, moreover,
  • $E$-(iv)
    $\hat{\Omega}_{E}(\mathbb{C}^{2},0)=\mathbb{C}[[x]]\Omega_{1,\Gamma}+\cdots+\mathbb{C}[[x]]\Omega_{n,\Gamma}+\mathcal{I}_{E,\Gamma}$ for generic $\Gamma\in\mathcal{C}$.
Notice that the only modification is to require that $E$ is invariant, in $E$-(i) and $E$-(ii).
The definitions of terminal and good need also adapting to $E$. For technical reasons which will become clear shortly, instead of starting the construction at level $l=0$, we need to set an initial level $l_{0}$, with $l_{0}\in\{0,1\}$, and we assume, for the moment, that there exists, for that $l_{0}$, a family satisfying the following conditions:
Definition 4.4.
A family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}\subset\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, with $l_{0}\leq l\leq g$, is $E$-terminal if
  • $E$-(a)
    $\mathcal{T}_{l}$ is an $E$-separated family of level $l$;
  • $E$-(b)
    On one hand $\nu_{\mathcal{C}}(\Omega_{2{\nu}_{l_{0}}})\leq n+\overline{\beta}_{l_{0}+1}$ if $l_{0}<g$, and on the other,
    $\nu_{\mathcal{C}}(\Omega_{2{\nu}_{l}})-\nu_{\mathcal{C}}(\Omega_{2{\nu}_{l-1}})\leq\overline{\beta}_{l+1}-\overline{\beta}_{l},\ \ \mathrm{if}\ l_{0}<l<g,$
    (this inequality the same as in Definition 3.10 except for the initial level $l_{0}$).
  • $E$-(c)
    The same as (c) in Definition 3.10: $\mathcal{T}_{l}\subset\cup_{\beta^{\prime}\in{\mathcal{E}}({\mathcal{C}})}\Delta_{\beta^{\prime},<}$. Moreover, we have $\mathcal{T}_{l}\subset\Delta_{\beta_{l+1},<}^{0}\cup\cup_{\beta^{\prime}\in{\mathcal{E}}({\mathcal{C}}),\ \beta^{\prime}<\beta_{l+1}}\Delta_{\beta^{\prime},<}$ for $0\leq l<g$.
Definition 4.5.
Given an $E$-terminal family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ of level $0\leq l<g$, we define an $E$-initial family ${\mathcal{G}}_{l+1}^{0}=(\Omega_{1}^{0},\ldots,\Omega_{2\nu_{l+1}}^{0})$ of stage 00 (in level $l+1$) by means of equation (5), exactly as in the case $E=\emptyset$.
Let $\mathcal{G}_{l+1}^{s}=(\Omega_{1}^{s},\ldots,\Omega^{s}_{2{\nu}_{l+1}})$ be given with $\Omega_{i}^{s}\in\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, for $0\leq l<g$ and $s\geq 1$.
Definition 4.6.
The family $\mathcal{G}_{l+1}^{s}$ is $E$-good (in level $l+1$, but we omit this) if the following properties hold:
  1. (I)
    The first $2{\nu}_{l}$ elements form an $E$-terminal family of level $l$. That is, there is a terminal family ${\mathcal{T}}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ of level $l$, such that $\Omega_{j}^{s}=\Omega_{j}$ for any $1\leq j\leq 2{\nu}_{l}$.
  2. (II)
    $(\Omega_{1}^{s},\ldots,\Omega_{2{\nu}_{l+1}}^{s})$ is $E$-ordered of level $l+1$.
  3. (III)
    Finally, conditions (III) and (IV) for good families hold for $\mathcal{G}^{s}_{l+1}$ (Definition 3.11).
4.2. The method for the case $E\neq\emptyset$
With the above definitions, all the results in subsection 3.5.3 carry over adding the prefix $E$ to each ordered, separated, terminal and good family, as well as for a nice basis, as they are only dependent on the notion of leading variable, which has not changed, and the conditions we have adapted have no bearing on the computations modulo $n$. However, we need to bootstrap the process at level $l_{0}$, which is the foundational step and is still missing. This is what we carry out in what follows.
When $E=\emptyset$, the terminal family of level 00 was ${\mathcal{T}}_{0}=\{dx,df_{1}\}$. From this family, applying the results in subsections 3.5.2 and 3.5.3 we ended up with the terminal family ${\mathcal{T}}_{g}$.
When $E\neq\emptyset$, the construction is totally similar: given an $E$-terminal family $\mathcal{T}_{l}$ of level $l_{0}\leq l<g$, we define ${\mathcal{G}}_{l+1}^{0}=(\Omega_{1}^{0},\ldots,\Omega^{0}_{2{\nu}_{l+1}})$ as in Definition 3.5. We have
$\nu_{\mathcal{C}}(\Omega_{2{\nu}_{l}+1}^{0})\geq\nu_{\mathcal{C}}(f_{l+1}dx)=n+\overline{\beta}_{l+1}\geq\nu_{\mathcal{C}}(\Omega_{2{\nu}_{l}})=\nu_{\mathcal{C}}(\Omega_{2{\nu}_{l}}^{0})$
and analogously we obtain $\nu_{\mathcal{C}}(\Omega_{2k{\nu}_{l}+1}^{0})\geq\nu_{\mathcal{C}}(\Omega_{2k{\nu}_{l}}^{0})$ for any $k\geq 1$. As a consequence, ${\mathcal{G}}_{l+1}^{0}$ is an ordered family. Then we just apply the results in subsection 3.5.3 to obtain an $E$-terminal family of level $l+1$. By repeating this process, we end up with an $E$-terminal family of level $g$ and hence an $E$-nice basis.
Remark 4.1.
Thus, we only need to find an $E$-terminal family of some level $l_{0}$. Actually, it suffices to find an $E$-good family ${\mathcal{G}}_{l_{0}}^{1}$ (with $E$-(I) being empty) since in such a case the results in subsection 3.5.3 provide an $E$-terminal family of level $l_{0}$.
The rest of this section is devoted to providing this initial $E$-terminal family. There are three cases to consider: $\delta_{y}=0$, so that $E=\{x=0\}$; $\delta_{y}=1$ with $\beta_{0}\in(n)$; and $\delta_{y}=1$ with $\beta_{0}\not\in(n)$.
4.2.1. Case $E=\{x=0\}$
In this case, we set:
$\Omega_{1}=dx,\quad\Omega_{2}=xdf_{1}.$
which satisfy $\nu(\Omega_{1})=0$, $\nu_{\mathcal{C}}(\Omega_{1})=n$, $\nu(\Omega_{2})=1$, $\nu_{\mathcal{C}}(\Omega_{2})=\beta_{1}+n$. The family $\mathcal{T}_{0}=\{\Omega_{1},\Omega_{2}\}$ is clearly $E$-terminal of level $l_{0}=0$ and thus, we obtain an $E$-nice basis $\mathbf{\Omega}$ by Remark 4.1.
4.2.2. Case $\delta_{y}=1$ and $\beta_{0}\in(n)$
Notice that $\beta_{0}\in(n)$ implies $\beta_{0}<\beta_{1}$. We have $\overline{\Omega}_{1}=x^{\delta_{x}}dy$ and $\overline{\Omega}_{2}=ydx$ and hence $\nu_{C}(\overline{\Omega}_{1})\in(n)$, $\nu_{C}(\overline{\Omega}_{2})\in(n)$. We set $l_{0}=0$, $\Omega_{1}^{0}=\overline{\Omega}_{1}$ and $\Omega_{2}^{0}=\overline{\Omega}_{2}$. Applying Lemma 3.5, with $\beta_{l+1}=\beta_{0}$, we obtain $\Omega_{1}^{1}=\Omega_{1}^{0}$ and $\Omega_{2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{0}),<}^{0}$ where $\nu_{\mathcal{C}}(\Omega_{2}^{1})=n+{\mathfrak{n}}(\beta_{0})$. The family ${\mathcal{G}}_{0}^{1}=\{\Omega_{1}^{1},\Omega_{2}^{1}\}$ is an $E$-good family of level 00. Moreover, it satisfies $\nu(\Omega_{1}^{1})=\delta_{x}$ and $\nu(\Omega_{2}^{1})=1$. Remark 4.1 guarantees that we can find an $E$-nice basis starting with $\mathcal{G}_{0}^{1}$.
4.2.3. Case $\delta_{y}=1$ and $\beta_{0}\not\in(n)$
Notice that $\beta_{0}=\beta_{1}$ in this case. Recall Definition 4.1 for $\overline{\Omega}_{1},\overline{\Omega}_{2}$ and $\overline{\Omega}_{j}$. As $\overline{\Omega}_{1}=x^{\delta_{x}}dy$ and $\overline{\Omega}_{2}=ydx$, we have $[\nu_{C}(\overline{\Omega}_{1})]_{n}=[\nu_{C}(\overline{\Omega}_{2})]_{n}=[\beta_{1}]_{n}$, so that $\{\overline{\Omega}_{1},\overline{\Omega}_{2}\}$ is not an $E$-separated family of level 00. Using Lemma 3.5 and Proposition 3.3 we obtain $\Omega_{1}^{1}=\overline{\Omega}_{1}$ and $\Omega_{2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{1}),<}^{0}$ with $\nu_{\mathcal{C}}(\Omega_{2}^{1})=n+{\mathfrak{n}}(\beta_{1})$. This family ${\mathcal{G}}_{0}^{1}=\{\Omega_{1}^{1},\Omega_{2}^{1}\}$ is $E$-separated of level 00 but does not satisfy $E$-(b) or $E$-(c) in Definition 4.4, so it is not an $E$-terminal family of level 00. This is why, in this case, we need to start at $l_{0}=1$. By Remark 4.1, we just need to provide an $E$-good family of level $1$.
Consider the $E$-ordered family ${\mathcal{G}}_{1}^{0}=\{\Omega_{1}^{0},\ldots,\Omega_{2{\nu}_{1}}^{0}\}$ given by $\Omega_{j}^{0}=\overline{\Omega}_{j}$ for $1\leq j\leq 2{\nu}_{1}$, which satisfies $\nu(\Omega_{j}^{0})=\lfloor\frac{j+\delta_{x}}{2}\rfloor$ for $1\leq j\leq 2{\nu}_{1}$. Note that
(23) $[\nu_{C}(\overline{\Omega}_{2j+1})]_{n}=[\nu_{C}(\overline{\Omega}_{2j+2})]_{n}=[(j+1)\beta_{1}]_{n}.$
if $1\leq 2j+1<2j+2\leq 2{\nu}_{1}$. All congruences $[\beta_{1}]_{n},[2\beta_{1}]_{n},\ldots,[\nu_{1}\beta_{1}]_{n}$ are pairwise different, so the unique situation in which two $1$-forms have congruent ${\mathcal{C}}$-orders is described in equation (23). Since $dy\in\tilde{\Delta}_{\beta_{1},<}^{0}$, we have $\Omega_{2j+1}^{0}\in\tilde{\Delta}_{\beta_{1},<}$ by Corollary 3.3. We also have $\Omega_{2j+2}^{0}\in\tilde{\Delta}_{\beta_{1},\leq}$ by Corollary 3.2. We apply Lemma 3.5 and Proposition 3.3 to obtain
$\Omega_{2j+1}^{1}=\Omega_{2j+1}^{0}\in\Delta_{\beta_{1},<}\quad\mathrm{and}\quad\nu_{\mathcal{C}}(\Omega_{2j+1}^{1})=n\delta_{x}+(j+1)\beta_{1}$
for $2j+1\leq 2{\nu}_{1}$; we also get
$\Omega_{2j+2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{1}),<}^{0},\ \ \nu(\Omega_{2j+2}^{1})=j+1\ \ \mathrm{and}\ \ \nu_{\mathcal{C}}(\Omega_{2j+2}^{1})=n+j\beta_{1}+{\mathfrak{n}}(\beta_{1})$
for $2j+2\leq 2{\nu}_{1}$. As a consequence, the family ${\mathcal{G}}_{1}^{1}=\{\Omega_{1}^{1},\ldots,\Omega_{2{\nu}_{1}}^{1}\}$ is $E$-good of level $1$ (with $E$-(I) being empty). Note that ${\mathcal{L}}_{1}$ is the subset of odd numbers of $\{1,\ldots,2\nu_{1}\}$ whereas ${\mathcal{U}}_{1}$ is its subset of even numbers. Moreover, we obtain $\nu(\Omega_{j}^{1})=\lfloor\frac{j+\delta_{x}}{2}\rfloor$ for $1\leq j\leq 2{\nu}_{1}$ by construction. As stated above, Remark 4.1 ends our argument.
4.3. Existence of a nice $E$-basis
The previous discussion provides the proof for the next result.
Theorem 4.1.
In any case, starting with the $E$-initial family of level $l_{0}\in\left\{0,1\right\}$, the procedure of section 3.5 provides an $E$-terminal family and hence a nice $E$-basis $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ for the equisingularity class ${\mathcal{C}}$ of $(\Gamma,E)$.
Remark 4.2.
The analogue of Proposition 3.7 holds in the general case, i.e.
$\nu_{\mathcal{C}}(\Omega_{k})=\max\{\nu_{\mathcal{C}}(\Omega^{\prime}):\Omega^{\prime}\in\hat{M}_{k}^{E}\}.$
The proof is the same.
Remark 4.3.
The set $\Lambda_{\mathcal{C}}^{E}$ (Definition 2.7) depends on $n,\beta_{0},\beta_{1},\ldots,\beta_{g}$ and indeed it is straightforward to write an algorithm whose input is $(\delta_{x},\delta_{y},n,\beta_{0},\ldots,\beta_{g})$ and whose output is $\Lambda_{\mathcal{C}}^{E}$. The argument is analogous to the one in Remark 3.18.
5. Families of equisingular curves
In order to reason using the resolution of singularities of a generic curve in an equisingularity class, we need to study the relation between an analytic family of curves and another family in the blow-up of the generic element. It turns out that these families are intimately related and a nice basis for the former will essentially provide a nice basis for the latter. The study of this relation is the purpose of this section.
5.1. Suitable pairs
Fix a local system of coordinates $(x,y)$ in $(\mathbb{C}^{2},0)$. From now on, we shall use the expression pair to refer to a pair $(\Gamma,E)$ where $\Gamma$ is a germ of irreducible curve at $0\in\mathbb{C}^{2}$ (or at any point in a non-singular complex analytic surface) and $E$ is a germ of normal crossings divisor.
Definition 5.1.
Given a pair $(\Gamma,E)$, let $\pi_{1}$ be the blow-up of the origin $P_{0}=(0,0)$ of $\mathbb{C}^{2}$ and let $E_{1}=\pi_{1}^{-1}(E)$ and $P_{1}$ be the point in the strict transform $\Gamma_{1}$ of $\Gamma$ that belongs to the divisor $\pi_{1}^{-1}(0,0)$. We consider ${\Gamma}_{1}$ as a germ of irreducible curve defined in a neighborhood of ${P}_{1}$.
The following is the key notion which will enable us to relate a nice basis before blow-up with another one after.
Definition 5.2.
We say that the pair $(\Gamma,E)$ is suitable if
  • $E\subset\{xy=0\}$,
  • $x=0$ is not the tangent cone of $\Gamma$ at the origin and
  • if $E=\{y=0\}$ then $y=0$ is the tangent cone of $\Gamma$ at the origin.
A pair $(\Gamma,E)$ which is not suitable will be called unsuitable.
Remark 5.1.
Let ${\mathcal{C}}$ be the equisingularity class of a pair $(\Gamma,E)$ (with $E\subset\{xy=0\}$). The property of being suitable only depends on the order of the coordinates $(x,y)$. Indeed, $(\Gamma,E)$ is suitable in coordinates $({\rm x}_{0},{\rm y}_{0})$, where $({\rm x}_{0},{\rm y}_{0})=(x,y)$ if $(\Gamma,E)$ is suitable in coordinates $(x,y)$; otherwise, it is suitable in coordinates $({\rm x}_{0},{\rm y}_{0})=(y,x)$.
Consider the equisingularity class ${\mathcal{C}}$ of a suitable pair $(\Gamma_{\circ},E)$ in coordinates $({\rm x}_{0},{\rm y}_{0})=(x,y)$. Any curve $\Gamma\in\mathcal{C}$ has a Puiseux parametrization of the form
(24) $\Gamma(t)=\left(t^{n},\sum_{\beta\geq\beta_{0}}a_{\beta,\Gamma}t^{\beta}\right)$
where $n=\nu(\Gamma)$ depends just on ${\mathcal{C}}$ and so we can write $n=\nu({\mathcal{C}})$, $a_{\beta_{0},\Gamma}\neq 0$ and $\beta_{0}\geq n$. That parametrization will be called a suitable parametrization of $\Gamma$, leaving $E$ implicit. Consider coordinates $x=x_{1}$, $y=x_{1}y_{1}$ in the first chart of the blow-up $\pi_{1}$ of the origin. Since $x=0$ is not the tangent cone of $\Gamma$, its strict transform ${\Gamma}_{1}$ has a parametrization
(25) ${\Gamma}_{1}(t)=\left(t^{n},\sum_{\beta\geq\beta_{0}}a_{\beta,\Gamma}t^{\beta-n}\right)$
in coordinates $(x_{1},y_{1})$. We now relate the main invariants of $(\Gamma,E)$ to those of $(\Gamma_{1},\pi_{1}^{-1}(E))$ in a neighborhood of $P_{1}$.
5.1.1. Approximate roots
Take $1\leq j\leq g$. Given an approximate root
$f_{j,\Gamma}=y^{\nu_{j-1}}+a_{\nu_{j-1}-1}(x)y^{\nu_{j-1}-1}+\ldots+a_{0}(x)$
of $\Gamma\in\mathcal{C}$, the $j$th-approximate root ${f}_{1,j,\Gamma_{1}}$ of ${\Gamma}_{1}$ in the local system of coordinates above satisfies
${f}_{1,j,\Gamma_{1}}=\frac{f_{j,\Gamma}\circ\pi_{1}}{x_{1}^{\nu_{j-1}}}.$
Let ${\mathcal{C}}_{1}$ be the equisingularity class of $({\Gamma}_{1},E_{1})$. Since by construction, $\nu(\Gamma_{1})$ is the multiplicity of $\Gamma_{1}$ at $P_{1}$, and on the other hand the values $\nu_{\Gamma_{1}}(f_{1,j,\Gamma_{1}})$ depend just on ${\mathcal{C}}_{1}$, we shall denote them $\nu({\mathcal{C}}_{1})$ and $\nu_{{\mathcal{C}}_{1}}({f}_{1,j})$ respectively, for simplicity. Certainly,
$\nu_{{\mathcal{C}}_{1}}({f}_{1,j})=\nu_{\mathcal{C}}(f_{j})-\nu_{j-1}n=\overline{\beta}_{j}-\nu_{j-1}n$
holds. Moreover, for any $1\leq j\leq g$, the $y$-degree of both $f_{j}$ and $f_{1,j}$ is equal to $\nu_{j-1}$, and their leading variables are $\beta_{j}$, where we are considering coordinates $(x,y)$ and $(x_{1},y_{1})$ and hence the parametrizations (24) and (25).
5.1.2. Preparing the pair $({\Gamma}_{1},{E}_{1})$
In this subsection we find coordinates $({\rm x}_{1},{\rm y}_{1})$ in which $\Gamma_{1}$ has a suitable parametrization that we relate with the parametrization (24) of $\Gamma$.
The pair $({\Gamma}_{1},{E}_{1})$ is not necessary suitable in coordinates $(x_{1},y_{1})$; for instance, if $a_{n}\neq 0$ then $(\Gamma_{1},E_{1})$ is not even a germ of curve at $(x_{1},y_{1})=(0,0)$. This is easily solved by means of the change of coordinates $(x_{1},y_{1}^{\prime})=(x_{1},y_{1}-a_{n})$. Thus,
(26) $\left(t^{n},\sum_{\beta\geq\beta^{\prime}}a_{\beta,\Gamma}t^{\beta-n}\right)$
is a parametrization of ${\Gamma}_{1}$ in coordinates $(x_{1},y_{1}^{\prime})$, where $\beta^{\prime}$ is the first index, greater than $n$, such that $a_{\beta^{\prime},\Gamma}\neq 0$. Now $({\Gamma}_{1},{E}_{1})$ is suitable if and only if $\beta^{\prime}-n\geq n$. If this is the case, then we define coordinates $({\rm x}_{1},{\rm y}_{1})=(x_{1},y_{1}^{\prime})$ and (26) provides a suitable parametrization
$\left(t^{n},\sum_{\beta\geq\beta^{\prime}-n}b_{\beta,\Gamma_{1}}t^{\beta}\right)=\left(t^{\nu({\mathcal{C}}_{1})},\sum_{\beta\geq\beta^{\prime}-n}b_{\beta,\Gamma_{1}}t^{\beta}\right),$
in coordinates $({\rm x}_{1},{\rm y}_{1})$, where
(27) $b_{\beta-\nu({\mathcal{C}}_{1}),\Gamma_{1}}=b_{\beta-n,\Gamma_{1}}=a_{\beta,\Gamma}$
for any $\beta\in{\mathcal{E}}_{E}(\mathcal{C})\setminus\{n\}$.
Remark 5.2.
Notice that $({\Gamma}_{1},{E}_{1})$ is suitable, in coordinates $(x_{1},y_{1}^{\prime})$, either for all $\Gamma\in\mathcal{C}$ or for none. If there exists an exponent in ${\mathcal{E}}_{E}(\mathcal{C})$ greater than $n$ and less than $2n$ then $\beta^{\prime}$ is equal to the first Puiseux exponent $\beta_{1}$. Since $a_{\beta_{1}}$ never vanishes on $\mathcal{C}$, it follows that $({\Gamma}_{1},{E}_{1})$ is unsuitable for any $\Gamma\in\mathcal{C}$. If no such a exponent exists, clearly $({\Gamma}_{1},{E}_{1})$ is suitable for any $\Gamma\in\mathcal{C}$.
Assume that $({\Gamma}_{1},{E}_{1})$ is unsuitable in coordinates $(x_{1},y_{1}^{\prime})$ for $\Gamma\in\mathcal{C}$. The pair $({\Gamma}_{1},{E}_{1})$ is suitable in coordinates $({\rm x}_{1},{\rm y}_{1})=(y_{1}^{\prime},x_{1})$ by Remark 5.1. Let us provide a suitable parametrization of $\Gamma_{1}$ in coordinates $({\rm x}_{1},{\rm y}_{1})$. Notice that
$\left(\sum_{\beta\geq\beta_{1}}a_{\beta,\Gamma}t^{\beta-n},t^{n}\right)$
is a (possibly unsuitable) parametrization of ${\Gamma}_{1}$ in coordinates $({\rm x}_{1},{\rm y}_{1})$. From this one, we can obtain, by taking $(\beta_{1}-n)$-th roots of unity, the suitable parametrization
(28) $\left(t^{\beta_{1}-n},\sum_{\beta\geq n}b_{\beta,\Gamma_{1}}t^{\beta}\right)=\left(t^{\nu({\mathcal{C}}_{1})},\sum_{\beta\geq n}b_{\beta,\Gamma_{1}}t^{\beta}\right).$
with the important properties:
(29) $b_{n,\Gamma_{1}}=a_{\beta_{1},\Gamma}^{-\frac{n}{\beta_{1}-n}}$
and
(30) $b_{\beta-\nu({\Gamma}_{1}),\Gamma_{1}}=-\frac{n}{\beta_{1}-n}a_{\beta_{1},\Gamma}^{-\frac{\beta-n}{\beta_{1}-n}}a_{\beta,\Gamma}+Q_{\beta}(a_{\beta_{1},\Gamma}^{\frac{-1}{\beta_{1}-n}},a_{{\mathfrak{n}}(\beta_{1}),\Gamma},\ldots,a_{{\mathfrak{p}}(\beta),\Gamma}),$
where for any $\beta\in{\mathcal{E}}_{E}(\mathcal{C})$ with $\beta>\beta_{1}$, $Q_{\beta}$ is a polynomial. In particular $b_{\beta-\nu({\Gamma}_{1})}$ depends on $a_{\beta_{1}},a_{{\mathfrak{n}}(\beta_{1})},\ldots,a_{\beta}$ and in degree $1$ on $a_{\beta}$.
Remark 5.3.
Let $\Gamma\in\mathcal{C}$, and consider a suitable pair $(\Gamma,E)$ in coordinates $({\rm x}_{0},{\rm y}_{0})=(x,y)$. Let
(31) $(\mathbb{C}^{2},0)\stackrel{{\scriptstyle\pi_{1}}}{{\longleftarrow}}\mathcal{X}_{1}\stackrel{{\scriptstyle\pi_{2}}}{{\longleftarrow}}\cdots\mathcal{X}_{k-1}\stackrel{{\scriptstyle\pi_{k}}}{{\longleftarrow}}\mathcal{X}_{k}\stackrel{{\scriptstyle\pi_{k+1}}}{{\longleftarrow}}\cdots$
be the local blowing-up process following the infinitely near points of $\Gamma=\Gamma_{0}$, and denote by $\Gamma_{i}$ the strict transform of $\Gamma$ in $\mathcal{X}_{i}$. Let $E_{0}=E$ and $E_{i}=\pi_{i}^{-1}\circ\cdots\circ\pi_{1}^{-1}(E)$ be the (germ of the) exceptional divisor in each stage $i\geq 1$. Denote $P_{0}=(0,0)\in\mathbb{C}^{2}$ and let $P_{i}$ be the point of $\Gamma_{i}$ in $E_{i}$. We find coordinates $({\rm x}_{s},{\rm y}_{s})$ in which the pair $(\Gamma_{s},E_{s})$ is suitable for $s\geq 1$ by just iterating the construction that was described for $s=1$. For any germ of curve $\Gamma^{\prime}$ at $P_{s}$ such that $(\Gamma^{\prime},E_{s})$ is in the equisingularity class ${\mathcal{C}}_{s}$ of the pair $(\Gamma_{s},E_{s})$ there is a suitable parametrization of $\Gamma^{\prime}$ in coordinates $({\rm x}_{s},{\rm y}_{s})$.
In the remainder of this section, we fix the coordinates $({\rm x}_{s},{\rm y}_{s})$ and suitable parametrizations of elements in ${\mathcal{C}}_{s}$ are considered in such coordinates for $s\geq 0$. Given $\Gamma_{s}^{\prime}\in{\mathcal{C}}_{s}$, the suitable parametrization of the strict transform $\Gamma_{s+1}^{\prime}$ of $\Gamma_{s}^{\prime}$ by $\pi_{s+1}$ is obtained from the suitable parametrization of $\Gamma_{s}^{\prime}$ by either equation (27) or equations (29), (30).
5.1.3. Fanning exponents
Consider now a one-parameter analytic family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of curves in ${\mathcal{C}}$: that is, a family $\Gamma^{\epsilon}$ admitting parametrizations
$\Gamma^{\epsilon}(t)=\left(t^{n},\sum_{\beta\geq\beta_{0}}a_{\beta}(\epsilon)t^{\beta}\right)$
where $a_{\beta_{0}},a_{{\mathfrak{n}}(\beta_{0})},\ldots$ are analytic in a common neighborhood of $\epsilon=0$. Suppose that the family is non-constant and let
$\theta(\Gamma^{\epsilon})=\min\{\beta\in{\mathcal{E}}_{E}(\mathcal{C}):a_{\beta}\ \mathrm{is}\ \mathrm{non-constant}\}$
be the fanning exponent of the family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$. Given such a family, we obtain by pull-back a corresponding family $(\Gamma^{\epsilon}_{1},E_{1})$ in $P_{1}$ with suitable parametrizations described above. The next remarks follow from the previous discussions.
Remark 5.4.
[12, Lemma 3.19] There are two possibilities:
  • Either $\theta(\Gamma^{\epsilon})=n$ and the map $\epsilon\mapsto{\Gamma}_{1}^{\epsilon}\cap\pi_{1}^{-1}(0)$ is non-constant: in other words the base point of ${\Gamma}_{1}^{\epsilon}$ in $\pi_{1}^{-1}(0)$ varies in the family,
  • Or $\theta(\Gamma^{\epsilon})>n$ and the first non-constant coefficient of the suitable parametrization of the family $({\Gamma}_{1}^{\epsilon},{E}_{1})$ is $b_{\theta(\Gamma^{\epsilon})-\nu({\mathcal{C}}_{1})}(\epsilon)$.
Remark 5.5.
Fix a parametrization $(t^{\nu({\mathcal{C}})},\sum a_{\beta,{\Gamma}}t^{\beta})$ of $\Gamma\in{\mathcal{C}}$. Given $\beta^{\prime}\in{\mathcal{E}}_{E}({\mathcal{C}})$, consider the family $({\Gamma}_{\beta^{\prime}}^{\epsilon})_{\epsilon\in(\mathbb{C},0)}$ of curves in ${\mathcal{C}}$ such that $(t^{\nu({\mathcal{C}})},\epsilon t^{\beta^{\prime}}+\sum a_{\beta,{\Gamma}}t^{\beta})$ is a parametrization of $\Gamma_{\beta^{\prime}}^{\epsilon}$. The family of strict transforms $({\Gamma}_{1,\beta^{\prime}}^{\epsilon})_{\epsilon\in(\mathbb{C},0)}$ is a family of curves in ${\mathcal{C}}_{1}$ as long as $\beta^{\prime}>\nu({\mathcal{C}})$. Remark 5.4 implies also that if $\beta^{\prime}>\nu({\mathcal{C}})$, then $\beta^{\prime}-\nu({\mathcal{C}}_{1})\in{\mathcal{E}}_{E_{1}}({\mathcal{C}}_{1})$.
Remark 5.6.
Take, on the other hand, a suitable curve $\Gamma_{1}$ such that $\pi_{1}(\Gamma_{1})=\Gamma\in\mathcal{C}$, with suitable parametrization $(t^{\nu({\mathcal{C}}_{1})},\sum b_{\beta,{\Gamma}_{1}}t^{\beta})$ in coordinates $({\rm x}_{1},{\rm y}_{1})$. Given $\beta^{\prime}\in{\mathcal{E}}_{{E}_{1}}({\mathcal{C}}_{1})$, consider the family $({\Gamma}_{1,\beta^{\prime}}^{\epsilon})_{\epsilon\in(\mathbb{C},0)}$ of curves in ${\mathcal{C}}$ such that $(t^{\nu({\mathcal{C}}_{1})},\epsilon t^{\beta^{\prime}}+\sum b_{\beta,{\Gamma}_{1}}t^{\beta})$ is a parametrization of ${\Gamma}_{1,\beta^{\prime}}^{\epsilon}$. This defines by blow-down a family $({\Gamma}^{\epsilon})_{\epsilon\in(\mathbb{C},0)}$ of equisingular curves such that $\Gamma^{0}=\Gamma$. As a consequence, the fanning coefficient of $(\Gamma^{\epsilon})$ satisfies $\theta(\Gamma^{\epsilon})\in{\mathcal{E}}_{E}(\mathcal{C})\setminus\{\nu({\mathcal{C}})\}$. Even more, $\theta(\Gamma^{\epsilon})=\beta^{\prime}+\nu({\mathcal{C}}_{1})$ by Remark 5.4.
Remark 5.7.
Remarks 5.5 and 5.6 imply that the map
$\begin{array}[]{ccc}{\mathcal{E}}_{E}(\mathcal{C})\setminus\{\nu({\mathcal{C}})\}&\to&{\mathcal{E}}_{{E}_{1}}({\mathcal{C}}_{1})\\ \theta&\mapsto&\theta-\nu({\mathcal{C}}_{1})\end{array}$
is a bijection.
Remark 5.8.
Let ${\mathcal{C}}$ be the equisingularity class of a suitable pair $(\Gamma,E)$ with $\Gamma$ generic in $\mathcal{C}$. The strict transform ${\Gamma}_{1}$ is a generic curve of ${\mathcal{C}}_{1}$. This holds because $b_{\beta-\nu({\mathcal{C}}_{1})}$ depends on the all coefficients $(a_{\beta^{\prime}})_{\nu({\mathcal{C}})<\beta^{\prime}\leq\beta}$ and in degree $1$ on $a_{\beta}$ for any $\beta\in{\mathcal{E}}_{E}(\mathcal{C})\setminus\{\nu({\mathcal{C}})\}$ except when $\beta=n+\nu({\mathcal{C}}_{1})$ and $({\Gamma}_{1},{E}_{1})$ is unsuitable in coordinates $(x_{1},y_{1}^{\prime})$. In this case, $b_{n}$ is a power of the linear function $a_{\beta_{1}}=a_{n+\nu({\mathcal{C}}_{1})}$. This dependence on the highest index variable $a_{\beta}$ is preserved by further blow-ups, as is the genericity of the strict transforms.
5.2. Families of equisingular curves and desingularization
Consider a curve $\Gamma\in\mathcal{C}$, a suitable pair $(\Gamma,E)$ and the setting in Remark 5.3. Let $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ be a a non-constant analytic family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of curves with $\Gamma^{0}=\Gamma$ with suitable parametrizations of the form $\Gamma^{\epsilon}(t)=(t^{\nu({\mathcal{C}})},\sum_{\beta\geq\beta_{0}}a_{\beta}(\epsilon)t^{\beta})$. Define $(\Gamma_{0}^{\epsilon})=(\Gamma^{\epsilon})$ and $(\Gamma_{i}^{\epsilon})$ the strict transform of $(\Gamma^{\epsilon})$ in $\mathcal{X}_{i}$. As $(\Gamma^{\epsilon})$ is non-constant, there exists an integer $\iota(\Gamma^{\epsilon})\geq 1$ such that $\Gamma_{i}^{\epsilon}$ passes through $P_{i}$ for all $0\leq i<\iota(\Gamma^{\epsilon})$, and $\epsilon$ in a neighborhood of 00 but $\Gamma_{\iota}^{\epsilon}$ does not pass through $P_{\iota(\Gamma^{\epsilon})}$ for any $\epsilon$ in a neighborhood of 00.
Definition 5.3.
We say that $\iota(\Gamma^{\epsilon})$ is the sliding divisor of the family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$.
Denote $D_{i}=\pi_{i}^{-1}(P_{i-1})$ for $i\geq 1$. The sliding divisor $\iota(\Gamma^{\epsilon})$ indicates the divisor $D_{\iota(\Gamma^{\epsilon})}$ where the curves in the family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ have no common point, hence the name (the point $\Gamma_{\iota(\Gamma^{\epsilon})}^{\epsilon}\cap D_{\iota(\Gamma^{\epsilon})}$ slides along that divisor).
Remark 5.9.
Let $\iota=\iota(\Gamma^{\epsilon})$ for clarity in this remark. Since the family $(\Gamma_{\iota-1}^{\epsilon},E_{\iota-1})_{\epsilon\in({\mathbb{C}},0)}$ is contained in an equisingularity class and their points $P_{\iota,\epsilon}=\Gamma_{\iota}^{\epsilon}\cap E_{\iota}$ vary continuously in a non-constant way, there is exactly one irreducible component of $E_{\iota}$ passing through $P_{\iota}$, so that $P_{\iota}$ is necessarily a trace point of $E_{\iota}$.
Remark 5.10.
The sliding divisor $\iota(\Gamma_{\epsilon})$ is equal to $1$ if and only if the fanning exponent $\theta(\Gamma^{\epsilon})$ is equal to $\nu({\mathcal{C}})$, by Remark 5.4.
We now proceed to show how the sliding divisor determines the fanning exponent. Let $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ be an analytic family in $\mathcal{C}$ and assume the notations above for $(\Gamma^{\epsilon}_{i})$, etc. Set $\Gamma_{0}^{\epsilon}=\Gamma^{\epsilon}$ for simplicity. The following definition is closely related to Corollary 3.20 of [12].
Definition 5.4.
Let ${\mathcal{C}}$ be the equisingularity class of a suitable pair $(\Gamma^{0},E)$. We define $\mathbf{Slid}_{\mathcal{C}}$ as the set of sliding divisors associated to analytic families $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ in ${\mathcal{C}}$. We define the map
$\hat{\theta}(\iota)=\nu_{P_{\iota-1}}(\Gamma_{\iota-1}^{0})+\sum_{j=1}^{\iota-1}\nu_{P_{j}}(\Gamma_{j}^{0}).$
for $\iota\in\mathbf{Slid}_{\mathcal{C}}$ (cf. Definition 2.1).
Lemma 5.1.
Let $\iota=\iota(\Gamma^{\epsilon})$ be the sliding divisor of $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$. Then
$\theta(\Gamma^{\epsilon}):=\theta(\Gamma_{0}^{\epsilon})=\hat{\theta}(\iota)$
where an empty sum is 00.
Proof.
The equality is trivial if $\iota=1$ by Remark 5.4, so we can assume $\iota>1$. Since the sliding divisor of the family $(\Gamma_{\iota-1}^{\epsilon})$ is equal to one, its fanning exponent is the multiplicity $\nu_{P_{\iota-1}}(\Gamma_{\iota-1}^{0})$ by Remark 5.10. The result follows now applying Remark 5.4 to the families $(\Gamma_{\iota-2}^{\epsilon})_{\epsilon\in({\mathbb{C}},0)},\ldots,(\Gamma_{0}^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$. ∎
Before proceeding, we study which exceptional divisors in the desingularization of a curve are candidates to sliding divisors:
Lemma 5.2.
Let ${\mathcal{C}}$ be the equisingularity class of a suitable pair $(\Gamma,E)$. Then
$\mathbf{Slid}_{\mathcal{C}}=\{\iota\geq 1:P_{\iota}\ \mbox{\emph{is a trace point of}}\ E_{\iota}\}.$
Proof.
Let us recall that $P_{i}$ is the $i$th-infinitely near point of $\Gamma$. Notice that whether or not $P_{i}$ is a trace point of $E_{i}$ does not depend on the choice of $\Gamma\in{\mathcal{C}}$. By Remark 5.9, if $\iota$ is a sliding divisor then $P_{\iota}$ is a trace point of $E_{\iota}$, which gives the direct inclusion.
To prove the reverse inclusion, assume that $P_{\iota}$ is a trace point of $E_{\iota}$ and let $\mu_{\iota-1}=\nu_{P_{\iota-1}}(\Gamma_{\iota-1})$. Consider a suitable parametrization $(t^{\mu_{\iota-1}},\sum_{\beta\geq\mu_{\iota-1}}a_{\beta,\Gamma}t^{\beta})$ of $(\Gamma_{\iota-1},E_{\iota-1})$ in coordinates $({\rm x}_{\iota-1},{\rm y}_{\iota-1})$, and define the analytic family $(\Gamma_{\iota-1}^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ with $\Gamma_{\iota-1}^{\epsilon}$ defined in the neighborhood of $P_{\iota-1}$ having suitable parametrization
$\Gamma^{\epsilon}_{\iota-1}\equiv\bigg{(}t^{\mu_{\iota-1}},\epsilon t^{\mu_{\iota-1}}+\sum_{\beta\geq\mu_{\iota-1}}a_{\beta,\Gamma^{0}}t^{\beta}\bigg{)}.$
Obviously, $\Gamma_{\iota-1}^{0}=\Gamma_{\iota-1}$. Since $P_{\iota}$ is a trace point, the curves in the family belong to the equisingularity class of $(\Gamma_{\iota-1},E_{\iota-1})$. The family $(\Gamma_{\iota-1}^{\epsilon})$ induces by blow-down another family $(\Gamma_{0}^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ defined in a neighborhood of the origin, and each curve $\Gamma_{0}^{\epsilon}$ belongs to $\mathcal{C}$. By construction, $\iota$ is the sliding divisor of such a family, and we are done. ∎
Lemma 5.3.
Let ${\mathcal{C}}$ be the equisingularity class of a suitable pair $(\Gamma,E)$. Consider $\iota,\iota^{\prime}\in\mathbf{Slid}_{\mathcal{C}}$ with $\iota<\iota^{\prime}$. Then $\hat{\theta}(\iota)<\hat{\theta}(\iota^{\prime})$.
Proof.
Denote $\mu_{j}=\nu_{P_{j}}(\Gamma_{j})$ for $j\geq 0$ and consider the family $(\Gamma_{0}^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ associated to $\iota^{\prime}$ in the proof of Lemma 5.2 with $\Gamma^{0}_{0}=\Gamma$. Since the sliding divisor $\iota(\Gamma_{\iota-1}^{\epsilon})$ is greater than $1$ by definition, we have, by Remark 5.4:
$\theta(\Gamma_{\iota-1}^{\epsilon})\in{\mathcal{E}}_{E_{\iota-1}}(\Gamma_{\iota-1}^{0})\setminus\{\mu_{\iota-1}\}$
and, in particular, $\theta(\Gamma^{\epsilon}_{\iota-1})>\mu_{\iota-1}$. By applying Remark 5.4 $\iota-1$ times, we conclude that
$\hat{\theta}(\iota^{\prime})>\mu_{\iota-1}+\sum_{j=1}^{\iota-1}\mu_{j}$
Hence, $\hat{\theta}(\iota)<\hat{\theta}(\iota^{\prime})$ by Lemma 5.1. ∎
Proposition 5.1.
Let ${\mathcal{C}}$ be the equisingularity class of a suitable pair $(\Gamma,E)$. Then $\hat{\theta}:\mathbf{Slid}_{\mathcal{C}}\to{\mathcal{E}}_{E}(\mathcal{C})$ is a monotone bijection.
Proof.
By Lemma 5.3, it suffices to prove that $\hat{\theta}$ is surjective. This is clear since, by Remark 5.5, any $k\in{\mathcal{E}}_{E}(\mathcal{C})$ is the fanning exponent of a family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ in ${\mathcal{C}}$. ∎
5.3. Contact orders between curves and vector fields
Our study of the analytic moduli requires us to restrict the analytic families we have studied to be associated to flows of vector fields. This is what we introduce in this subsection.
Consider the equisingularity class ${\mathcal{C}}$ of a suitable pair $(\Gamma,E)$. Recall that $P_{0},P_{1},\ldots$ is the sequence of infinitely near points of $\Gamma$ following the blow-ups $\pi_{i}$ for $i\geq 1$.
Definition 5.5.
For any integer $i\geq 0$, we denote by ${\mathfrak{X}}_{E,i}$ the set of analytic vector fields defined in a neighborhood of $P_{i}$ in $\mathcal{X}_{i}$ that
  • leave invariant every irreducible component of $E_{i}$ at $P_{i}$ (see subsection 5.2 for notation), and
  • are singular at $P_{i}$ for $i=0$.
We denote by ${\mathfrak{X}}_{E,i}^{\circ}$ the subset of ${\mathfrak{X}}_{E,i}$ consisting of vector fields singular at $P_{i}$.
Given $X\in{\mathfrak{X}}_{E,i}$, with $X(P_{i})=0$ or $i\geq 1$, such that $\Gamma_{i}$ is not $X$-invariant, we consider the flow $(\phi^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of $X$ in a neighborhood of $P_{i}$ and define the curve $\Gamma_{i}^{\epsilon}=\phi^{\epsilon}(\Gamma_{i})$ for $\epsilon\in\mathbb{C}$ in a neighborhood of 00. The pair $(\Gamma_{i}^{\epsilon},E_{i})$ is in the same equisingularity class as $(\Gamma_{i},E_{i})$ for $\epsilon$ in a neighborhood of 00. By blowing-down we obtain the family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ induced by $X$, which is an analytic family in the equisingularity class of $\Gamma\in{\mathcal{C}}$.
Notation. The fanning exponent of the family $(\Gamma^{\epsilon})$ just constructed will be denoted $\theta_{X}(\Gamma)$ in order to stress that $\Gamma^{\epsilon}$ is defined by $X$.
Definition 5.6.
For any $i\geq 0$, we define the sets of fanning exponents associated to analytic vector fields and to singular analytic vector fields:
$\begin{split}&\Theta_{i+1}(\Gamma)=\{\theta_{X}(\Gamma):X\in{\mathfrak{X}}_{E,i+1}\},\\ &\Theta_{i}^{\circ}(\Gamma)=\{\theta_{X}(\Gamma):X\in{\mathfrak{X}}_{E,i}^{\circ}\}.\end{split}$
Both are included in ${\mathcal{E}}_{E}({\mathcal{C}})$. For completeness, we set $\Theta_{0}(\Gamma)=\Theta_{0}^{\circ}(\Gamma)$. We shall call them the sets of flow-fanning exponents and the singular flow-fanning exponents at indices $i+1$ and $i$, respectively.
Remark 5.11.
[12, Corollary 4.11 and Lemma 4.13] When $E=\emptyset$, the number $\sharp({\mathcal{E}}_{\emptyset}({\mathcal{C}})\setminus\Theta_{0}^{\circ}(\Gamma))$ is the dimension of the space of analytic classes of elements in ${\mathcal{C}}$ for any generic $\Gamma\in{\mathcal{C}}$. Moreover, we have
$\Theta_{0}^{\circ}(\Gamma)=\{\nu_{\Gamma}(\omega)-\nu({\mathcal{C}})\ |\ w\in\hat{\Omega}(\mathbb{C}^{2},0)\ \mathrm{and}\ \omega(0)=0\}$
by [12, Theorem 3.21].
The following long result is the foundation of the enumeration of the elements in $\Lambda_{\Gamma}^{E}$ which will be done in the next section.
Proposition 5.2.
Consider the equisingularity class ${\mathcal{C}}$ of a suitable pair $(\Gamma,E)$. Then:
  1. (1)
    Singular flow-fanning exponents at index $i$ are flow-fanning exponents at index $i+1$: $\Theta_{i}^{\circ}(\Gamma)\subset\Theta_{i+1}(\Gamma)$ for $i\geq 0$.
  2. (2)
    For indices not corresponding to sliding divisors, flow-fanning exponents and singular flow-fanning exponents are the same: $\Theta_{i}(\Gamma)=\Theta_{i}^{\circ}(\Gamma)$ if $i\in\mathbb{Z}_{\geq 1}\setminus\mathbf{Slid}_{\mathcal{C}}$.
  3. (3)
    Assume $i$ is a sliding divisor. Then flow-fanning exponents of index $i$ are greater than or equal to $\hat{\theta}(i)$. Moreover, $\hat{\theta}(i)$ is the unique non-singular flow-fanning exponent of index $i$: if $i\in\mathbf{Slid}_{\mathcal{C}}$, then $\Theta_{i}(\Gamma)\subset\mathbb{Z}_{\geq\hat{\theta}(i)}$ and $\Theta_{i}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)=\{\hat{\theta}(i)\}$.
  4. (4)
    For any sliding divisor $i$ and $j>i$, the flow-fanning exponents of index $j$ are strictly greater than $\hat{\theta}(i)$: $\Theta_{j}(\Gamma)\subset\mathbb{Z}_{>\hat{\theta}(i)}$ if $i\in\mathbf{Slid}_{\mathcal{C}}$ and $i<j$.
  5. (5)
    Flow-fanning exponents of index $i+1$ which are not singular flow-fanning exponents of index $i$ are the same thing as flow-fanning exponents of index $i+1$ which are not of index $i$, for $i\geq 0$: $\Theta_{i+1}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)=\Theta_{i+1}(\Gamma)\setminus\Theta_{i}(\Gamma)$ for $i\geq 0$.
  6. (6)
    The set of exponents $\mathcal{E}_{E}(\mathcal{C})$ is the union of the flow-fanning exponents of all indices: ${\mathcal{E}}_{E}({\mathcal{C}})=\cup_{i\geq 0}\Theta_{i}(\Gamma)$.
  7. (7)
    Actually, one has the following union where $\sqcup$ indicates disjoint union:
    $\mathcal{E}_{E}(\mathcal{C})=\Theta_{0}^{\circ}(\Gamma)\bigsqcup\sqcup_{i\geq 0}(\Theta_{i+1}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)).$
  8. (8)
    Finally, whenever $\Gamma_{i}$ is transverse to a nonempty $E_{i}$ (in particular $P_{i}$ is a trace point of $E_{i}$ and $\Gamma_{i}$ is non-singular at $P_{i}$), then all flow-fanning exponents of index $i+1$ correspond to singular flow-fanning exponents of index $i$. That is, if $\tau$ is the first index such that $\Gamma_{\tau}$ is transverse to a nonempty $E_{\tau}$, then $\Theta_{i+1}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)=\emptyset$ for $i\geq\tau$.
Proof.
Property (1) holds because any $X\in{\mathfrak{X}}_{E,i}^{\circ}$ can be lifted by $\pi_{i+1}$ to a vector field in ${\mathfrak{X}}_{E,i+1}$ because $X$ is singular at $P_{i}$.
Let $i\in\mathbb{Z}_{\geq 1}\setminus\mathbf{Slid}_{\mathcal{C}}$ be a non-sliding divisor. By Lemma 5.2, $P_{i}$ is not a trace point of $E_{i}$, so that $E_{i}$ has two irreducible components at $P_{i}$. By definition, any $X\in{\mathfrak{X}}_{E,i}$ preserves both components, and must be singular, which gives Property (2).
To prove (3), let $i\in\mathbf{Slid}_{\mathcal{C}}$ and $k\in\Theta_{i}(\Gamma)$. By definition, $k=\theta_{X}(\Gamma)$ for some $X\in{\mathfrak{X}}_{E,i}$. The flow of $X$ induces a family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ whose sliding divisor $i^{\prime}$ is $i^{\prime}=i$ when $X\not\in{\mathfrak{X}}_{E,i}^{\circ}$, and $i^{\prime}>i$ otherwise. In any case, $k=\hat{\theta}(i^{\prime})\geq\hat{\theta}(i)$ by Proposition 5.1, which is the first assertion. We also know that $\Theta_{i}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)\subset\{\hat{\theta}(i)\}$ and $\Theta_{i}^{\circ}(\Gamma)\subset\mathbb{Z}_{>\hat{\theta}(i)}$. As $\hat{\theta}(i)$ is equal to $\theta_{X}(\Gamma)$ for any $X\in{\mathfrak{X}}_{E,i}\setminus{\mathfrak{X}}_{E,i}^{\circ}$ it belongs to $\Theta_{i}(\Gamma)$, and we have the second assertion.
For (4), take $k\in\Theta_{j}(\Gamma)$ with $i<j$ so that $k=\theta_{X}(\Gamma)$ for some $X\in{\mathfrak{X}}_{E,j}$. The family induced by the flow of $X$ on $\Gamma$ has a sliding divisor $j^{\prime}>i$, which gives $k=\hat{\theta}(j^{\prime})>\hat{\theta}(i)$, and we are done.
Consider (5). The inclusion $\supset$ is clear since $\Theta_{i}^{\circ}(\Gamma)\subset\Theta_{i}(\Gamma)$. Let us prove the direct inclusion $\subset$. Take $k\in\Theta_{i+1}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)$. It suffices to show $k\not\in\Theta_{i}(\Gamma)$. If it were not so, then $k\in\Theta_{i}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)$, so that $i\in\mathbf{Slid}_{\mathcal{C}}$ and $k=\hat{\theta}(i)$ by (2) and (3). Since $\Theta_{i+1}(\Gamma)\subset\mathbb{Z}_{>\hat{\theta}(i)}$ by (4), we obtain a contradiction.
The inclusion $\supset$ in (6) is obvious. For the direct inclusion, take $k\in{\mathcal{E}}_{E}({\mathcal{C}})$. By Proposition 5.1 it is of the form $k=\hat{\theta}(i)$, and thus it belongs to $\Theta_{i}(\Gamma)$.
Property (7) is the longest to prove. First, we claim that if $k\in\Theta_{i}(\Gamma)\cap\Theta_{j}(\Gamma)$ with $0\leq i<j$ then $k\in\Theta_{i}^{\circ}(\Gamma)\cap\Theta_{i+1}(\Gamma)$. By (1), It suffices to show that $k\in\Theta_{i}^{\circ}(\Gamma)$. This holds for $i=0$ by definition, so we assume $i\geq 1$. Suppose, aiming at contradiction, that $k\not\in\Theta_{i}^{\circ}(\Gamma)$. Now, by (2), we get $i\in\mathbf{Slid}_{\mathcal{C}}$; by (3), $k\in\Theta_{i}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)=\{\hat{\theta}(i)\}$; and, finally, by (4), $k\in\Theta_{j}(\Gamma)\subset\mathbb{Z}_{>\hat{\theta}(i)}$, which is the desired contradiction. Thus, $k\in\Theta_{i}^{\circ}(\Gamma)$.
Now if $k\in\Theta_{i}(\Gamma)\cap\Theta_{j}(\Gamma)$ for some $0\leq i<j$, then $k\in\Theta_{l}^{\circ}(\Gamma)$ for any $i\leq l<j$ by the above claim and induction on $l$. We can now prove (7). The equality follows from (6). It remains to show that the union is disjoint. Given $k\in\Theta_{j}(\Gamma)\setminus\Theta_{j-1}^{\circ}(\Gamma)$, if it belonged to $\Theta_{i}(\Gamma)\setminus\Theta_{i-1}^{\circ}(\Gamma)$ for $i<j$ or to $\Theta_{0}^{\circ}(\Gamma)$, then $k\in\Theta_{j-1}^{\circ}(\Gamma)$, which is impossible. Hence, the union is disjoint.
Finally, Property (8) holds because the sets $\Theta_{i+1}(\Gamma)$ and $\Theta_{i}^{\circ}(\Gamma)$ are both equal to $\mathbb{Z}_{\geq\hat{\theta}(i+1)}$. ∎
Definition 5.7.
Let $\Gamma$ be a germ of irreducible curve and $E$ a germ of normal crossings divisor. We define
$\Lambda_{\Gamma,\circ}^{E}=\{\nu_{\Gamma}(\omega)\ |\ w\in\hat{\Omega}_{E}(\mathbb{C}^{2},0)\ \mathrm{and}\ \omega(0)=0\}.$
Proposition 5.3.
Consider the equisingularity class ${\mathcal{C}}$ of a suitable pair $(\Gamma,E)$. Then
$\begin{split}&\Theta_{i}^{\circ}(\Gamma)=-\nu_{P_{i}}(\Gamma_{i})+\sum_{l=1}^{i}\nu_{P_{l}}(\Gamma_{l})+\Lambda_{\Gamma_{i},\circ}^{E_{i}},\\ &\Theta_{i+1}(\Gamma)=\sum_{l=1}^{i}\nu_{P_{l}}(\Gamma_{l})+\Lambda_{\Gamma_{i+1}}^{E_{i+1}}\end{split}$
for any $i\geq 0$.
Proof.
Denote $\mu_{j}=\nu_{P_{j}}(\Gamma_{j})$ and fix $i\geq 0$. We have $\Lambda_{\Gamma_{i},\circ}^{E_{i}}=\mu_{i}+\Theta_{0}^{\circ}(\Gamma_{i})$ [12, Theorem 3.21]. This gives
(32) $\Theta_{i}^{\circ}(\Gamma)=\sum_{l=1}^{i}\mu_{l}+\Theta_{0}^{\circ}(\Gamma_{i})=-\mu_{i}+\sum_{l=1}^{i}\mu_{l}+\Lambda_{\Gamma_{i},\circ}^{E_{i}}$
by Remark 5.4. Notice that $\Lambda_{\Gamma_{i+1}}^{E_{i+1}}\setminus\Lambda_{\Gamma_{i+1},\circ}^{E_{i+1}}=\emptyset$ if $i+1\not\in\mathbf{Slid}_{\mathcal{C}}$ and $\Lambda_{\Gamma_{i+1}}^{E_{i+1}}\setminus\Lambda_{\Gamma_{i+1},\circ}^{E_{i+1}}=\{\mu_{i}\}$ otherwise. As a consequence, if $i+1\not\in\mathbf{Slid}_{\mathcal{C}}$, we have
$\Theta_{i+1}(\Gamma)=\Theta_{i+1}^{\circ}(\Gamma)=\sum_{l=1}^{i}\mu_{l}+\Lambda_{\Gamma_{i+1},\circ}^{E_{i+1}}=\sum_{l=1}^{i}\mu_{l}+\Lambda_{\Gamma_{i+1}}^{E_{i+1}}$
by (2) in Proposition 5.2 and equation (32). On the other hand, if $i+1\in\mathbf{Slid}_{\mathcal{C}}$, we have
$\Theta_{i+1}(\Gamma)=\Theta_{i+1}^{\circ}(\Gamma)\cup\{\hat{\theta}(i+1)\}=\sum_{l=1}^{i}\mu_{l}+(\Lambda_{\Gamma_{i+1},\circ}^{E_{i+1}}\cup\{\mu_{i}\})=\sum_{l=1}^{i}\mu_{l}+\Lambda_{\Gamma_{i+1}}^{E_{i+1}}$
by (3) in Proposition 5.2, equation (32) and Definition 5.4. ∎
Proposition 5.4.
Consider the equisingularity class ${\mathcal{C}}$ of a suitable pair $(\Gamma,E)$. We have
(33) ${\mathcal{E}}_{E}({\mathcal{C}})\setminus\Theta_{0}^{\circ}(\Gamma)=\bigsqcup_{0\leq i<\tau}(\Theta_{i+1}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma))=\bigsqcup_{0\leq i<\tau}(\Theta_{i+1}(\Gamma)\setminus\Theta_{i}(\Gamma))$
where $\tau$ is the first index such that $\Gamma_{\tau}$ is non-singular and transverse to a non-empty $E_{\tau}$. Moreover,
$\sharp({\mathcal{E}}_{E}({\mathcal{C}})\setminus\Theta_{0}^{\circ}(\Gamma))=\sum_{0\leq i<\tau}\sharp([\nu_{P_{i}}(\Gamma_{i})+\Lambda_{\Gamma_{i+1}}^{E_{i+1}}]\setminus\Lambda_{\Gamma_{i},\circ}^{E_{i}}).$
Proof.
Denote $\mu_{j}=\nu_{P_{j}}(\Gamma_{j})$. Equation (33) is a consequence of (7), (8) and (5) in Proposition 5.2. We know that
$\Theta_{i+1}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma)=(\sum_{l=1}^{i}\mu_{l}+\Lambda_{\Gamma_{i+1}}^{E_{i+1}})\setminus(-\mu_{i}+\sum_{l=1}^{i}\mu_{l}+\Lambda_{\Gamma_{i},\circ}^{E_{i}})$
by Proposition 5.3, so that
$\sharp(\Theta_{i+1}(\Gamma)\setminus\Theta_{i}^{\circ}(\Gamma))=\sharp([\mu_{i}+\Lambda_{\Gamma_{i+1}}^{E_{i+1}}]\setminus\Lambda_{\Gamma_{i},\circ}^{E_{i}})$
for any $i\geq 0$, which completes the proof. ∎
6. Contacts of Kähler differentials and blow-ups
We have finally arrived at the endpoint of our work. Here, we provide an application of our methods. By studying the behavior of contacts of Kähler differentials under blow-up, using the results about flows of vector fields, we can give an alternative proof of Genzmer’s formula [13].
As above, $\Gamma$ is a germ of irreducible curve and $E$ a germ of normal crossings divisor at $(0,0)$. We keep the notation of Definition 5.1.
Definition 6.1.
Let ${\mathcal{C}}$ be the equisingularity class of $(\Gamma,E)$, we denote
$\Lambda_{\mathcal{C},\circ}^{E}=\Lambda_{\Gamma^{\prime},\circ}^{E}\quad\mathrm{and}\quad\Lambda_{{\mathcal{C}}_{1}}^{{E}_{1}}=\Lambda_{{\Gamma}_{1}^{\prime}}^{{E}_{1}}$
for any generic $\Gamma^{\prime}\in\mathcal{C}$.
6.1. Main results
Proposition 5.4 motivates us to compare $\Lambda_{\mathcal{C},\circ}^{E}$ with $\Lambda_{{\mathcal{C}}_{1}}^{{E_{1}}}$. We first state the key results, and delay their proofs to the forthcoming subsections.
Definition 6.2.
Given $n\in\mathbb{Z}_{\geq 2}$, we define
$\sigma(n)=\left\{\begin{array}[]{l}\frac{(n-2)(n-4)}{4}\;\mathrm{if}\;n\equiv 0\mod(2)\\[5.0pt] \frac{(n-3)^{2}}{4}\;\mathrm{if}\;n\equiv 1\mod(2).\end{array}\right.$
The next result is the induction step required to prove Genzmer’s formula. Recall that $E=\{x^{\delta_{x}}y^{\delta_{y}}=0\}$ with $\delta_{x},\delta_{y}\in\left\{0,1\right\}$.
Proposition 6.1.
The inclusion $\Lambda_{\mathcal{C},\circ}^{E}\subset\nu(\mathcal{C})+\Lambda_{{\mathcal{C}}_{1}}^{{E_{1}}}$ holds, and we can count the difference:
(34) $\sharp([\nu(\mathcal{C})+\Lambda_{{\mathcal{C}}_{1}}^{{E_{1}}}]\setminus\Lambda_{\mathcal{C},\circ}^{E})=\sigma(\nu(\mathcal{C})+\delta_{x}+\delta_{y}).$
Proposition 6.1 is a consequence of Proposition 6.2 below. We introduce now the concept of “parallel families” of $1$-forms for $\mathcal{C}$ and $\mathcal{C}_{1}$, as the main tool in the comparison of $\Lambda_{\mathcal{C},\circ}^{E}$ and $\Lambda_{{\mathcal{C}}_{1}}^{{E_{1}}}$.
Definition 6.3.
We say that an $E$-good (resp. $E$-initial, $E$-terminal, $E$-nice basis) family ${\mathcal{G}}_{l}^{s}=(\Omega_{1}^{s},\ldots,\Omega_{r}^{s})$ of level $l$ of $\mathcal{C}$ and an ${E}_{1}$-good (resp. $E_{1}$-initial, $E_{1}$-terminal, $E_{1}$-nice basis) family $\tilde{\mathcal{G}}_{l}^{s}=(\tilde{\Omega}_{1}^{s},\ldots,\tilde{\Omega}_{r}^{s})$ of the same level $l$ of $\mathcal{C}_{1}$ are parallel if the following properties hold:
  • The leading variables of $\Omega_{j}^{s}$ and $\tilde{\Omega}_{j}^{s}$ coincide for any $1\leq j\leq r$;
  • $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j}^{s})=\nu_{{\mathcal{C}}}({\Omega}_{j}^{s})-n\lfloor\frac{j-1+\delta_{x}+\delta_{y}}{2}\rfloor$ for any $1\leq j\leq r$.
Proposition 6.2.
There exist an $E$-nice basis $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ for ${\mathcal{C}}$ and an ${E}_{1}$-nice basis $\mathbf{\tilde{\Omega}}=(\tilde{\Omega}_{1},\ldots,\tilde{\Omega}_{n})$ for ${\mathcal{C}}_{1}$ that are parallel. Moreover,
(35) $\nu(\Omega_{j})=\left\lfloor\frac{j-1+\delta_{x}+\delta_{y}}{2}\right\rfloor$
holds for any $1\leq j\leq n$.
Corollary 6.1.
Let $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ be an $E$-nice basis. Define
$\hat{\Omega}_{j}=(\pi_{1}^{*}\Omega_{j})/x_{1}^{\nu(\Omega_{j})}$
for $1\leq j\leq n$. Then $(\hat{\Omega}_{1},\ldots,\hat{\Omega}_{n})$ is an $E_{1}$-Apery basis of ${\mathcal{C}}_{1}$. In other words, we have
$\hat{\Omega}_{E_{1}}(\mathcal{X}_{1},P_{1})=\mathbb{C}[[x_{1}]]\hat{\Omega}_{1,\Gamma^{\prime}}+\ldots+\mathbb{C}[[x_{1}]]\hat{\Omega}_{n,\Gamma^{\prime}}+{\mathcal{I}}_{E_{1},\Gamma^{\prime}}$
for any generic $\Gamma^{\prime}\in{\mathcal{C}}_{1}$.
Proof.
By construction $\hat{\Omega}_{j}$ is a formal $1$-form defined at ${P}_{1}$ and preserving ${E}_{1}$ for any $1\leq j\leq n$. The equality $\nu_{{\mathcal{C}}_{1}}(\hat{\Omega}_{j})=\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j})$ for any $1\leq j\leq n$ holds because $\mathbf{\Omega}$ and $\mathbf{\tilde{\Omega}}$ are parallel by Proposition 6.2. As a consequence, $(\hat{\Omega}_{1},\ldots,\hat{\Omega}_{n})$ is an Apery basis of ${\mathcal{C}}_{1}$. ∎
6.2. Proof of Proposition 6.2
Before proceeding to the core of the proof, we consider the different cases arising in it in order to construct parallel families of the starting level $l_{0}$ for $\Gamma$ and $\Gamma_{1}$ in each case. Once these are constructed, we can give a common argument. Assume that $(\Gamma,E)$ is a suitable pair in coordinates $(x,y)$. We consider coordinates $(x_{1},y_{1})$ to study the pair $(\Gamma_{1},E_{1})$ even if in such coordinates it could be unsuitable.
6.2.1. Case $E=\emptyset$ or $E=\{x=0\}$
In this case, $l_{0}=0$ for $\mathcal{C}$ and ${E}_{1}=\{x_{1}=0\}$. A terminal family ${\mathcal{T}}_{0}$ of level 00 of $\mathcal{C}$ is $(\Omega_{1},\Omega_{2})=(dx,x^{\delta_{x}}df_{1})$ whereas a terminal family $\tilde{\mathcal{T}}_{0}$ of level 00 of ${\mathcal{C}}_{1}$ is $(\tilde{\Omega}_{1},\tilde{\Omega}_{2})=(dx_{1},x_{1}df_{1,1})$ (see subsection 5.1.1). Of course, $dx,dx_{1}\in\Delta_{n,<}$ and $df_{1},x_{1}d{f}_{1,1}\in\Delta_{\beta_{1},<}^{0}$, and also
$\nu_{\mathcal{C}}(dx)=\nu_{{\mathcal{C}}_{1}}(dx_{1})=n\quad\mathrm{and}\quad\nu_{\mathcal{C}}(x^{\delta_{x}}df_{1})-\nu_{{\mathcal{C}}_{1}}(x_{1}d{f}_{1,1})=n\delta_{x}.$
Thus, ${\mathcal{T}}_{0}$ and $\tilde{\mathcal{T}}_{0}$ are parallel and equation (35) holds for $j\in\{1,2\}$.
6.2.2. Case $\delta_{y}=1$ and $\beta_{0}\in(n)\setminus\{n\}$
In this case, $l_{0}=0$ also, and ${E}_{1}=\{x_{1}y_{1}=0\}$. As explained in subsubsection 4.2.2, ${\mathcal{T}}_{0}=(\Omega_{1},\Omega_{2})$ where
$\Omega_{1}=x^{\delta_{x}}dy\in\Delta_{\beta_{0},<}^{0},\ \Omega_{2}\in\Delta_{\beta_{1},<}^{0},\ \nu_{\mathcal{C}}(\Omega_{1})=n\delta_{x}+\beta_{0}\ \mathrm{and}\ \nu_{\mathcal{C}}(\Omega_{2})=n+\beta_{1}.$
Analogously, $\tilde{\mathcal{T}}_{0}=(\tilde{\Omega}_{1},\tilde{\Omega}_{2})$ where $\tilde{\Omega}_{1}=x_{1}dy_{1}\in\Delta_{\beta_{0},<}^{0}$, $\tilde{\Omega}_{2}\in\Delta_{\beta_{1},<}^{0}$, $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{1})=\beta_{0}$ and $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2})=\beta_{1}$. Both families $\mathcal{T}_{0}$ and $\tilde{\mathcal{T}}_{0}$ are terminal and indeed parallel, and equation (35) holds for any $j\in\{1,2\}$.
6.2.3. $\delta_{y}=1$ and $\beta_{0}=n$
Again, $l_{0}=0$ for $\mathcal{C}$, $E=\{xy=0\}$ and ${E}_{1}=\{x_{1}=0\}$. As described in subsubsection 4.2.2, ${\mathcal{T}}_{0}=(\Omega_{1},\Omega_{2})$ where $\Omega_{1}=xdy\in\Delta_{n,<}^{0}$, $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$, $\nu_{\mathcal{C}}(\Omega_{1})=2n$ and $\nu_{\mathcal{C}}(\Omega_{2})=n+\beta_{1}$. We also have $\tilde{\mathcal{T}}_{0}=(\tilde{\Omega}_{1},\tilde{\Omega}_{2})$ where $\tilde{\Omega}_{1}=dx_{1}\in\Delta_{n,<}$, $\tilde{\Omega}_{2}=x_{1}d{f}_{1,1}\in\Delta_{\beta_{1},<}^{0}$, $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{1})=n$ and $\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2})=\beta_{1}$. Once more, the families $\mathcal{T}_{0}$ and $\tilde{\mathcal{T}}_{0}$ are both terminal, and parallel. Equation (35) for $j\in\{1,2\}$ holds too.
6.2.4. $\delta_{y}=1$ and $\beta_{0}\not\in(n)$
In this case $l_{0}=1$ for $(\Gamma,E)$. We have $E=\{x^{\delta_{x}}y=0\}$ and ${E}_{1}=\{x_{1}y_{1}=0\}$. Consider the $E$-good family ${\mathcal{G}}_{1}^{1}$ for ${\mathcal{C}}$ and the ${E}_{1}$-good family $\tilde{\mathcal{G}}_{1}^{1}$ for ${\mathcal{C}}_{1}$ as described in subsubsection 4.2.3. We have
$\Omega_{2j+1}^{1}=x^{\delta_{x}}y^{j}dy\in\Delta_{\beta_{1},<},\quad\tilde{\Omega}_{2j+1}^{1}=x_{1}y_{1}^{j}dy_{1}\in\Delta_{\beta_{1},<}$
and
$\nu_{\mathcal{C}}(\Omega_{2j+1}^{1})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2j+1}^{1})=n\nu(\Omega_{2j+1}^{1})=n\left\lfloor\frac{(2j+1)-1+\delta_{x}+\delta_{y}}{2}\right\rfloor$
if $1\leq 2j+1\leq 2{\nu}_{1}$. We also get $\Omega_{2j+2}^{1},\tilde{\Omega}_{2j+2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{1}),<}^{0}$ and
$\nu_{\mathcal{C}}(\Omega_{2j+2}^{1})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2j+2}^{1})=(n+j\beta_{1}+{\mathfrak{n}}(\beta_{1}))-(n+j(\beta_{1}-n)+({\mathfrak{n}}(\beta_{1})-n)).$
As a consequence
$\nu_{\mathcal{C}}(\Omega_{2j+2}^{1})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2j+2}^{1})=n\nu(\Omega_{2j+2}^{1})=n\left\lfloor\frac{(2j+2)-1+\delta_{x}+\delta_{y}}{2}\right\rfloor$
if $1\leq 2j+2\leq 2{\nu}_{1}$. Thus ${\mathcal{G}}_{1}^{1}$ and $\tilde{\mathcal{G}}_{1}^{1}$ are parallel families (although not necessarily terminal). The members of the family $(\Omega_{j}^{1},\ldots,\Omega_{2{\nu}_{1}}^{1})$ satisfy equation (35).
6.2.5. End of the proof of Proposition 6.2
We apply the procedure of Section $1$ to build terminal families $\mathcal{T}_{l}$ and $\tilde{\mathcal{T}}_{l}$ for ${\mathcal{C}}$ and ${\mathcal{C}}_{1}$ respectively for $0\leq l\leq g$. We are going to show that $\mathcal{T}_{l}$ and $\tilde{\mathcal{T}}_{l}$ are parallel for any $0\leq l\leq g$ by induction on $l$.
Case $l_{0}=0$. As seen above, $\mathcal{T}_{0}$ and $\tilde{\mathcal{T}}_{0}$ are parallel. Assume that there exist parallel families ${\mathcal{T}}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ and $\tilde{\mathcal{T}}_{l}=(\tilde{\Omega}_{1},\ldots,\tilde{\Omega}_{2{\nu}_{l}})$ of level $l<g$ such that ${\mathcal{T}}_{l}$ satisfies equation (35). Define the families ${\mathcal{G}}_{l+1}^{0}=(\Omega_{1}^{0},\ldots\Omega_{2{\nu}_{l+1}}^{0})$ and $\tilde{\mathcal{G}}_{l+1}^{0}=(\tilde{\Omega}_{1}^{0},\ldots\tilde{\Omega}_{2{\nu}_{l+1}}^{0})$ setting $\Omega_{2a\nu_{l}+j}^{0}=f_{l+1}^{a}\Omega_{j}$ and $\tilde{\Omega}_{2a\nu_{l}+j}^{0}={f}_{1,l+1}^{a}\tilde{\Omega}_{j}$. This gives
(36) $\nu(\Omega_{2a\nu_{l}+j}^{0})=a\nu_{l}+\left\lfloor\frac{j-1+\delta_{x}+\delta_{y}}{2}\right\rfloor=\left\lfloor\frac{(2a\nu_{l}+j)-1+\delta_{x}+\delta_{y}}{2}\right\rfloor$
for $1\leq 2a\nu_{l}+j\leq 2{\nu}_{l+1}$. Those two families are respectively $E$-initial and ${E}_{1}$-initial for ${\mathcal{C}}$ and ${\mathcal{C}}_{1}$. Since ${\mathcal{T}}_{l}$ and $\tilde{\mathcal{T}}_{l}$ are parallel, and $\beta_{l+1}$ is the leading variable of $f_{l+1}^{a}\Omega_{j}$, and ${f}_{1,l+1}^{a}\tilde{\Omega}_{j}$ for $1\leq j\leq\tilde{\nu}_{l}$ and $a>0$, it follows that the leading variables of $\Omega_{j}^{0}$ and $\tilde{\Omega}_{j}^{0}$ coincide for any $1\leq j\leq\tilde{\nu}_{l+1}$. Moreover,
$\nu_{{\mathcal{C}}}({\Omega}_{2a\nu_{l}+j}^{0})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2a\nu_{l}+j}^{0})=a(\overline{\beta}_{l+1}-(\overline{\beta}_{l+1}-\nu_{l}n))+n\left\lfloor\frac{j-1+\delta_{x}+\delta_{y}}{2}\right\rfloor$
and thus
$\nu_{{\mathcal{C}}}({\Omega}_{2a\nu_{l}+j}^{0})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2a\nu_{l}+j}^{0})=n\left\lfloor\frac{2a\nu_{l}+j-1+\delta_{x}+\delta_{y}}{2}\right\rfloor.$
As a consequence, ${\mathcal{G}}_{l+1}^{0}$ and $\tilde{\mathcal{G}}_{l+1}^{0}$ are parallel.
Take $1\leq k\leq 2{\nu}_{l+1}$, and assume $\Omega_{k}^{1}\neq\Omega_{k}^{0}$. Then there exists $1\leq j<k$ such that $[\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}$. Since
$[\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}=[\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{k}^{0})]_{n},$
we deduce that $\tilde{\Omega}_{k}^{1}\neq\tilde{\Omega}_{k}^{0}$. The leading variable of $\Omega_{k}^{0}$ is $\beta=\beta_{l+1}$ and, as ${\mathcal{G}}_{l+1}^{0}$ and $\tilde{\mathcal{G}}_{l+1}^{0}$ are parallel, we conclude that the leading variables of $\tilde{\Omega}_{k}^{1}$ and $\Omega_{k}^{1}$ are equal to ${\mathfrak{n}}(\beta)$ and also that
$\nu_{{\mathcal{C}}}({\Omega}_{k}^{1})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{k}^{1})=(\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})+m)-(\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{k}^{0})+m)=\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{k}^{0}),$
where $m={\mathfrak{n}}(\beta)-\beta$. At the same time,
$\Omega_{k}^{1}=C_{\beta^{\prime},j}\Omega_{k}^{0}-C_{\beta,k}x^{d}\Omega_{j}^{0}\quad\mathrm{and}\quad\tilde{\Omega}_{k}^{1}=\tilde{C}_{\beta^{\prime},j}\tilde{\Omega}_{k}^{0}-\tilde{C}_{\beta,k}x^{\tilde{d}}\tilde{\Omega}_{j}^{0},$
(see equation (17)), where $d=\tilde{d}+(\nu(\Omega_{k}^{0})-\nu(\Omega_{j}^{0}))$ by the parallelism of ${\mathcal{G}}_{l+1}^{0}$ and $\tilde{\mathcal{G}}_{l+1}^{0}$, and by equation (36). Notice also that $\tilde{d}\geq 0$ by construction. We deduce that $\nu(\Omega_{k}^{0})\leq\nu(x^{d}{\Omega}_{j}^{0})$. Since $\Omega_{j}^{0}\in M_{j}^{E}$ and $\Omega_{k}^{0}\in\hat{M}_{k}^{E}$, we get $\nu(\Omega_{k}^{1})=\nu(\Omega_{k}^{0})$. Thus equation (35) for $(\Omega_{1}^{1},\ldots,\Omega_{2{\nu}_{l+1}}^{1})$ is a consequence of equation (36). In the same way, $\tilde{\Omega}_{k}^{1}\neq\tilde{\Omega}_{k}^{0}$ implies $\Omega_{k}^{1}\neq\Omega_{k}^{0}$.
By repeating this argument finitely many times, we obtain that ${\mathcal{T}}_{l+1}$ and $\tilde{\mathcal{T}}_{l+1}$ are parallel and ${\mathcal{T}}_{l+1}$ satisfies equation (35). Thus, by induction on $l$, we obtain parallel terminal families ${\mathcal{T}}_{g}$ and $\tilde{\mathcal{T}}_{g}$ and hence parallel nice bases ${\bf\Omega}$ and $\tilde{\bf{\Omega}}$ such that ${\bf\Omega}$ satisfies equation (35).
Case $l_{0}=1$. In this remaining case, when $\delta_{y}=1$ and $\beta_{0}\not\in(n)$, the families ${\mathcal{G}}_{1}^{1}$ and $\tilde{\mathcal{G}}_{1}^{1}$ in subsection 6.2.4 are parallel and ${\mathcal{G}}_{1}^{1}$ satisfies equation (35). Following the same argument as above, we obtain parallel terminal families of level $l_{0}=1$, ${\mathcal{T}}_{1}$ and $\tilde{\mathcal{T}}_{1}$ with equation (35) holding for ${\mathcal{T}}_{1}$. The rest of the proof is in the same as in the previous case.
6.3. Proof of Proposition 6.1
We can assume that $(\Gamma,E)$ is a suitable pair by Remark 5.1. We have $\Lambda_{\mathcal{C},\circ}^{E}\subset\Lambda_{{\mathcal{C}}_{1}}^{{E}_{1}}+\nu(\mathcal{C})$ because $\Theta_{0}^{0}(\Gamma)\subset\Theta_{1}(\Gamma)$ by Proposition 5.2 and Proposition 5.3.
We know that there exist parallel nice bases ${\bf\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ and ${\bf\tilde{\Omega}}=(\tilde{\Omega}_{1},\ldots,\tilde{\Omega}_{n})$ by Proposition 6.2. Again, we consider the different cases for $E$.
Case $E=\emptyset$. An $E$-nice basis for singular $1$-forms for ${\mathcal{C}}$ is clearly obtained by replacing $(\Omega_{1},\Omega_{2})$ with $(xdx,xdf_{1})$ in $\mathbf{\Omega}$. We have
$\nu_{{\mathcal{C}}}(xdx)-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{1})=n\quad\mathrm{and}\quad\nu_{{\mathcal{C}}}(xdf_{1})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{2})=n.$
Since $\nu_{{\mathcal{C}}}({\Omega}_{j})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j})=n\lfloor\frac{j-1}{2}\rfloor$ for $j\geq 3$, it follows that
$\sharp([\Lambda_{{\mathcal{C}}_{1}}^{{E}_{1}}+n]\setminus\Lambda_{\mathcal{C},\circ}^{E})=\underbrace{0+0+0+0+1+1+2+2+\ldots}_{n\ \mathrm{terms}}=\sigma(n).$
Case $E\equiv\{x=0\}$ or $E\equiv\{y=0\}$. This means $\delta_{x}+\delta_{y}=1$. An $E$-nice basis for singular forms $1$-forms for ${\mathcal{C}}$ is obtained by simply replacing $\Omega_{1}$ with $x\Omega_{1}$ in $\mathbf{\Omega}$. Since $\nu_{{\mathcal{C}}}(x\Omega_{1})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{1})=n$ and $\nu_{{\mathcal{C}}}({\Omega}_{j})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j})=n\lfloor\frac{j}{2}\rfloor$ for $j\geq 2$, we deduce
$\sharp([\Lambda_{{\mathcal{C}}_{1}}^{{E}_{1}}+n]\setminus\Lambda_{\mathcal{C},\circ}^{E})=\underbrace{0+0+0+1+1+2+2+\ldots}_{n\ \mathrm{terms}}=\sigma(n+1).$
Case $E\equiv\{xy=0\}$. That is, $\delta_{x}+\delta_{{y}}=2$. Since all the $1$-forms that preserve $E$ are singular and $\nu_{{\mathcal{C}}}({\Omega}_{j})-\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j})=n\lfloor\frac{j+1}{2}\rfloor$ for $j\geq 1$, it follows that
$\sharp([\Lambda_{{\mathcal{C}}_{1}}^{{E}_{1}}+n]\setminus\Lambda_{\mathcal{C},\circ}^{E})=\underbrace{0+0+1+1+2+2+\ldots}_{n\ \mathrm{terms}}=\sigma(n+2).$
This completes the proof.
6.4. The dimension formula
The next formula is proved in [13]. It follows also from all our previous results.
Theorem 6.1.
Let $\Gamma$ be a germ of irreducible curve defined in a neighborhood of $0\in\mathbb{C}^{2}$. The dimension of the generic component of the analytic moduli of the equisingularity class of $\Gamma$ is equal to
$\sum_{0\leq i<\tau}\sigma(\nu_{P_{i}}(\Gamma_{i})+\delta_{i}),$
where $P_{0},\ldots,P_{\tau}$ are infinitely near points of $\Gamma$, $\tau$ is the first index such that $\Gamma_{\tau}$ is smooth, and $\delta_{i}$ is the number of irreducible components of $E_{i}$ passing through $P_{i}$.
Proof.
Notice that $\sigma(2)=\sigma(3)=0$, so that we do not need to assume that $\Gamma_{\tau}$ is transverse to $E_{\tau}$, just that it is non-singular. Up to replacing $\Gamma$ with another curve in its equisingularity class ${\mathcal{C}}$, we can assume that $\Gamma_{i}$ is generic in the equisingularity class of $(\Gamma_{i},E_{i})$ for any $0\leq i<\tau$ by Remark 5.8 . Now, Propositions 5.4 and 6.1 imply the result. ∎

### v3

@@
-Construction of a Basis of Kähler differentials for generic plane branches
+Explicit Computation of The Generic Component of the Analytic Moduli of a Plane Branch
 Abstract.
@@
 We compute a basis of the ${\mathbb{C}}[[x]]$-module of Kähler differentials for generic $\Gamma\in{\mathcal{C}}$, algorithmically, and study its behaviour under blow-up.
-As an application, we give an alternative proof for a formula of Genzmer, that provides the dimension of the generic component of the moduli of analytic classes in the equisingularity class of $\Gamma$.
-Key words and phrases:
-plane branch, approximate roots, moduli, Kähler differentials, flows of holomorphic vector fields
+As a first application, we give an algorithm providing the generic semimodule in an equisingularity class in terms of its multiplicity and its Puiseux characteristic exponents.
+As another application, we give an alternative proof for a formula of Genzmer, that provides the dimension of the moduli of analytic classes in the equisingularity class of $\Gamma$.
 1991 Mathematics Subject Classification:
-32S05, 32S65, 14H20, 32G13
+32G13, 32S05, 32S65, 14H20
 1. Introduction
@@
 There exists a Zariski open subset $U$ of ${\mathbb{C}}^{p}$ consisting of orbits of $G$ such that all orbits of $G$ in $U$ have the same dimension $d$ by a theorem of Rosenlicht [18].
-Indeed, $U/G$ can be identified with a Zariski open subset of an algebraic variety $W$ of dimension $p-d$ (cf.
+Indeed, $U/G$ can be identified with a Zariski open subset of an algebraic variety $W$ of dimension pdp-d (cf.
 [19, Chapter VI.1]) that represents the generic component of the moduli space $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$.
-The generic dimension $p-d$ of the moduli was studied in the case of one Puiseux characteristic exponent by Delorme [9] and Briançon, Granger and Maisonobe [4] who gave an algorithm for its calculation.
+The generic dimension pdp-d of the moduli was studied in the case of one Puiseux characteristic exponent by Delorme [9] and Briançon, Granger and Maisonobe [4] who gave an algorithm for its calculation.
 Genzmer provided a general formula for the generic dimension of the moduli space of any plane branch that depends on simple combinatorial data associated to the topological class of $\Gamma_{\circ}$ [13].
@@
 Consider, for simplicity, the subset $\mathcal{C}$ of the topological class $\tilde{\mathcal{C}}$ of $\Gamma_{\circ}$ of plane branches $\Gamma$ whose tangent cone is different from $x=0$.
-It is obvious that every analytic class of an element of $\tilde{\mathcal{C}}$ intersects $\mathcal{C}$ and hence $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}=\mathcal{C}/\sim_{\mathrm{ana}}$.
+Clearly every analytic class of an element of $\tilde{\mathcal{C}}$ intersects $\mathcal{C}$ and hence $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}=\mathcal{C}/\sim_{\mathrm{ana}}$.
 Given $\Gamma\in\mathcal{C}$, set ${\mathcal{I}}_{\Gamma}=\{\omega\in\hat{\Omega}(\mathbb{C}^{2},0):\Gamma(t)^{*}\omega\equiv 0\}$.
-We shall compute a basis of the ${\mathbb{C}}[[x]]$-module of Kähler differentials of $\Gamma$; more precisely, if $n$ is the multiplicity of $\Gamma_{\circ}$ at the origin, we shall find an Apery basis of $\Gamma$, i.e.
-germs of $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ such that
+We shall compute a “nice” basis of the ${\mathbb{C}}[[x]]$-module of Kähler differentials for generic $\Gamma\in\tilde{\mathcal{C}}$.
+Specifically:
+Theorem 1.1.
+There is a nice basis for $\mathcal{C}$}: If $n$ is the multiplicity of $\Gamma_{\circ}$ at the origin, there exists an open subset $U$ of $\tilde{\mathcal{C}}$ and germs of $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$, for $\Gamma\in U$, such that
   • $\hat{\Omega}(\mathbb{C}^{2},0)=\mathbb{C}[[x]]\Omega_{1,\Gamma}\oplus\ldots\oplus\mathbb{C}[[x]]\Omega_{n,\Gamma}\oplus{\mathcal{I}}_{\Gamma}$;
   • The values $\nu_{\Gamma}(\Omega_{1,\Gamma}),\ldots,\nu_{\Gamma}(\Omega_{n,\Gamma})$ define an Apery set of $\Lambda_{\Gamma}$ [2], i.e.
-the classes of $\nu_{\Gamma}(\Omega_{j,\Gamma})$ and $\nu_{\Gamma}(\Omega_{k,\Gamma})$ modulo $n$ are different for all $1\leq j<k\leq n$.
-To begin with, define $\overline{\Omega}_{2j+1}=y^{j}dx$ and $\overline{\Omega}_{2j+2}=y^{j}dy$ for any $j\in\mathbb{Z}_{\geq 0}$.
-We prove that can choose the elements of the basis $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ with sharp extra properties for any generic $\Gamma\in\mathcal{C}$:
-minimality:
-$\Omega_{j,\Gamma}\in\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$;
-monotonicity:
+the classes of $\nu_{\Gamma}(\Omega_{j,\Gamma})$ and $\nu_{\Gamma}(\Omega_{k,\Gamma})$ modulo $n$ are different for all $1\leq j<k\leq n$;
+  • $\Omega_{j,\Gamma}\in\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$, where $\overline{\Omega}_{2k+1}=y^{k}dx$ and $\overline{\Omega}_{2k+2}=y^{k}dy$ for any $k\in\mathbb{Z}_{\geq 0}$;
   • $\nu_{\Gamma}(\Omega_{j,\Gamma})<\nu_{\Gamma}(\Omega_{k,\Gamma})$ for all $1\leq j<k\leq n$.
-Let us comment these properties.
+Moreover, $\Omega_{j,\Gamma}$ depends polynomially on $\Gamma\in U$ and $\nu_{\Gamma}(\Omega_{1,\Gamma}),\ldots,\nu_{\Gamma}(\Omega_{n,\Gamma})$ do not depend on $\Gamma\in U$.
+The first two properties determine that $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ is an Apery basis [2] of $\Gamma$ for $\Gamma\in U$.
+The last two itens provide properties of minimality and monotonicity of the Apery basis respectively.
+Let us expand on this remark.
 By the properties of an Apery basis, it follows that $\oplus_{k=1}^{j}{\mathbb{C}}[[x]]\Omega_{k,\Gamma}$ is a free $\mathbb{C}[[x]]$-module of rank $j$.
@@
 Let us remark that the minimality property is related to “maximal contact” properties in the spirit of Hironaka [17] for the resolution of algebraic singularities and of Aroca, Hironaka and Vicente [3] for the complex analytic case.
-Maximal contact varieties are essential for the resolution of singularities in characteristic 00.
+Maximal contact varieties are essential for the resolution of singularities in characteristic 0.
 The $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ have maximal contact in the following sense: $\nu_{\Gamma}(\Omega_{j,\Gamma})$ is the maximum of the $\Gamma$-orders of the $1$-forms in $\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$ (Proposition 3.7).
@@
 its number of Puiseux characteristic exponents.
-We define the terminal family of $1$-forms of level 00 as the set consisting of the elementary $1$-forms $dx$ and $dy$.
+We define the terminal family of $1$-forms of level 0 as the set consisting of the elementary $1$-forms dxdx and $dy$.
 Given a terminal family of level $l$, the initial family of level $l+1$, with $0\leq l<g$, is obtained by combining that terminal family of level $l$ with data of topological flavor.
 Then we describe a straightforward algorithm that provides a terminal family of level $l+1$.
-At the end of level $g$ we obtain an Apery basis.
+At the end of level $g$ we obtain the Apery basis provided by Theorem 1.1.
 richness:
 The algorithm of construction of an Apery basis preserves a very rich structure.
-For instance, the $\Gamma$-values of all the $1$-forms involved in the algorithm can be calculated in terms of combinatorial data of the equisingularity class of $\Gamma_{\circ}$ (Propositions 3.3 and 3.4).
-In particular, we obtain a combinatorial algorithm that provides the generic semimodule $\Lambda_{\Gamma}$ for $\Gamma\in\mathcal{C}$.
+Indeed, we profit from our constructions to obtain that the $\Gamma$-values of all the $1$-forms involved in the construction of the Apery basis can be calculated in terms of $n,\beta_{1},\ldots,\beta_{g}$ (Propositions 3.3 and 3.4).
+Theorem 1.2.
+Algorithm 1 provides the generic semimodule $\Lambda_{\Gamma}$, for $\Gamma\in U$, in terms of $n$ and the Puiseux characteristic exponents $\beta_{1},\ldots,\beta_{g}$ of the equisingularity class $\mathcal{C}$.
 Being more precise, the initial family of level $l+1$, with $0\leq l<g$, is obtained by combining the terminal family of level $l$ with the approximate root $f_{l+1,\Gamma}$ (cf.
@@
 They provide sufficient control to calculate the generic semigroup $\Lambda_{\Gamma}$, for $\Gamma\in\mathcal{C}$, without calculating the $1$-forms in the Apery basis or any other $1$-forms along the process.
-The existence of an Apery basis with the minimality and monotonicity properties, for generic $\Gamma\in\mathcal{C}$, is proved in Theorem 3.1.
 Simplicity and richness are natural byproducts of our algorithm (Propositions 3.3 and 3.4).
-Moreover, the Apery bases given by Theorem 3.1 depend polynomially on $\Gamma$ for generic $\Gamma\in\mathcal{C}$.
+1.1. Nice bases and blow-up
 Our second aim is to relate the semimodule of $\Gamma$ with analogue concepts for the strict transform $\tilde{\Gamma}$ of $\Gamma$ by the blow-up of the origin.
@@
 Theorem 4.1 is the analogue of Theorem 1.1 for this setting.
+Note that Algorithm 1 (cf.
+Theorem 1.2) is generalized to this case.
 Minimality, monotonicity, simplicity and richness are all used to provide a simple comparison of the semimodules $\Lambda_{\Gamma}^{E}$ and $\Lambda_{\tilde{\Gamma}}^{\tilde{E}}$ (Proposition 6.1), where $\tilde{E}$ is the total transform of $E$.
-Let us provide an application of our construction.
-Consider the sequence $P_{0}=(0,0)\in\mathbb{C}^{2},P_{1},P_{2}\ldots$ of infinitely near points of a plane branch $\Gamma$ (Remark 5.3).
+As an application of our construction, consider the sequence $P_{0}=(0,0)\in\mathbb{C}^{2},P_{1},P_{2}\ldots$ of infinitely near points of a plane branch $\Gamma$ (Remark 5.3).
 Let $\pi_{j}$ be the blow-up of the point $P_{j-1}$.
@@
 Now, let us consider actions provided by the flow of a vector field on $\Gamma_{j}$.
-Actions of vector fields as a tool for the study of the analytic classification problem were introduced by the first author in [11].
+Actions of vector fields as a tool for the study of the analytic classification problem were introduced in [11].
 Consider a holomorphic vector field $X$ defined in a neighborhood of $P_{j}$ such that $E_{j}:=\tau_{j}^{-1}(0,0)$ is $X$-invariant but $\Gamma_{j}$ is not.
-We obtain a family of curves $\Gamma_{j}^{\epsilon}=\mathrm{exp}(\epsilon X)(\Gamma_{j})$ defined for $\epsilon\in\mathbb{C}$ close to 00.
+We obtain a family of curves $\Gamma_{j}^{\epsilon}=\mathrm{exp}(\epsilon X)(\Gamma_{j})$ defined for $\epsilon\in\mathbb{C}$ close to 0.
 By blow-down by $\tau_{j}$, we get a family ${(\Gamma^{\epsilon})}_{\epsilon\in(\mathbb{C},0)}$ in the equisingularity class of $\Gamma$.
@@
 Any $\beta$ obtained this way, by varying $X$, is called a flow-fanning exponents at index $j$, and its set is denoted by $\Theta_{j}(\Gamma)$ (Definition 5.6).
-The generic dimension of the moduli is equal to
-$\sum_{j\geq 1}\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma))$
-by Remark 5.11 and Proposition 5.4, where we assume that $\Gamma$ is generic in its equisingularity class $\mathcal{C}$ (otherwise we replace $\Gamma$ with a generic element).
-Such formula is obtained by introducing a correspondence between exponents of the Puiseux expansion of $\Gamma$ and divisors of the desingularization process (Proposition 5.1).
-We get a geometric interpretation of the generic dimension of the moduli in which $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ is the set of exponents of a Puiseux parametrization that can be changed by the action of a vector field defined in a neighborhood of $P_{j}$ but can not be changed by the action of a vector field defined in a neighborhood of any of the points $P_{0},P_{1},\ldots,P_{j-1}$.
+We profit from this construction to reinterpret the formula for the generic dimension of the moduli of $\Gamma$ by Genzmer.
+Theorem ([13]).
+Let $\Gamma$ be a germ of irreducible curve defined in a neighborhood of $0\in\mathbb{C}^{2}$.
+The dimension of the generic component of the analytic moduli of the equisingularity class of $\Gamma$ is equal to
+$\sum_{0\leq j<\tau}\sigma(\nu_{P_{j}}(\Gamma_{j})+\delta_{j}),$
+where $P_{0},\ldots,P_{\tau}$ are infinitely near points of $\Gamma$, $\tau$ is the first index such that $\Gamma_{\tau}$ is smooth, and $\delta_{i}$ is the number of irreducible components of $E_{i}$ passing through $P_{i}$ (cf.
+Definition 6.2).
+Theorem 1.3.
+Let $\Gamma$ be a germ of irreducible curve defined in a neighborhood of $0\in\mathbb{C}^{2}$.
+The dimension of the generic component of the analytic moduli of the equisingularity class of $\Gamma$ is equal to
+$\sum_{j\geq 1}\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)),$
+and indeed $\sharp(\Theta_{j+1}(\Gamma)\setminus\Theta_{j}(\Gamma))=\sigma(\nu_{P_{j}}(\Gamma_{j})+\delta_{j})$ for any $j\geq 0$, where we assume that $\Gamma$ is generic in its equisingularity class $\mathcal{C}$ (otherwise we replace $\Gamma$ with a generic element).
+We get a geometric interpretation of the generic dimension of the moduli space $\mathcal{C}/\sim_{\mathrm{ana}}$ in which $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ is the set of exponents of a Puiseux parametrization that can be changed by the action of a vector field defined in a neighborhood of $P_{j}$ but can not be changed by the action of a vector field defined in a neighborhood of any of the points $P_{0},P_{1},\ldots,P_{j-1}$.
+Thus, in particular we obtain a geometric interpretation of Genzmer’s formula.
+It is obtained by introducing a correspondence between exponents of the Puiseux expansion of $\Gamma$ and divisors of the desingularization process (Proposition 5.1).
 It is possible to compute $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ in terms of the semimodules $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 5.4) by applying the results on the correspondence between orders of Kähler differentials on $\Gamma$ and the tangency orders of vector fields with $\Gamma$ described in [12].
-For the case where $\Gamma$ is generic in its equisingularity class, our construction of the Apery basis and its properties allow to calculate in parallel $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 6.2), to obtain $\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma))$ for $j\in\mathbb{Z}_{\geq 1}$.
-This term is indeed the $j$th-term $\sigma(\nu_{P_{j-1}}(\Gamma_{j-1}\cup\tau_{j}^{-1}(0,0)))$ in Genzmer’ formula (Proposition 6.1) (cf.
+For the case where $\Gamma$ is generic in its equisingularity class, our construction of the Apery basis and its properties allow to calculate in parallel $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 6.2), to obtain $\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma))$ for $j\in\mathbb{Z}_{\geq 1}$ and then show that it is the $j$th-term $\sigma(\nu_{P_{j-1}}(\Gamma_{j-1}\cup\tau_{j}^{-1}(0,0)))$ in Genzmer’ formula (Proposition 6.1) (cf.
 Definitions 6.2 and 2.1).
-So, we obtain a geometric interpretation of Genzmer’s formula for the generic dimension of the moduli space $\mathcal{C}/\sim_{\mathrm{ana}}$ (Theorem 6.1).
 We introduce the setting of the paper in section 2.
 The construction of the Apery basis of a generic plane branch in an equisingularity class is carried out in section 3.
-This is generalized for the case of Kähler differentials preserving a normal crossings divisor $E$ in section 4.
+This is generalized for the case of Kähler differentials preserving a normal crossings divisor $E$ in section 4, where we also provide Algorithm 1 that gives the generic semimodule in an equisingularity class.
 Section 5 is devoted to introducing the correspondence between exponents of the Puiseux expansion and divisors of the desingularization process of a plane branch $\Gamma$ and to use such a setting to introduce the sets of flow-fanning exponents and describe their properties.
@@
 The calculation of the $\mathbb{C}[[x]]$-module of Kähler differentials on $\Gamma_{\circ}$ of $1$-forms preserving $E$ is simple if $\Gamma_{\circ}$ is smooth and transverse to $x=0$.
-Indeed, it is generated by $dx$, $dy$ or $xdy$ depending on whether $\{y=0\}\not\subset E$, $E=\{y=0\}$ or $E=\{xy=0\}$ respectively.
+Indeed, it is generated by $dx$, dydy or xdyxdy depending on whether $\{y=0\}\not\subset E$, $E=\{y=0\}$ or $E=\{xy=0\}$ respectively.
 Thus, we will assume from now on $n>1$.
@@
 We shall only use the fact that if $\ell\in\Lambda_{\Gamma}^{E}$, then $n+\ell\in\Lambda_{\Gamma}^{E}$.
-If we write $[\ell]_{m}$ for the class modulo $m$ of $\ell$, then: if $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ is a family of $1$-forms such that $\nu_{\Gamma}(\Omega_{i})$ is minimum in $[\nu_{\Gamma}(\Omega_{i})]_{n}\cap\Lambda_{\Gamma}^{E}$ for $i=1,\ldots,n$ and $[\nu_{\Gamma}(\Omega_{i})]_{n}\neq[\nu_{\Gamma}(\Omega_{j})]_{n}$ for $i\neq j$, then certainly $\mathbf{\Omega}$ gives, by push-back to $\Gamma$, a basis of Kähler $1$-forms over $\Gamma$ and preserving $E$, i.e.
+If we write $\ell]_{m}$ for the class modulo $m$ of $\ell$, then: if $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{n})$ is a family of $1$-forms such that $\nu_{\Gamma}(\Omega_{i})$ is minimum in $\nu_{\Gamma}(\Omega_{i})]_{n}\cap\Lambda_{\Gamma}^{E}$ for $i=1,\ldots,n$ and $\nu_{\Gamma}(\Omega_{i})]_{n}\neq[\nu_{\Gamma}(\Omega_{j})]_{n}$ for $i\neq j$, then certainly $\mathbf{\Omega}$ gives, by push-back to $\Gamma$, a basis of Kähler $1$-forms over $\Gamma$ and preserving $E$, i.e.
 $\Gamma^{*}\hat{\Omega}_{E}(\mathbb{C}^{2},0)=\Gamma^{*}({\mathbb{C}}[[x]]\Omega_{1}+\ldots+{\mathbb{C}}[[x]]\Omega_{n}).$
@@
 2.4. Approximate roots of $(\Gamma,E)$
-In order to provide bases of Kähler differentials for generic curves in $\mathcal{C}$, we shall need to use both the elementary $1$-forms $dx$ and $dy$ and the approximate roots of $\Gamma$ where $\Gamma\in{\mathcal{C}}$.
-Fix $11\leq j\leq g+$ 1. We define
+In order to provide bases of Kähler differentials for generic curves in $\mathcal{C}$, we shall need to use both the elementary $1$-forms dxdx and $dy$ and the approximate roots of $\Gamma$ where $\Gamma\in{\mathcal{C}}$.
+Fix $1\leq j\leq g+1$. We define
 $f_{j}=\prod_{k=0}^{\nu_{j-1}-1}(y-\sum_{\beta<\beta_{j}}e^{\frac{2\pi ik\beta}{n}}a_{\beta}x^{\beta/n}).$
@@
 By convention, we denote $\beta_{g+1}=\infty$ and we obtain $\Gamma_{<\beta_{g+1}}=\Gamma$.
-By construction, $f_{j,\Gamma}\circ\Gamma_{<\beta_{j}}\equiv 0$ holds for all $\Gamma\in\mathcal{C}$ and $11\leq j\leq g+$ 1.
+By construction, $f_{j,\Gamma}\circ\Gamma_{<\beta_{j}}\equiv 0$ holds for all $\Gamma\in\mathcal{C}$ and $1\leq j\leq g+1$.
 Remark 2.2.
@@
 3. Case $E=\emptyset$
-We divide our study into two cases: $E=\emptyset$, $E\neq\emptyset$.
+We divide our study into two cases: $\emptyset, EE\neq\emptyset$.
 The former will in fact contain all the important definitions and arguments, whereas the latter will just consist in a generalization in which everything except the bootstrapping process carries over without modification.
@@
 A $1$-form $\omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ is of level $l$ if $\omega\in M_{2{\nu}_{l}}$ for some $l\in\{0,\ldots,g\}$.
-All the $1$-forms we shall construct will be of some level $l$, as the reader will see. Certainly, $dx$ and $dy$ are both of level 00.
+All the $1$-forms we shall construct will be of some level $l$, as the reader will see. Certainly, $dx$ and $dy$ are both of level 0.
 3.2. Ordered families
@@
 section 1) for generic $\Gamma\in\mathcal{C}$ by Conditions (ii) and (iii).
-At level 00, the first condition will be accomplished by considering the 1-forms $dx$ and $df_{1}$, that constitute a separated family.
+At level 0, the first condition will be accomplished by considering the 1-forms $dx$ and $df_{1}$, that constitute a separated family.
 Then we describe a process in which $dx$ and $df_{1}$ appear alternatively and the approximate roots of $\Gamma$, with $\Gamma\in\mathcal{C}$, play a prominent role.
@@
 For now, a terminal family of level $l$ can be considered just as the product of our algorithm at the end of level $l$.
-We define $\mathcal{T}_{0}=\{dx,df_{1}\}$ as the terminal family of level 00.
+We define $\mathcal{T}_{0}=(dx,df_{1})$ as the terminal family of level 0.
 We are going to construct, iteratively, a terminal family $\mathcal{T}_{l+1}$ of level $l+1$ in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$, with $2{\nu}_{l+1}$ elements.
@@
 The construction of $\mathcal{T}_{l+1}$ requires several stages.
-The first one (stage 00) is elementary: we combine the $1$-forms in $\mathcal{T}_{l}$ with the approximate root $f_{l+1}$.
+The first one (stage 0) is elementary: we combine the $1$-forms in $\mathcal{T}_{l}$ with the approximate root $f_{l+1}$.
 More precisely, we define
@@
 Definition 3.5.
-Fix a terminal family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ We say that
-[…]
-is a the initial family (of stage 00) of level $l+1$.
+Fix a terminal family $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ . We say that
+$\mathcal{G}_{l+1}^{0}=(\Omega^{0}_{2a\nu_{l}+j})_{0\leq a<n_{l+1},\ 1\leq j\leq 2{\nu}_{l}}=(\Omega^{0}_{k})_{1\leq k\leq 2{\nu}_{l+1}}$
+is a the initial family (of stage 0) of level $l+1$.
 These $1$-forms belong to ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$, see equation (14), and are $2{\nu}_{l+1}$ in number. They will be called the stage 0 forms of level $l+1$.
@@
 Definition 3.6.
-Given $l\in{\mathbb{Z}}$ and $m\in{\mathbb{Z}}_{\geq 1}$, we denote by $[l]_{m}$ the class of $l$ modulo $m$. Moreover, given $S\subset{\mathbb{Z}}$, we define $[S]_{m}=\{[l]_{m}\ |\ \ l\in S\}$.
+Given $l\in{\mathbb{Z}}$ and $m\in{\mathbb{Z}}_{\geq 1}$, we denote by $l]_{m}$ the class of $l$ modulo $m$. Moreover, given $S\subset{\mathbb{Z}}$, we define $S]_{m}=\{[l]_{m}\ |\ \ l\in S\}$.
 Unfortunately, the stage 0 forms of level $l+1$ are ordered but usually not separated (condition (iii) in Definition 3.4).
@@
 Assume that $\mathcal{G}_{l+1}^{s}$ is non-separated.
-When this happens, we consider the pairs $j,k$ with $1\leq j<k\leq 2{\nu}_{l+1}$ such that $[\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$: whenever this equality holds, there is some $d\in\mathbb{Z}_{\geq 0}$ with $\nu_{\mathcal{C}}(\Omega^{s}_{k})=\nu_{\mathcal{C}}(\Omega_{j}^{s})+dn$.
+When this happens, we consider the pairs $j,k$ with $1\leq j<k\leq 2{\nu}_{l+1}$ such that $\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$: whenever this equality holds, there is some $d\in\mathbb{Z}_{\geq 0}$ with $\nu_{\mathcal{C}}(\Omega^{s}_{k})=\nu_{\mathcal{C}}(\Omega_{j}^{s})+dn$.
 Consider the pull-backs
@@
 (6) $\Omega^{s+1}_{k}:=C_{j,\eta_{j}}\Omega^{s}_{k}-C_{k,\eta_{k}}x^{d}\Omega^{s}_{j};$
-Setting also $\Omega^{s+1}_{j}=\Omega^{s}_{j}$, for the other $1$-form, we obtain the $2{\nu}_{l+1}$ $1$-forms in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ of stage $s+1$ (hence the exponent in $\Omega^{s+1}_{j}$), in the same level $l+1$.
+Setting also $\Omega^{s+1}_{j}=\Omega^{s}_{j}$, for the other $1$-form, we obtain the $2{\nu}_{l+1}$ 11-forms in ${\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ of stage $s+1$ (hence the exponent in $\Omega^{s+1}_{j}$), in the same level $l+1$.
 Remark 3.3.
@@
 Remark 3.4.
-We have $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})>\nu_{\mathcal{C}}(\Omega^{s}_{k})$ and $[\nu_{\mathcal{C}}(\Omega^{s+1}_{k})]_{n}\neq[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ as we shall see (Propositions 3.3 and 3.4).
+We have $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})>\nu_{\mathcal{C}}(\Omega^{s}_{k})$ and $\nu_{\mathcal{C}}(\Omega^{s+1}_{k})]_{n}\neq[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ as we shall see (Propositions 3.3 and 3.4).
 Remark 3.5.
-The algorithm described above to obtain a separated family $\mathcal{G}_{l+1}^{r}$ was called simple in section 1.
+The algorithm described above to obtain a separated family $\mathcal{G}_{l+1}^{r}$ was described as simple in section 1.
 One reason is that $\mathcal{G}_{l+1}^{0}$ is built from $\mathcal{T}_{l}$ and data related to the equisingularity class, namely the approximate root $f_{l+1}$.
@@
   • If $\beta\not\in\left\{\beta_{1},\ldots,\beta_{g}\right\}$ (i.e.
-it is not a Puiseux characteristic exponent) and $\ell=\max\{l\geq 0:\beta_{l}\leq\beta\}$, then $m=0$ and $[k]_{e_{\ell}}=[0]_{e_{\ell}}$.
-  • If $\beta=\beta_{\ell+1}$ (i.e. it is a Puiseux characteristic exponent), then $[k]_{e_{\ell}}=[m\beta]_{e_{\ell}}$.
+it is not a Puiseux characteristic exponent) and $\ell=\max\{l\geq 0:\beta_{l}\leq\beta\}$, then $m=0$ and $k]_{e_{\ell}}=[0]_{e_{\ell}}$.
+  • If $\beta=\beta_{\ell+1}$ (i.e. it is a Puiseux characteristic exponent), then $k]_{e_{\ell}}=[m\beta]_{e_{\ell}}$.
 The sequence $(c_{{\mathfrak{n}}^{r}(\beta)})_{r\geq 0}$ will be called the $c$-sequence of $g$.
@@
 belong to $\mathbb{C}[a_{9},a_{10},\ldots][[t]]$, where we skip the subindexes $\Gamma$ in the second expression.
-In the former case, $t\frac{\partial(f_{1}\circ\Gamma(t))}{\partial t}$ has leading variable $a_{9}$ of degree 00 and we have $k=0$, $e_{0}=6$, $\tilde{c}=9$, $R\equiv 1$, $Q_{9}\equiv Q_{10}\equiv\ldots\equiv 0$, $c_{9}=1$, $c_{10}=10/9$, $c_{11}=11/9$ and so on.
+In the former case, $t\frac{\partial(f_{1}\circ\Gamma(t))}{\partial t}$ has leading variable $a_{9}$ of degree 0 and we have $k=0$, $e_{0}=6$, $\tilde{c}=9$, $R\equiv 1$, $Q_{9}\equiv Q_{10}\equiv\ldots\equiv 0$, $c_{9}=1$, $c_{10}=10/9$, $c_{11}=11/9$ and so on.
 In the latter case, one can verify that $f_{2}^{2}\circ\Gamma(t)$ has leading variable $a_{10}$ of degree $1$.
@@
 The nonzero polynomial $R(a_{\beta_{0}},\ldots,a_{{\mathfrak{p}}(\beta)})$ is the same for all the coefficients $C_{\beta}$ (this is what makes the definition of the $c$-sequence possible).
-Also, despite saying that $a_{\beta}$ is of “degree $m$”, the first element $C_{\beta}$ is
+Also, despite saying that $a_{\beta}$ is of “degree mm”, the first element $C_{\beta}$ is
 $C_{\beta}=\tilde{c}a_{\beta}R(a_{\beta_{0}},\ldots,a_{{\mathfrak{p}}(\beta)})a_{\beta}^{m}+Q_{\beta}.$
-But the terminology reflects the fact that $a_{\beta}$ is the maximal variable in the coefficient of the least order (the “leader”), and if $m>0$, then it subsequently appears explicitly always to the power $m$ in $C_{{\mathfrak{n}}^{r}(\beta)}$, for any $r>0$.
+But the terminology reflects the fact that $a_{\beta}$ is the variable of maximal index in the coefficient of the least order (the “leader”), and if $m>0$, then it subsequently appears explicitly always to the power $m$ in $C_{{\mathfrak{n}}^{r}(\beta)}$, for any $r>0$.
 If $\beta$ is not a Puiseux characteristic exponent, then $m=0$.
@@
 This property, together with Remark 3.2, determine the $\mathcal{C}$-orders of all the $1$-forms appearing in the construction of a nice basis of $\mathcal{C}$.
-We obtain an algorithm to calculate such $\mathcal{C}$-orders without calculating the corresponding $1$-forms.
+We obtain Algorithm 1 that calculates such $\mathcal{C}$-orders without calculating the corresponding $1$-forms.
 This is the property of rich structure of the construction mentioned in section 1.
@@
 Assume $\Omega\in\hat{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ (resp.
-$f\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$) belongs to $\Delta_{\beta_{l},\leq}^{m}$ for some $1\leq l\leq g$ and $20\leq m\leq n_{l}-$ 2.
+$f\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$) belongs to $\Delta_{\beta_{l},\leq}^{m}$ for some $1\leq l\leq g$ and $0\leq m\leq n_{l}-2$.
 Then $\Omega\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$ (resp.
@@
 The proof for $f\in\mathbb{C}{[a_{\beta}]}_{\beta\in\mathcal{E}(\mathcal{C})}[[x,y]]$ is analogous but it can also be deduced from the case of $1$-forms by considering $df=\frac{\partial f}{\partial x}dx+\frac{\partial f}{\partial y}dy$.
-By definition, we know that $[k]_{e_{l-1}}=[m\beta_{l}]_{e_{l-1}}$. Let $\Gamma$ be a generic curve with irreducible parametrization
+By definition, we know that $k]_{e_{l-1}}=[m\beta_{l}]_{e_{l-1}}$. Let $\Gamma$ be a generic curve with irreducible parametrization
 $\Gamma(t)=\left(t^{n/e_{l-1}},\sum_{\beta_{0}\leq\beta<\beta_{l}}a_{\beta,\Gamma}t^{\beta/e_{l-1}}\right).$
@@
 This implies $C_{\beta_{l}}\neq 0$ for a generic curve as above.
-As the left hand side has order a multiple of $e_{l-1}$, we obtain $[0]_{e_{l-1}}=[k+\beta_{l}]_{e_{l-1}}$ and thus $(m+1)\beta_{l}$ is a multiple of $e_{l-1}$, which is impossible because $m+1$ is not a multiple of $n_{l}$.
+As the left hand side has order a multiple of $e_{l-1}$, we obtain $0]_{e_{l-1}}=[k+\beta_{l}]_{e_{l-1}}$ and thus $(m+1)\beta_{l}$ is a multiple of $e_{l-1}$, which is impossible because $m+1$ is not a multiple of $n_{l}$.
 ∎
@@
 This is the aim of the following technical lemmas.
-We first prove that product of several (say $a$) power series starting at a Puiseux characteristic exponent $\beta_{l}$ with degree 00 and $b$ power series starting at $\beta<\beta_{l}$ gives a power series with leading variable $\beta_{l}$ of degree $a-1$.
+We first prove that product of several (say $a$) power series starting at a Puiseux characteristic exponent $\beta_{l}$ with degree 0 and $b$ power series starting at $\beta<\beta_{l}$ gives a power series with leading variable $\beta_{l}$ of degree a1a-1.
 We can also control the $c$-sequence of such product.
@@
 implies that $k$ is a multiple of $n/\gcd(n,\beta_{1},\ldots,\beta_{l-1})=\nu_{l-1}$, we conclude that
-[…]
+$g_{i}(t)\in\bigcup_{\beta^{\prime}<\beta_{l}}\Delta^{0}_{\beta^{\prime}}\ \mathrm{for}\ 1\leq i\leq\nu_{l-1}-1.$
 The result is, thus, a direct consequence of Lemma 3.2, setting $a=m$ and $b=(\nu_{l-1}-1)m$, and the multinomial formula. ∎
@@
 Then $f_{l}^{m}\Omega\in\tilde{\Delta}_{\beta_{l},\leq}^{m-1}$ and the $c$-sequence of $f_{l}^{m}\Omega$ is equal to $1,m,m,\ldots$.
-Finally, for products of powers of approximate roots and $1$-forms with the same leading variable with degree 00, we can also control the outcome, and the $c$-sequence.
+Finally, for products of powers of approximate roots and $1$-forms with the same leading variable with degree 0, we can also control the outcome, and the $c$-sequence.
 The next result is a consequence of Corollary 3.1, Lemma 3.2 and Proposition 3.1.
@@
 Then $f_{l}^{m}\Omega\in\tilde{\Delta}_{\beta_{l},\leq}^{m}$ and the $c$-sequence $(c_{{\mathfrak{n}}^{r}(\beta_{l})}^{\prime})_{r\geq 0}$ of $f_{l}^{m}\Omega$ has $c_{\beta_{l}}^{\prime}=1$ and $c_{{\mathfrak{n}}^{r}(\beta_{l})}^{\prime}=c_{{\mathfrak{n}}^{r}(\beta_{l})}+m$ for $r\geq 1$ where $(c_{{\mathfrak{n}}^{r}(\beta_{l})})_{r\geq 0}$ is the $c$-sequence of $\Omega$.
-All these results will be relevant for what will be stage 00 of each level $l+1$, in which we shall make use of both the $1$-forms constructed in levels up to $l$, and the approximate root $f_{l+1}$.
+All these results will be relevant for what will be stage 0 of each level $l+1$, in which we shall make use of both the $1$-forms constructed in levels up to $l$, and the approximate root $f_{l+1}$.
 3.5. Terminal families
@@
 Definition 3.10.
-We say that $\mathcal{T}_{l}=\{\Omega_{1},\ldots\Omega_{2{\nu}_{l}}\}\subset{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ is a terminal family of level $l$ if the following conditions hold:
+We say that $\mathcal{T}_{l}=(\Omega_{1},\ldots\Omega_{2{\nu}_{l}})\subset{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ is a terminal family of level $l$ if the following conditions hold:
   1. (a)
@@
 Condition (b) indicates that the sequence of multiplicities “does not go too far” and indeed that all of them are less than or equal to the multiplicity of $df_{l+1}$ which is a form of level $l$ since it belongs to $\hat{M}_{2{\nu}_{l}}$.
-Condition (c) implies, on one hand, that the leading variables are not greater than $\beta_{l+1}$ for $l<g$, and that if they are $\beta_{l+1}$, then they appear with degree 00; and on the other hand, that the $c$-sequences are required to be strictly increasing.
+Condition (c) implies, on one hand, that the leading variables are not greater than $\beta_{l+1}$ for $l<g$, and that if they are $\beta_{l+1}$, then they appear with degree 0; and on the other hand, that the $c$-sequences are required to be strictly increasing.
 The last property is inherent to the study the semi-module of contacts of holomorphic $1$-forms; notice that Expression (3) describes the semigroup of the curves in $\mathcal{C}$} but $f_{0}^{k_{0}}f_{1}^{k_{1}}\cdots f_{r}^{k_{r}}\in{\Delta}_{\beta_{r},\leq}^{k_{r}-1}\setminus{\Delta}_{\beta_{r},<}^{k_{r}-1}$ if $r\geq 1$ and $k_{r}>0$ and thus those products do not have strictly increasing $c$-sequences.
 Condition (c) is used in the inductive steps of our construction.
-The family $\mathcal{T}_{0}=\{dx,df_{1}\}$ is terminal of level 00 because $\Omega_{1}\in\Delta_{n,<}$ and $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$. This is our starting point.
+The family $\mathcal{T}_{0}=(dx,df_{1})$ is terminal of level 0 because $\Omega_{1}\in\Delta_{n,<}$ and $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$. This is our starting point.
 Remark 3.12.
-In a terminal family ${\mathcal{T}}_{l}$ of level $l<g$, the multiplicity $\nu_{\mathcal{C}}(\Omega)$ of any $1$-form $\Omega\in\mathcal{T}_{l}$ belongs to $[\beta_{l+1}]_{e_{l}}$ or $[0]_{e_{l}}$ (because of their corresponding leading variables).
+In a terminal family ${\mathcal{T}}_{l}$ of level $l<g$, the multiplicity $\nu_{\mathcal{C}}(\Omega)$ of any $1$-form $\Omega\in\mathcal{T}_{l}$ belongs to $\beta_{l+1}]_{e_{l}}$ or $0]_{e_{l}}$ (because of their corresponding leading variables).
 This is a key property for our later arguments.
 More precisely, we have
-  • If $\Omega$ is of level $l-1$, then $\nu_{\mathcal{C}}(\Omega)\in(e_{l})$;
+  • If $\Omega$ is of level l1l-1, then $\nu_{\mathcal{C}}(\Omega)\in(e_{l})$;
   • Let be $\Omega$ of level $l$. Then $\Omega\in\cup_{\beta<\beta_{l+1}}\Delta_{\beta,<}\Leftrightarrow\nu_{\mathcal{C}}(\Omega)\in(e_{l})$;
-  • Let be $\Omega$ of level $l$. Then $\Omega\in\Delta_{\beta_{l+1},<}^{0}\Leftrightarrow\nu_{\mathcal{C}}(\Omega)\in[\beta_{l+1}]_{e_{l}}$
-Hence, a terminal family of level $l<g$ “covers” all the congruences modulo $n$ in $[(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ (cf. Definition 3.6).
+  • Let be $\Omega$ of level $l$. Then $\Omega\in\Delta_{\beta_{l+1},<}^{0}\Leftrightarrow\nu_{\mathcal{C}}(\Omega)\in[\beta_{l+1}]_{e_{l}}$ .
+Hence, a terminal family of level $l<g$ “covers” all the congruences modulo $n$ in $(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ (cf. Definition 3.6).
 Corollary 3.4.
-We have $[\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}=[(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ for $l<g$.
+We have $\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}=[(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ for $l<g$.
 Proof.
-This holds because, on one hand, there are $2\nu_{l}$ congruences modulo $n$ in the set $(e_{l})\cup\{\beta_{l+1}+(e_{l})\}$ ($\nu_{l}$ from $(e_{l})$ and other $\nu_{l}$ from $\beta_{l+1}+(e_{l})$); and on the other hand, $[\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}$ has $2\nu_{l}$ elements by definition of terminal family.
+This holds because, on one hand, there are $2\nu_{l}$ congruences modulo $n$ in the set $(e_{l})\cup\{\beta_{l+1}+(e_{l})\}$ ($\nu_{l}$ from $(e_{l})$ and other $\nu_{l}$ from $\beta_{l+1}+(e_{l})$); and on the other hand, $\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}$ has $2\nu_{l}$ elements by definition of terminal family.
 ∎
 3.5.2. From level $l$ to level $l+1$
-Suppose $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ is a terminal family of level $l$ for $0\leq l<g$.
+Suppose $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ is a terminal family of level $l$ for $0\leq l<g$.
 Our aim is to construct a terminal family of level $l+1$ from $\mathcal{T}_{l}$.
 This requires several stages.
-The first one (stage 00) is the construction of the initial family $\mathcal{G}_{l+1}^{0}$ of stage 00 (Definition 3.5).
+The first one (stage 0) is the construction of the initial family $\mathcal{G}_{l+1}^{0}$ of stage 0 (Definition 3.5).
 Lemma 3.3.
@@
 Remark 3.13.
-[…]
+Denote ${\mathfrak{S}}=\{(a,j):0\leq a<n_{l+1}\ \mathrm{and}\ 1\leq j\leq 2{\nu}_{l}\}$. Since
 $\nu_{\mathcal{C}}(f_{l+1}^{a}\Omega_{j})=a\overline{\beta}_{l+1}+\nu_{\mathcal{C}}(\Omega_{j})\in(e_{l+1})$
-for any $(a,j)\in{\mathfrak{S}}$, the set $[\nu_{\mathcal{C}}(\mathcal{G}_{l+1}^{0})]_{n}$ contains at most $\nu_{l+1}=n/e_{l+1}$ elements.
+for any $(a,j)\in{\mathfrak{S}}$, the set $\nu_{\mathcal{C}}(\mathcal{G}_{l+1}^{0})]_{n}$ contains at most $\nu_{l+1}=n/e_{l+1}$ elements.
 Notice also that, since $m\overline{\beta}_{l+1}\not\in(e_{l})$ for any $1\leq m<n_{l+1}$, we have
 $\nu_{\mathcal{C}}(f_{l+1}^{a}\Omega_{j})-\nu_{\mathcal{C}}(f_{l+1}^{b}\Omega_{k})\not\in(n)$
-if $(a,j),(b,k)\in{\mathfrak{S}}$, $(a,j)\neq(b,k)$ and $\nu_{\mathcal{C}}(\Omega_{j})$ and $\nu_{\mathcal{C}}(\Omega_{k})$ belong both to $(e_{l})$ or both to $[\beta_{l+1}]_{e_{l}}$.
-In particular any element of $[(e_{l+1})]_{n}$ has at most two representatives in $(\nu_{\mathcal{C}}({\Omega}_{j}^{0}))_{j=1}^{2\nu_{l+1}}$.
-Hence, $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{j=1}^{2\nu_{l+1}}$ defines exactly the $\nu_{l+1}$ classes modulo $n$ in $[(e_{l+1})]_{n}$ and each class has two representatives in it.
+if $\in{\mathfrak{S}}, (a,j)(b,k)(a,j)\neq(b,k)$ and $\nu_{\mathcal{C}}(\Omega_{j})$ and $\nu_{\mathcal{C}}(\Omega_{k})$ belong both to $(e_{l})$ or both to $\beta_{l+1}]_{e_{l}}$.
+In particular any element of $(e_{l+1})]_{n}$ has at most two representatives in $(\nu_{\mathcal{C}}({\Omega}_{j}^{0}))_{j=1}^{2\nu_{l+1}}$.
+Hence, $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{j=1}^{2\nu_{l+1}}$ defines exactly the $\nu_{l+1}$ classes modulo $n$ in $(e_{l+1})]_{n}$ and each class has two representatives in it.
 This property will allow us to choose, from each class modulo $n$, the $1$-form whose representative appears later in the sequence and substitute it for another one using the key formula (6), in order to obtain the $2{\nu}_{l+1}$ new forms of the next stage $1$.
@@
     We have
-    $\begin{split}\{[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}:j\in{\mathcal{L}}_{s}\}=[(e_{l+1})]_{n}&\ \ \mathrm{and}\\ \{[\nu_{\mathcal{C}}(\Omega_{j}^{s})-{\mathfrak{n}}^{s}(\beta_{l+1})]_{n}:j\in{\mathcal{U}}_{s}\}=[(e_{l+1})]_{n}.&\end{split}$
-    In particular $[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}$ if $j\neq k$ and both $j,k$ belong to ${\mathcal{L}}_{s}$ or both belong to ${\mathcal{U}}_{s}$.
+[…]
+    In particular $\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}$ if $j\neq k$ and both $j,k$ belong to ${\mathcal{L}}_{s}$ or both belong to ${\mathcal{U}}_{s}$.
 3.5.3. From stage $s$ to $s+1$ inside level $l+1$
@@
 Lemma 3.4.
-Assume that $[\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ for $j\neq k$, and that $\Omega^{s}_{k}\in\Delta_{{\mathfrak{n}}^{s}(\beta_{l+1}),<}$ and $\Omega^{s}_{j}\in\cup_{\beta^{\prime}<{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$.
+Assume that $\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ for $j\neq k$, and that $\Omega^{s}_{k}\in\Delta_{{\mathfrak{n}}^{s}(\beta_{l+1}),<}$ and $\Omega^{s}_{j}\in\cup_{\beta^{\prime}<{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$.
 Let $(c_{{\mathfrak{n}}^{r}(\beta_{l+1})})_{r\geq 0}$ be the $c$-sequence of $\Omega^{s}_{k}$.
@@
 Lemma 3.5.
-Suppose that $[\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}$ for $j<k$, and that $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ belong to $\tilde{\Delta}_{\beta_{l+1},\leq}$.
+Suppose that $\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}$ for $j<k$, and that $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ belong to $\tilde{\Delta}_{\beta_{l+1},\leq}$.
 Denote by $(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),j})_{r\geq 0}$ and $(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),k})_{r\geq 0}$ the $c$-sequences of $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ respectively.
@@
   • Let $1\leq j\leq 2\nu_{l}$ such that $\Omega_{j}\in\Delta_{\beta^{\prime},<}$ with $\beta^{\prime}=\beta_{l+1}$.
-Then $j+2(n_{l+1}-1)\nu_{l}\in{\mathcal{U}}_{1}$ and $j+2r\nu_{l}\in{\mathcal{L}}_{1}$ for $20\leq r\leq n_{l+1}-$ 2.
+Then $j+2(n_{l+1}-1)\nu_{l}\in{\mathcal{U}}_{1}$ and $j+2r\nu_{l}\in{\mathcal{L}}_{1}$ for $0\leq r\leq n_{l+1}-2$.
 Recall that the sets $M_{j}$ and $\hat{M}_{j}$ were defined in page 3.1.
@@
     $m_{k}=n_{l+1}-1$, $m_{j}=0$, $\nu_{\mathcal{C}}(\Omega_{b})\in[\beta_{l+1}]_{e_{l}}$ and $\nu_{\mathcal{C}}(\Omega_{a})\in(e_{l})$.
-Since any $[x]_{n}\in[(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{1\leq j\leq 2\nu_{l+1}}$ by Remark 3.13, we deduce ${\mathcal{U}}_{1}=\{1\leq r\leq 2\nu_{l+1}:\Omega_{r}^{1}\neq\Omega_{r}^{0}\}$.
+Since any $x]_{n}\in[(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{1\leq j\leq 2\nu_{l+1}}$ by Remark 3.13, we deduce ${\mathcal{U}}_{1}=\{1\leq r\leq 2\nu_{l+1}:\Omega_{r}^{1}\neq\Omega_{r}^{0}\}$.
 Next, let us show $\Omega_{j}^{1}\in\cup_{\beta^{\prime}\leq\beta_{l+1}}\Delta_{\beta^{\prime},<}\cap\hat{M}_{j}$, $\Omega_{k}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{l+1}),<}^{0}\cap\hat{M}_{k}$ and equation (18).
@@
 By construction, we have
-$[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{1})]_{n}$
+$\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{1})]_{n}$
 if $j\neq j^{\prime}$ and $j,j^{\prime}\in{\mathcal{L}}_{s}$.
-Thus $\{[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}:j\in{\mathcal{L}}_{1}\}$ and $[(e_{l+1})]_{n}$ are sets of cardinal $\nu_{l+1}$ such that the former is included in the latter and as a consequence they coincide.
-Analogously, $[\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k^{\prime}}^{0})]_{n}$ holds if $k\neq k^{\prime}$ and $k,k^{\prime}\in{\mathcal{U}}_{1}$.
-This property together with equation (18) imply that $\{[\nu_{\mathcal{C}}(\Omega_{k}^{1})-{\mathfrak{n}}(\beta_{l+1})]_{n}:k\in{\mathcal{U}}_{1}\}$ is a set of $\nu_{l+1}$ elements contained in $[(e_{l+1})]_{n}$ and hence they coincide.
+Thus $\{[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}:j\in{\mathcal{L}}_{1}\}$ and $(e_{l+1})]_{n}$ are sets of cardinal $\nu_{l+1}$ such that the former is included in the latter and as a consequence they coincide.
+Analogously, $\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k^{\prime}}^{0})]_{n}$ holds if $k\neq k^{\prime}$ and $k,k^{\prime}\in{\mathcal{U}}_{1}$.
+This property together with equation (18) imply that $\{[\nu_{\mathcal{C}}(\Omega_{k}^{1})-{\mathfrak{n}}(\beta_{l+1})]_{n}:k\in{\mathcal{U}}_{1}\}$ is a set of $\nu_{l+1}$ elements contained in $(e_{l+1})]_{n}$ and hence they coincide.
 ∎
@@
 As $\mathcal{G}_{l+1}^{s}$ is good, $\Omega_{j}^{s}\in\cup_{\beta^{\prime}\leq{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$ for any $1\leq j\leq 2{\nu}_{l+1}$.
-Since ${\mathfrak{n}}^{s}(\beta_{l+1})\in(e_{l+1})$ by hypothesis, any element of $[(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{s}))_{1\leq j\leq 2\nu_{l+1}}$ by property (IV) for $s$.
+Since ${\mathfrak{n}}^{s}(\beta_{l+1})\in(e_{l+1})$ by hypothesis, any element of $(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{s}))_{1\leq j\leq 2\nu_{l+1}}$ by property (IV) for $s$.
 Thus ${\mathcal{L}}_{s+1}$ and ${\mathcal{U}}_{s+1}$ have both $\nu_{l+1}$ elements.
@@
 By definition, the maximum $\mathcal{C}$}-order of a $1$-form in a terminal family $\mathcal{T}_{l}$ of level $l$ is $\overline{\beta}_{l+1}$.
-Since a $1$-form $\Omega_{k}^{0}$ of stage 00 and level $l+1$ is of the form $f_{l+1}^{a}\Omega$ where $0\leq a<n_{l+1}$ and $\Omega\in\mathcal{T}_{l}$, we obtain
+Since a $1$-form $\Omega_{k}^{0}$ of stage 0 and level $l+1$ is of the form $f_{l+1}^{a}\Omega$ where $0\leq a<n_{l+1}$ and $\Omega\in\mathcal{T}_{l}$, we obtain
 $\nu_{\mathcal{C}}(\Omega_{k}^{0})\leq\overline{\beta}_{l+1}+(n_{l+1}-1)\overline{\beta}_{l+1}=n_{l+1}\overline{\beta}_{l+1}$
@@
 Condition (b) is trivially satisfied.
-Let $\mathcal{B}_{s}$ be the set of congruences $[x]_{n}$ modulo $n$ such that there exist $\Omega^{s}_{j},\Omega_{k}^{s}$ with $[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}=[x]_{n}$ and $j<k\leq n$.
+Let $\mathcal{B}_{s}$ be the set of congruences $x]_{n}$ modulo $n$ such that there exist $\Omega^{s}_{j},\Omega_{k}^{s}$ with $\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}=[x]_{n}$ and $j<k\leq n$.
 We claim that $\mathcal{G}_{g}^{s}$ is terminal if and only if $\mathcal{B}_{s}=\emptyset$.
@@
 Since this implies
-[…]
+$\left\{[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\ |\ 1\leq j\leq n\right\}=n,$
 ${\mathcal{G}}_{g}^{s}$ is separated. Condition (c) is again a consequence of property (III).
@@
 3.6. The terminal family of level $g$
-Thus, beginning with the $1$-forms $\Omega_{1}=dx$ and $\Omega_{2}=df_{1}$ of level 00, we obtain a terminal family of $1$-forms $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{2n})$ of level $g$ by applying $g$ times Proposition 3.6.
+Thus, beginning with the $1$-forms $\Omega_{1}=dx$ and $\Omega_{2}=df_{1}$ of level 0, we obtain a terminal family of $1$-forms $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{2n})$ of level $g$ by applying $g$ times Proposition 3.6.
+The next result implies Theorem 1.1 by Remark 3.1.
 Theorem 3.1.
@@
 by Proposition 3.7, we obtain $\nu_{\mathcal{C}}(\Omega_{2\nu_{l}})=\overline{\beta}_{l+1}$. Thus, we could replace $\Omega_{2\nu_{l}}$ with $df_{l+1}$ in the nice basis.
-Remark 3.18.
-Since $\nu_{\mathcal{C}}(f_{l})=\overline{\beta}_{l}$ for $1\leq l\leq g$, it is easy to build an algorithm that provides $\Lambda_{\Gamma}$, for generic $\Gamma\in\mathcal{C}$, and whose input is $n,\beta_{1},\ldots,\beta_{g}$, or equivalently $n,\overline{\beta}_{1},\ldots,\overline{\beta}_{g}$, by Remark 3.8.
 4. Case $E\neq\emptyset$
@@
 A family $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{r})$ in $\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, with $r\in\{2\nu_{l},\min(2\nu_{l},n)\}$, is $E$-ordered of level $0\leq l\leq g$ for $\Gamma$ if:
-  • $E$-(i)
+  • EE-(i)
     $\Omega_{j}\in\hat{M}_{j}^{E}$ for $1\leq j\leq r$;
-  • $E$-(ii)
+  • EE-(ii)
     $\nu_{C}(\Omega_{k})\geq\nu_{C}(\Omega_{j})$ for $1\leq j\leq k\leq r$.
 Moreover, $\mathbf{\Omega}$ is $E$-separated of level $l$ if it is ordered and
-  • $E$-(iii)
+  • EE-(iii)
     $\nu_{\mathcal{C}}(\Omega_{1}),\ldots,\nu_{\mathcal{C}}(\Omega_{\min(2\nu_{l},n)})$ define pairwise different classes modulo $n$.
 Finally, for $l=g$ and $r=n$, $\mathbf{\Omega}$ is an $E$-nice basis if, moreover,
-  • $E$-(iv)
+  • EE-(iv)
     $\hat{\Omega}_{E}(\mathbb{C}^{2},0)=\mathbb{C}[[x]]\Omega_{1,\Gamma}+\cdots+\mathbb{C}[[x]]\Omega_{n,\Gamma}+\mathcal{I}_{E,\Gamma}$ for generic $\Gamma\in\mathcal{C}$.
-Notice that the only modification is to require that $E$ is invariant, in $E$-(i) and $E$-(ii).
+Notice that the only modification is to require that $E$ is invariant, in EE-(i) and EE-(ii).
 The definitions of terminal and good need also adapting to $E$.
@@
 Definition 4.4.
-A family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}\subset\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, with $l_{0}\leq l\leq g$, is $E$-terminal if
-  • $E$-(a)
+A family $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})\subset\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, with $l_{0}\leq l\leq g$, is $E$-terminal if
+  • EE-(a)
     $\mathcal{T}_{l}$ is an $E$-separated family of level $l$;
-  • $E$-(b)
+  • EE-(b)
     On one hand $\nu_{\mathcal{C}}(\Omega_{2{\nu}_{l_{0}}})\leq n+\overline{\beta}_{l_{0}+1}$ if $l_{0}<g$, and on the other,
@@
     (this inequality the same as in Definition 3.10 except for the initial level $l_{0}$).
-  • $E$-(c)
+  • EE-(c)
     The same as (c) in Definition 3.10: $\mathcal{T}_{l}\subset\cup_{\beta^{\prime}\in{\mathcal{E}}({\mathcal{C}})}\Delta_{\beta^{\prime},<}$.
@@
 Definition 4.5.
-Given an $E$-terminal family $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ of level $0\leq l<g$, we define an $E$-initial family ${\mathcal{G}}_{l+1}^{0}=(\Omega_{1}^{0},\ldots,\Omega_{2\nu_{l+1}}^{0})$ of stage 00 (in level $l+1$) by means of equation (5), exactly as in the case $E=\emptyset$.
+Given an $E$-terminal family $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ of level $0\leq l<g$, we define an $E$-initial family ${\mathcal{G}}_{l+1}^{0}=(\Omega_{1}^{0},\ldots,\Omega_{2\nu_{l+1}}^{0})$ of stage 0 (in level $l+1$) by means of equation (5), exactly as in the case $E=\emptyset$.
 Let $\mathcal{G}_{l+1}^{s}=(\Omega_{1}^{s},\ldots,\Omega^{s}_{2{\nu}_{l+1}})$ be given with $\Omega_{i}^{s}\in\Omega_{E,\mathcal{C}}(\mathbb{C}^{2},0)$, for $0\leq l<g$ and $s\geq 1$.
@@
 This is what we carry out in what follows.
-When $E=\emptyset$, the terminal family of level 00 was ${\mathcal{T}}_{0}=\{dx,df_{1}\}$.
+When $E=\emptyset$, the terminal family of level 0 was ${\mathcal{T}}_{0}=(dx,df_{1})$.
 From this family, applying the results in subsections 3.5.2 and 3.5.3 we ended up with the terminal family ${\mathcal{T}}_{g}$.
@@
 Thus, we only need to find an $E$-terminal family of some level $l_{0}$.
-Actually, it suffices to find an $E$-good family ${\mathcal{G}}_{l_{0}}^{1}$ (with $E$-(I) being empty) since in such a case the results in subsection 3.5.3 provide an $E$-terminal family of level $l_{0}$.
+Actually, it suffices to find an $E$-good family ${\mathcal{G}}_{l_{0}}^{1}$ (with EE-(I) being empty) since in such a case the results in subsection 3.5.3 provide an $E$-terminal family of level $l_{0}$.
 The rest of this section is devoted to providing this initial $E$-terminal family.
@@
 which satisfy $\nu(\Omega_{1})=0$, $\nu_{\mathcal{C}}(\Omega_{1})=n$, $\nu(\Omega_{2})=1$, $\nu_{\mathcal{C}}(\Omega_{2})=\beta_{1}+n$.
-The family $\mathcal{T}_{0}=\{\Omega_{1},\Omega_{2}\}$ is clearly $E$-terminal of level $l_{0}=0$ and thus, we obtain an $E$-nice basis $\mathbf{\Omega}$ by Remark 4.1.
+The family $\mathcal{T}_{0}=(\Omega_{1},\Omega_{2})$ is clearly $E$-terminal of level $l_{0}=0$ and thus, we obtain an $E$-nice basis $\mathbf{\Omega}$ by Remark 4.1.
 4.2.2. Case $\delta_{y}=1$ and $\beta_{0}\in(n)$
@@
 Applying Lemma 3.5, with $\beta_{l+1}=\beta_{0}$, we obtain $\Omega_{1}^{1}=\Omega_{1}^{0}$ and $\Omega_{2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{0}),<}^{0}$ where $\nu_{\mathcal{C}}(\Omega_{2}^{1})=n+{\mathfrak{n}}(\beta_{0})$.
-The family ${\mathcal{G}}_{0}^{1}=\{\Omega_{1}^{1},\Omega_{2}^{1}\}$ is an $E$-good family of level 00.
+The family ${\mathcal{G}}_{0}^{1}=(\Omega_{1}^{1},\Omega_{2}^{1})$ is an $E$-good family of level 0.
 Moreover, it satisfies $\nu(\Omega_{1}^{1})=\delta_{x}$ and $\nu(\Omega_{2}^{1})=1$.
@@
 Recall Definition 4.1 for $\overline{\Omega}_{1},\overline{\Omega}_{2}$ and $\overline{\Omega}_{j}$.
-As $\overline{\Omega}_{1}=x^{\delta_{x}}dy$ and $\overline{\Omega}_{2}=ydx$, we have $[\nu_{C}(\overline{\Omega}_{1})]_{n}=[\nu_{C}(\overline{\Omega}_{2})]_{n}=[\beta_{1}]_{n}$, so that $\{\overline{\Omega}_{1},\overline{\Omega}_{2}\}$ is not an $E$-separated family of level 00.
+As $\overline{\Omega}_{1}=x^{\delta_{x}}dy$ and $\overline{\Omega}_{2}=ydx$, we have $\nu_{C}(\overline{\Omega}_{1})]_{n}=[\nu_{C}(\overline{\Omega}_{2})]_{n}=[\beta_{1}]_{n}$, so that $\{\overline{\Omega}_{1},\overline{\Omega}_{2}\}$ is not an $E$-separated family of level 0.
 Using Lemma 3.5 and Proposition 3.3 we obtain $\Omega_{1}^{1}=\overline{\Omega}_{1}$ and $\Omega_{2}^{1}\in\Delta_{{\mathfrak{n}}(\beta_{1}),<}^{0}$ with $\nu_{\mathcal{C}}(\Omega_{2}^{1})=n+{\mathfrak{n}}(\beta_{1})$.
-This family ${\mathcal{G}}_{0}^{1}=\{\Omega_{1}^{1},\Omega_{2}^{1}\}$ is $E$-separated of level 00 but does not satisfy $E$-(b) or $E$-(c) in Definition 4.4, so it is not an $E$-terminal family of level 00.
+This family ${\mathcal{G}}_{0}^{1}=(\Omega_{1}^{1},\Omega_{2}^{1})$ is $E$-separated of level 0 but does not satisfy EE-(b) or EE-(c) in Definition 4.4, so it is not an $E$-terminal family of level 0.
 This is why, in this case, we need to start at $l_{0}=1$.
 By Remark 4.1, we just need to provide an $E$-good family of level $1$.
-Consider the $E$-ordered family ${\mathcal{G}}_{1}^{0}=\{\Omega_{1}^{0},\ldots,\Omega_{2{\nu}_{1}}^{0}\}$ given by $\Omega_{j}^{0}=\overline{\Omega}_{j}$ for $1\leq j\leq 2{\nu}_{1}$, which satisfies $\nu(\Omega_{j}^{0})=\lfloor\frac{j+\delta_{x}}{2}\rfloor$ for $1\leq j\leq 2{\nu}_{1}$.
+Consider the $E$-ordered family ${\mathcal{G}}_{1}^{0}=(\Omega_{1}^{0},\ldots,\Omega_{2{\nu}_{1}}^{0})$ given by $\Omega_{j}^{0}=\overline{\Omega}_{j}$ for $1\leq j\leq 2{\nu}_{1}$, which satisfies $\nu(\Omega_{j}^{0})=\lfloor\frac{j+\delta_{x}}{2}\rfloor$ for $1\leq j\leq 2{\nu}_{1}$.
 Note that
-(23) $[\nu_{C}(\overline{\Omega}_{2j+1})]_{n}=[\nu_{C}(\overline{\Omega}_{2j+2})]_{n}=[(j+1)\beta_{1}]_{n}.$
+(23) $\nu_{C}(\overline{\Omega}_{2j+1})]_{n}=[\nu_{C}(\overline{\Omega}_{2j+2})]_{n}=[(j+1)\beta_{1}]_{n}.$
 if $1\leq 2j+1<2j+2\leq 2{\nu}_{1}$.
-All congruences $[\beta_{1}]_{n},[2\beta_{1}]_{n},\ldots,[\nu_{1}\beta_{1}]_{n}$ are pairwise different, so the unique situation in which two $1$-forms have congruent ${\mathcal{C}}$-orders is described in equation (23).
+All congruences $\beta_{1}]_{n},[2\beta_{1}]_{n},\ldots,[\nu_{1}\beta_{1}]_{n}$ are pairwise different, so the unique situation in which two $1$-forms have congruent $\mathcal{C}$}-orders is described in equation (23).
 Since $dy\in\tilde{\Delta}_{\beta_{1},<}^{0}$, we have $\Omega_{2j+1}^{0}\in\tilde{\Delta}_{\beta_{1},<}$ by Corollary 3.3.
@@
 for $2j+2\leq 2{\nu}_{1}$.
-As a consequence, the family ${\mathcal{G}}_{1}^{1}=\{\Omega_{1}^{1},\ldots,\Omega_{2{\nu}_{1}}^{1}\}$ is $E$-good of level $1$ (with $E$-(I) being empty).
+As a consequence, the family ${\mathcal{G}}_{1}^{1}=(\Omega_{1}^{1},\ldots,\Omega_{2{\nu}_{1}}^{1})$ is $E$-good of level $1$ (with EE-(I) being empty).
 Note that ${\mathcal{L}}_{1}$ is the subset of odd numbers of $\{1,\ldots,2\nu_{1}\}$ whereas ${\mathcal{U}}_{1}$ is its subset of even numbers.
@@
 The proof is the same.
-Remark 4.3.
-The set $\Lambda_{\mathcal{C}}^{E}$ (Definition 2.7) depends on $n,\beta_{0},\beta_{1},\ldots,\beta_{g}$ and indeed it is straightforward to write an algorithm whose input is $(\delta_{x},\delta_{y},n,\beta_{0},\ldots,\beta_{g})$ and whose output is $\Lambda_{\mathcal{C}}^{E}$.
-The argument is analogous to the one in Remark 3.18.
+4.4. An explicit algorithm
+As a byproduct of our work, we present Algorithm 1 in pseudo-code which, given the sequence $(n,\beta_{1},\ldots,\beta_{g})$ of the multiplicity and the Puiseux exponents of an equisingularity class $\mathcal{C}$}, returns the orders $a(1),a(2),\ldots,a(n)$ corresponding to the nice (Apery) basis, $a(i)=\nu_{{\mathcal{C}}}(\Omega_{i})$, for $i=1,\ldots,n$.
+In particular, the algorithm provides the generic semimodule $\Lambda_{\mathcal{C}}$ and gives Theorem 1.2.
+Indeed, the powerful properties of our construction allow to describe $\Lambda_{\mathcal{C}}$ without calculating explicitly a nice basis.
+Notice how, despite the elaborate nature of the previous arguments, the algorithm is straightforward and is coding very simple.
+Note that the value $b(i)$ is the leading variable of $\Omega_{i}$.
+The algorithm provides ${\Lambda}_{\mathcal{C}}$ but also gives ${\Lambda}_{\mathcal{C}}^{E}$ for $E=\{x^{\delta_{x}}y^{\delta_{y}}=0\}$ by making the following minor changes:
+  • For $\delta_{y}=0$, we consider $a(1)=b(1)=n$, $a(2)=n\delta_{x}+\beta_{1}$ and $b(2)=\beta_{1}$;
+  • For $\delta_{y}=1$, we consider $a(1)=n\delta_{x}+\beta_{0}$, $b(1)=\beta_{0}$, $a(2)=n+\beta_{1}$ and $b(2)=\beta_{1}$.
+Algorithm 1 Computation of the orders of an Apery basis
+Input: $\beta_{0}=n,\beta_{1},\ldots,\beta_{g}$
+Values corresponding to $\mathcal{T}_{0}$:
+$a(1)\leftarrow n,\,b(1)\leftarrow n,\,a(2)\leftarrow\beta_{1},\,b(2)\leftarrow\beta_{1}$
+for $l=0,\ldots,g-1$ do
+  Compute the orders corresponding to $\mathcal{G}^{0}_{l+1}$:
+  for $k=1,\ldots,n_{l+1}$ do
+   for $s=1,\ldots,2\nu_{l}$ do
+     $a(2k\nu_{l}+s)\leftarrow\overline{\beta}_{l+1}k+a(s)$
+     $b(2k\nu_{l}+s)\leftarrow\beta_{l+1}$
+   end for
+  end for
+  $d\leftarrow\beta_{l+1},\,c\leftarrow 1$
+  while $d<\beta_{l+2}\land c\neq 0$ do
+   Compute the orders corresponding to $\mathcal{G}^{r+1}_{l+1}$ from those of $\mathcal{G}^{r}_{l+1}$:
+   $c\leftarrow 0$
+   while $\exists k<s\,\big|(a(s)-a(k)\in(n)\land\max(b(s),b(k))=d)$ do
+     $a(s)\leftarrow a(s)+{\mathfrak{n}}(d)-d$
+     $b(s)\leftarrow{\mathfrak{n}}(d)$
+     if $s\leq n$ then
+      $c\leftarrow c+1$
+     end if
+   end while
+   $d\leftarrow\mathfrak{n}(d)$
+  end while
+end for
+Output: $a(1),\ldots,a(n)$, the orders of the nice (Apery) basis
 5. Families of equisingular curves
@@
 Define $(\Gamma_{0}^{\epsilon})=(\Gamma^{\epsilon})$ and $(\Gamma_{i}^{\epsilon})$ the strict transform of $(\Gamma^{\epsilon})$ in $\mathcal{X}_{i}$.
-As $(\Gamma^{\epsilon})$ is non-constant, there exists an integer $\iota(\Gamma^{\epsilon})\geq 1$ such that $\Gamma_{i}^{\epsilon}$ passes through $P_{i}$ for all $0\leq i<\iota(\Gamma^{\epsilon})$, and $\epsilon$ in a neighborhood of 00 but $\Gamma_{\iota}^{\epsilon}$ does not pass through $P_{\iota(\Gamma^{\epsilon})}$ for any $\epsilon$ in a neighborhood of 00.
+As $(\Gamma^{\epsilon})$ is non-constant, there exists an integer $\iota(\Gamma^{\epsilon})\geq 1$ such that $\Gamma_{i}^{\epsilon}$ passes through $P_{i}$ for all $0\leq i<\iota(\Gamma^{\epsilon})$, and $\epsilon$ in a neighborhood of 0 but $\Gamma_{\iota}^{\epsilon}$ does not pass through $P_{\iota(\Gamma^{\epsilon})}$ for any $\epsilon$ in a neighborhood of 0.
 Definition 5.3.
@@
 $\theta(\Gamma^{\epsilon}):=\theta(\Gamma_{0}^{\epsilon})=\hat{\theta}(\iota)$
-where an empty sum is 00.
+where an empty sum is 0.
 Proof.
@@
 Denote $\mu_{j}=\nu_{P_{j}}(\Gamma_{j})$ for $j\geq 0$ and consider the family $(\Gamma_{0}^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ associated to $\iota^{\prime}$ in the proof of Lemma 5.2 with $\Gamma^{0}_{0}=\Gamma$.
-Since the sliding divisor $\iota(\Gamma_{\iota-1}^{\epsilon})$ is greater than $1$ by definition, we have, by Remark 5.4:
+Since the sliding divisor $\iota(\Gamma_{\iota-1}^{\epsilon})$ is greater than 11 by definition, we have, by Remark 5.4:
 $\theta(\Gamma_{\iota-1}^{\epsilon})\in{\mathcal{E}}_{E_{\iota-1}}(\Gamma_{\iota-1}^{0})\setminus\{\mu_{\iota-1}\}$
@@
 We denote by ${\mathfrak{X}}_{E,i}^{\circ}$ the subset of ${\mathfrak{X}}_{E,i}$ consisting of vector fields singular at $P_{i}$.
-Given $X\in{\mathfrak{X}}_{E,i}$, with $X(P_{i})=0$ or $i\geq 1$, such that $\Gamma_{i}$ is not $X$-invariant, we consider the flow $(\phi^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of $X$ in a neighborhood of $P_{i}$ and define the curve $\Gamma_{i}^{\epsilon}=\phi^{\epsilon}(\Gamma_{i})$ for $\epsilon\in\mathbb{C}$ in a neighborhood of 00.
-The pair $(\Gamma_{i}^{\epsilon},E_{i})$ is in the same equisingularity class as $(\Gamma_{i},E_{i})$ for $\epsilon$ in a neighborhood of 00.
+Given $X\in{\mathfrak{X}}_{E,i}$, with $X(P_{i})=0$ or $i\geq 1$, such that $\Gamma_{i}$ is not $X$-invariant, we consider the flow $(\phi^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ of $X$ in a neighborhood of $P_{i}$ and define the curve $\Gamma_{i}^{\epsilon}=\phi^{\epsilon}(\Gamma_{i})$ for $\epsilon\in\mathbb{C}$ in a neighborhood of 0.
+The pair $(\Gamma_{i}^{\epsilon},E_{i})$ is in the same equisingularity class as $(\Gamma_{i},E_{i})$ for $\epsilon$ in a neighborhood of 0.
 By blowing-down we obtain the family $(\Gamma^{\epsilon})_{\epsilon\in({\mathbb{C}},0)}$ induced by $X$, which is an analytic family in the equisingularity class of $\Gamma\in{\mathcal{C}}$.
@@
 In this case, $l_{0}=0$ for $\mathcal{C}$ and ${E}_{1}=\{x_{1}=0\}$.
-A terminal family ${\mathcal{T}}_{0}$ of level 00 of $\mathcal{C}$ is $(\Omega_{1},\Omega_{2})=(dx,x^{\delta_{x}}df_{1})$ whereas a terminal family $\tilde{\mathcal{T}}_{0}$ of level 00 of ${\mathcal{C}}_{1}$ is $(\tilde{\Omega}_{1},\tilde{\Omega}_{2})=(dx_{1},x_{1}df_{1,1})$ (see subsection 5.1.1).
+A terminal family ${\mathcal{T}}_{0}$ of level 0 of $\mathcal{C}$ is $(\Omega_{1},\Omega_{2})=(dx,x^{\delta_{x}}df_{1})$ whereas a terminal family $\tilde{\mathcal{T}}_{0}$ of level 0 of ${\mathcal{C}}_{1}$ is $(\tilde{\Omega}_{1},\tilde{\Omega}_{2})=(dx_{1},x_{1}df_{1,1})$ (see subsection 5.1.1).
 Of course, $dx,dx_{1}\in\Delta_{n,<}$ and $df_{1},x_{1}d{f}_{1,1}\in\Delta_{\beta_{1},<}^{0}$, and also
@@
 6.2.3. $\delta_{y}=1$ and $\beta_{0}=n$
-Again, $l_{0}=0$ for $\mathcal{C}$, $E=\{xy=0\}$ and ${E}_{1}=\{x_{1}=0\}$.
+Again, we have $l_{0}=0$ for $\mathcal{C}$, $E=\{xy=0\}$ and ${E}_{1}=\{x_{1}=0\}$.
 As described in subsubsection 4.2.2, ${\mathcal{T}}_{0}=(\Omega_{1},\Omega_{2})$ where $\Omega_{1}=xdy\in\Delta_{n,<}^{0}$, $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$, $\nu_{\mathcal{C}}(\Omega_{1})=2n$ and $\nu_{\mathcal{C}}(\Omega_{2})=n+\beta_{1}$.
@@
 Take $1\leq k\leq 2{\nu}_{l+1}$, and assume $\Omega_{k}^{1}\neq\Omega_{k}^{0}$.
-Then there exists $1\leq j<k$ such that $[\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}$.
+Then there exists $1\leq j<k$ such that $\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}$.
 Since
-$[\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}=[\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{k}^{0})]_{n},$
+$\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{j}^{0})]_{n}=[\nu_{{\mathcal{C}}}({\Omega}_{k}^{0})]_{n}=[\nu_{{\mathcal{C}}_{1}}(\tilde{\Omega}_{k}^{0})]_{n},$
 we deduce that $\tilde{\Omega}_{k}^{1}\neq\tilde{\Omega}_{k}^{0}$.
@@
 6.4. The dimension formula
-The next formula is proved in [13]. It follows also from all our previous results.
-Theorem 6.1.
-Let $\Gamma$ be a germ of irreducible curve defined in a neighborhood of $0\in\mathbb{C}^{2}$.
-The dimension of the generic component of the analytic moduli of the equisingularity class of $\Gamma$ is equal to
-$\sum_{0\leq i<\tau}\sigma(\nu_{P_{i}}(\Gamma_{i})+\delta_{i}),$
-where $P_{0},\ldots,P_{\tau}$ are infinitely near points of $\Gamma$, $\tau$ is the first index such that $\Gamma_{\tau}$ is smooth, and $\delta_{i}$ is the number of irreducible components of $E_{i}$ passing through $P_{i}$.
-Proof.
-Notice that $\sigma(2)=\sigma(3)=0$, so that we do not need to assume that $\Gamma_{\tau}$ is transverse to $E_{\tau}$, just that it is non-singular.
+Now, we prove Theorem 1.3.
+Proof of Theorem 1.3.
 Up to replacing $\Gamma$ with another curve in its equisingularity class $\mathcal{C}$}, we can assume that $\Gamma_{i}$ is generic in the equisingularity class of $(\Gamma_{i},E_{i})$ for any $0\leq i<\tau$ by Remark 5.8 .
-Now, Propositions 5.4 and 6.1 imply the result.
+Now, Remark 5.11 and Propositions 5.4 and 6.1 imply the result.
 ∎